### v1

# Xuezhang Chen  and Shihong Zhang

Abstract: We introduce the notion of a biharmonic Poisson kernel associated with certain pair of conformal boundary operators and present its explicit formula. With this powerful tool, we next establish the classification theorems of nonnegative solutions to a biharmonic equation on the upper half-space and unit ball with proper pairs of conformal boundary operators.
$\mathbf{2020}$ MSC: 35B53,53C18 (35G30,35J08)
Keywords: Conformal boundary operators, biharmonic Poisson kernel, biharmonic Green function.
1 Introduction
Caffarelli-Gidas-Spruck [5] established the celebrated Liouville-type theorem: For $n\geq 2$, a $C^{2}$ solution of
$-\Delta u=(n+1)(n-1)u^{\frac{n+3}{n-1}},~{}~{}u>0\qquad\mathrm{~{}~{}in~{}~{}}\mathbb{R}^{n+1}$
is of the form
$u(X)=\left(\frac{\varepsilon}{\varepsilon^{2}+|X-X_{0}|^{2}}\right)^{\frac{n-1}{2}},\qquad\varepsilon\in\mathbb{R}_{+},X_{0}\in\mathbb{R}^{n+1}.$
Under an additional hypothesis $u(X)=O(|X|^{1-n})$ as $|X|\to\infty$, the same result was also proved earlier by Obata [28] and Gidas-Ni-Nirenberg [16]. Let $\mathbb{R}_{+}^{n+1}=\{X=(x,t)|~{}x\in\mathbb{R}^{n},t\in\mathbb{R}_{+}\}$ denote the upper half-space with boundary $\partial\mathbb{R}_{+}^{n+1}=\mathbb{R}^{n}$ and $\mathbb{B}^{n+1}$ denote the unit ball in $\mathbb{R}^{n+1}$. Escobar [13] used the Obata-type argument to characterize minimizers of a sharp Sobolev trace inequality on $\mathbb{B}^{n+1}$, via a conformal map between $\mathbb{R}_{+}^{n+1}$ and $\mathbb{B}^{n+1}$, which is equivalent to seeking smooth solutions to
$\begin{cases}\displaystyle\Delta u=0,\quad u>0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{R}_{+}^{n+1},\\ \displaystyle-\partial_{t}u(x,0)=(n-1)u^{\frac{n+1}{n-1}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\end{cases}$ (1.1)
under an extra hypothesis that $u(X)=O(|X|^{1-n})$ as $|X|\to\infty$. Y. Y. Li-Zhu [26] (see also, Y. Y. Li-Zhang [25]) employed the moving sphere method to prove the following Liouville-type theorem without the above hypothesis: A $C^{2}$ solution of (1.1) is of the form
$u(x,t)=\left(\frac{\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-1}{2}},\qquad\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}.$
Such Liouville-type theorems of general critical growth second order semilinear elliptic equations on $\mathbb{R}_{+}^{n+1}$ are referred to Escobar [14], Y. Y. Li-Zhang [25] etc.
Sun-Xiong [30] classified nonnegative smooth solutions of the following conformally invariant polyharmonic boundary value problem: For $m\geq 2,n+1>2m$ and $1<p\leq\frac{n+2m-1}{n+1-2m}$,
$\begin{cases}\displaystyle\Delta^{m}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\partial_{t}\Delta^{k}u(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle(-1)^{m}\partial_{t}\Delta^{m-1}u(x,0)=u^{p}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\end{cases}$ (1.2)
where $20\leq k\leq m-$ 2. When $m$ is even, assume in addition that $u(X)=o(|X|^{2m-1})$ as $|X|\to\infty$, then for the critical exponent $p=\frac{n+2m-1}{n-(2m-1)}$,
$u(x,t)=$ $c(n,m)\int_{\mathbb{R}^{n}}\frac{t^{2m-1}}{(|x-y|^{2}+t^{2})^{\frac{n+2m-1}{2}}}\left(\frac{\lambda}{\lambda^{2}+|y|^{2}}\right)^{\frac{n-2m+1}{2}}\mathrm{d}y+\sum_{k=1}^{m-1}t^{2k}P_{2k}(x),$
where $x_{0}\in\mathbb{R}^{n},\lambda\in\mathbb{R}_{+}$, $c(n,m)$ is a normalized constant and $P_{2k}(x)$ denotes a polynomial of degree $2k$ in the variable $x$; for $1<p<\frac{n+2m-1}{n+1-2m}$,
$u(x,t)=\sum_{k=1}^{m-1}t^{2k}P_{2k}(x).$
A key step of their approach is to improve the integrability of a Kevin transformation $u^{\ast}$ for $(-\Delta)^{m}$ via an iteration method, and eventually to show that the singularity of $u^{\ast}$ can be removed at $t=0$; in the next step, the integral equation of $u^{\ast}$ on $\partial\mathbb{R}_{+}^{n+1}$ is solved by the classification theorems due to Chen-Li-Ou [9] and Y. Y. Li [24], then the standard method can give the classification of nonnegative solutions to (1.2).
Conformal boundary operators $\mathscr{B}_{k}^{3}$ for $k\in\{0,1,2,3\}$ associated with the Paneitz operator have attracted a lot of attention, see [18, 4, 6, 11] etc. In the study of zeta functional determinants on a four-manifold with boundary, Chang-Qing [8] discovered a third order conformally covariant boundary operator $\mathscr{B}_{3}^{3}$. The discovery of conformal boundary operators $\mathscr{B}_{k}^{3}$ to an $(n+1)$-manifold $(X^{n+1},g)$ with boundary $\partial X^{n+1}=M^{n}$ is due to Branson-Gover [3], Juhl [23], Gover-Peterson [19], Grant [20], Stafford [29] etc. The Paneitz operator $P_{4}^{g}$ and $\mathscr{B}_{k}^{3}$ enjoy the property (cf. [7, Theorem 1.1]) that if $\tilde{g}=e^{2\tau}g$, then for all $\Psi\in C^{\infty}(\overline{X^{n+1}})$,
$\tilde{P}_{4}(\Psi)=e^{-\frac{n+5}{2}\tau}P_{4}^{g}(e^{\frac{n-3}{2}\tau}\Psi)$ (1.3)
and
~$\tilde{\mathscr{B}}_{k}^{3}(\Psi)=e^{-\frac{n+2k-3}{2}\tau}\mathscr{B}_{k}^{3}(e^{\frac{n-3}{2}\tau}\Psi).$ (1.4)
Readers are referred to Case [7] or authors [11] for explicit formulae of $P_{4}^{g}$ and $\mathscr{B}_{k}^{3}$ on manifolds with boundary. We are concerned with two conformally equivalent model spaces: $\mathbb{R}_{+}^{n+1}$ and $\mathbb{B}^{n+1}$. On the upper half-space (+$(\mathbb{R}^{n+1}_{+},\partial\mathbb{R}^{n+1}_{+},|\mathrm{d}X|^{2})$, the explicit formulae of conformal boundary operators are $\mathscr{B}^{3}_{0}(u)=u$ and
$\mathscr{B}^{3}_{1}(u)=$ $-\partial_{t}u;$
$\mathscr{B}^{3}_{2}(u)=$ $\partial_{t}^{2}u-\overline{\Delta}u;$
$\mathscr{B}^{3}_{3}(u)=$ $\partial_{t}\Delta u+2\overline{\Delta}\partial_{t}u$
for all $u\in C^{\infty}(\overline{\mathbb{R}_{+}^{n+1}})$. On the unit ball $(\mathbb{B}^{n+1},\mathbb{S}^{n},|\mathrm{d}\xi|^{2})$, these conformal boundary operators are $\mathscr{B}^{3}_{0}(U)=U$ and
$\mathscr{B}^{3}_{1}(U)=$ $\frac{\partial U}{\partial r}+\frac{n-3}{2}U,$
$\mathscr{B}^{3}_{2}(U)=$ $\frac{\partial^{2}U}{\partial r^{2}}-\Delta_{\mathbb{S}^{n}}U+(n-2)\frac{\partial U}{\partial r}+\frac{(n-3)(n-1)}{2}U,$
$\mathscr{B}^{3}_{3}(U)=$ $-\frac{\partial\Delta U}{\partial r}-2\Delta_{\mathbb{S}^{n}}\frac{\partial U}{\partial r}-\frac{n-3}{2}\frac{\partial^{2}U}{\partial r^{2}}-\frac{3n-5}{2}\Delta_{\mathbb{S}^{n}}U$
$+\frac{n-3}{2}\frac{\partial U}{\partial r}+\frac{(n^{2}-1)(n-3)}{4}U$
for all $U\in C^{\infty}(\overline{\mathbb{B}^{n+1}})$, where $r=|\xi|$ for $\xi\in\overline{\mathbb{B}^{n+1}}$; the corresponding boundary curvatures $(T_{i}^{3})_{g_{\mathbb{S}^{n}}}$ are constant and set $\mathbb{T}^{3}_{i}=\frac{n-3}{2}(T_{i}^{3})_{g_{\mathbb{S}^{n}}}$ for $1\leq i\leq 3$, explicitly,
$\mathbb{T}^{3}_{1}=\frac{n-3}{2},\qquad\mathbb{T}^{3}_{2}=\frac{(n-1)(n-3)}{2}\qquad\mathrm{and}\qquad\mathbb{T}^{3}_{3}=\frac{(n^{2}-1)(n-3)}{4}$
will be used throughout the paper.
Convention..
Let $\Delta$ and $\overline{\Delta}$ denote the Euclidean Laplacians on $\mathbb{R}^{n+1}$ and $\mathbb{R}^{n}$, respectively. Let $p_{k}^{\ast}=\frac{n+2k-3}{n-3}$ for $k\in\{0,1,2,3\}$, as $p_{k}^{\ast}+1$ is the critical Sobolev exponent associated to $\mathscr{B}^{3}_{k}$.
This motivates us to classify nonnegative smooth solutions to biharmonic boundary value problems on $\mathbb{R}_{+}^{n+1}$ associated with pairs of conformal boundary operators $(\mathscr{B}^{3}_{i},\mathscr{B}^{3}_{j})$ for $i<j\in\{1,2,3\}$: For $p_{i},p_{j}>0$, provided $u(x,0)$ has some suitable decay at infinity,
$\begin{cases}\displaystyle\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\quad~{}~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(u)=\mathbb{T}^{3}_{i}u^{p_{i}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{j}^{3}(u)=\mathbb{T}^{3}_{j}u^{p_{j}}&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (1.5)
The nonlinear interaction among $\mathscr{B}^{3}_{i}$ and $\mathscr{B}^{3}_{j}$ brings us new obstacles, compared with vanishing lower order part of derivatives in [30]. So, it has great difficulty in applying the approach in [30] to (1.5), as well as some peculiar phenomena occur.
An extremely powerful tool in our approach is so-called biharmonic Poisson kernel consisting of two kernel functions corresponding to a proper pair of conformal boundary operators $(\mathscr{B}_{i}^{3},\mathscr{B}_{j}^{3})$. In the following, we separate our discussion into two cases: $n\neq 3$ and $n=3$. In the latter case, it involves the critical dimension for both the Paneitz operator and $\mathscr{B}_{3}^{3}$.
Definition 1.1 (Poisson kernel for a biharmonic boundary value problem).
For $i,j\in\{0,1,2,3\}$ and $i<j$, a pair of kernel functions $P_{i}^{3}\oplus P_{j}^{3}(X)$ for $X\in\mathbb{R}_{+}^{n+1}$ is called a biharmonic Poisson kernel associated to a pair of conformal boundary operators $(\mathscr{B}_{i}^{3},\mathscr{B}_{j}^{3})$, if the following Poisson integral
$v(X)=$ $[P_{i}^{3}\ast f_{i}+P_{j}^{3}\ast f_{j}](X)$
$=$ $\int_{\partial\mathbb{R}_{+}^{n+1}}P_{i}^{3}(x-y,t)f_{i}(y)\mathrm{d}y+\int_{\partial\mathbb{R}_{+}^{n+1}}P_{j}^{3}(x-y,t)f_{j}(y)\mathrm{d}y$
makes sense, provided $f_{i},f_{j}\in C^{2}(\mathbb{R}^{n})$ have suitable decays at infinity, and gives a classical solution to
$\begin{cases}\displaystyle\Delta^{2}v=0&\mathrm{~{}~{}in~{}~{}}\quad~{}~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(v)=f_{i}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{j}^{3}(v)=f_{j}\qquad\,\,\,\,\,\,\,&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (1.6)
Obviously, the notion of biharmonic Poisson kernel could be naturally extended to $\mathbb{B}^{n+1}$ with boundary condition $(\mathscr{B}_{i}^{3},\mathscr{B}_{j}^{3})$.
Theorem 1.1.
For $n\neq 3$ and $X=(x,t)\in\mathbb{R}_{+}^{n+1}$, we let
$P_{0}^{3}(X)=$ $\frac{2(n+1)}{|\mathbb{S}^{n}|}\frac{t^{3}}{(t^{2}+|x|^{2})^{\frac{n+3}{2}}};$
$P_{1}^{3}(X)=$ $-\frac{2}{|\mathbb{S}^{n}|}\frac{t^{2}}{(t^{2}+|x|^{2})^{\frac{n+1}{2}}};$
$P_{2}^{3}(X)=$ $-\frac{1}{(n-1)|\mathbb{S}^{n}|}\frac{t}{(t^{2}+|x|^{2})^{\frac{n-1}{2}}};$
$P_{3}^{3}(X)=$ $\frac{1}{(n-1)(n-3)|\mathbb{S}^{n}|}\frac{1}{(t^{2}+|x|^{2})^{\frac{n-3}{2}}}.$
Suppose (i) $f_{0}\in C^{2}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$ ; (ii) $k\in\{1,2,3\}$, $f_{k}\in C^{2}(\mathbb{R}^{n})$ and $f_{k}=O(|x|^{-k-\delta_{k}})$ as $|x|\to\infty$ for some $\delta_{k}\in\mathbb{R}_{+}$. Then for $i,j\in\{0,1,2,3\}$ and $i+j\neq 3$, the explicit formula of Poisson kernel corresponding to biharmonic boundary value problem (1.6) is $P_{i}^{3}\oplus P_{j}^{3}$.
Remark 1.1.
It is straightforward to verify that other than two pairs $(\mathscr{B}_{0}^{3},\mathscr{B}_{3}^{3})$ and $(\mathscr{B}_{1}^{3},\mathscr{B}_{2}^{3})$, each $(\mathscr{B}_{i}^{3},\mathscr{B}_{j}^{3})$ satisfies ‘the complementing condition’ introduced by Agmon-Douglis-Nirenberg [1, Section 10], see also [15, Section 2.3].
For a smooth compact Riemannian manifold $(X^{n+1},g)$ with boundary $M^{n}$ and $n\geq 4$, under the assumption that a certain conformal invariant $\lambda_{1}(P_{4}^{g})>0$ (see [7]), the existence and uniqueness of smooth solutions to the following boundary value problem was studied in [7, Proposition 1.3]: For $\bar{f}_{0},\bar{f}_{1}\in C^{\infty}(M^{n})$,
$\begin{cases}\displaystyle P_{4}^{g}U=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad X^{n+1},\\ \displaystyle\mathscr{B}^{3}_{0}U=\bar{f}_{0}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,M^{n},\\ \displaystyle\mathscr{B}^{3}_{1}U=\bar{f}_{1}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,M^{n}.\end{cases}$
On the model space $(\mathbb{B}^{n+1},\mathbb{S}^{n},|\mathrm{d}\xi|^{2})$, for each proper pair of conformal boundary operators, such a biharmonic boundary value problem is expected to be better understood, serving as a first step to the prescribing boundary curvature problem on manifolds with boundary.
Theorem 1.2.
Suppose $n\neq 3$ and $\bar{f}_{i}\in C^{\infty}(\mathbb{S}^{n})$ for $i\in\{0,1,2,3\}$. For $\xi\in\mathbb{B}^{n+1}$ and $\eta\in\mathbb{S}^{n}$, we let
$\bar{P}_{0}^{3}(\xi,\eta)=$ $\frac{n+1}{4|\mathbb{S}^{n}|}\frac{\left(1-|\xi|^{2}\right)^{3}}{|\xi-\eta|^{n+3}};$
$\bar{P}_{1}^{3}(\xi,\eta)=$ $-\frac{1}{2|\mathbb{S}^{n}|}\frac{(1-|\xi|^{2})^{2}}{|\xi-\eta|^{n+1}};$
$\bar{P}_{2}^{3}(\xi,\eta)=$ $-\frac{1}{2(n-1)|\mathbb{S}^{n}|}\frac{1-|\xi|^{2}}{|\xi-\eta|^{n-1}};$
$\bar{P}_{3}^{3}(\xi,\eta)=$ $\frac{1}{(n-1)(n-3)|\mathbb{S}^{n}|}\frac{1}{|\xi-\eta|^{n-3}}.$
Then, for $i<j\in\{0,1,2,3\}$ we have
  • (i)
    For $i+j\neq 3$, the solution of the follow biharmonic boundary value problem
    $\begin{cases}\displaystyle\Delta^{2}U=0\qquad\quad&\mathrm{in}\qquad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}U=\bar{f}_{i}\qquad&\mathrm{on}\qquad\mathbb{S}^{n},\\ \displaystyle\mathscr{B}^{3}_{j}U=\bar{f}_{j}\qquad&\mathrm{on}\qquad\mathbb{S}^{n},\end{cases}$ (1.7)
    is unique and has the form of
    $U(\xi)=\int_{\mathbb{S}^{n}}\bar{P}_{i}^{3}(\xi,\eta)\bar{f}_{i}(\eta)\mathrm{d}V_{\mathbb{S}^{n}}(\eta)+\int_{\mathbb{S}^{n}}\bar{P}_{j}^{3}(\xi,\eta)\bar{f}_{j}(\eta)\mathrm{d}V_{\mathbb{S}^{n}}(\eta).$
  • (ii)
    For pairs $(0,3)$ and $(1,2)$, problem (1.7) has multiple smooth solutions.
Remark 1.2.
It is notable that restricting $\xi$ to $\mathbb{S}^{n}$, $2\bar{P}_{3}^{3}(\xi,\eta)$ coincides with the Green function of fractional GJMS operator $P_{3}$ on $\mathbb{S}^{n}$ (cf. the first author and Shi [10, Theorem 1(3)]). As an application, we can recover a special case in Case [7, Theorem 1.4] about the relationship between extrinsic and intrinsic GJMS operators, see Proposition 3.1 and its biharmonic Poisson kernel proof in Appendix A. Moreover, the non-uniqueness in Theorem 1.2(ii) actually arises from the invertibility of fractional GJMS operators $P_{3}$ and $P_{1}$, see [7, 10] for the definition.
We are ready to introduce a biharmonic Green function on $\mathbb{B}^{n+1}$ with each of four proper pairs of conformal boundary operators. Let
$\Gamma(\xi-\eta)=\Gamma(|\xi-\eta|)=\frac{1}{2(n-1)(n-3)|\mathbb{S}^{n}|}|\xi-\eta|^{3-n}\qquad\mathrm{~{}~{}when~{}~{}}n\neq 3$
and
$\Gamma(\xi-\eta)=\Gamma(|\xi-\eta|)=-\frac{1}{4|\mathbb{S}^{3}|}\log|\xi-\eta|\qquad\qquad\mathrm{~{}~{}when~{}~{}}n=3$
denote the fundamental solution of $\Delta^{2}$ in $\mathbb{R}^{n+1}$. We consider an nonhomogeneous biharmonic boundary value problem with vanishing boundary conditions
$\begin{cases}\Delta^{2}U=\bar{f}\qquad&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{B}^{n+1},\\ \mathscr{B}_{i}^{3}U=0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{S}^{n},\\ \mathscr{B}_{j}^{3}U=0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{S}^{n}.\end{cases}$ (1.8)
Remark 1.3.
A clever observation in (4.1) provides a traditional way in understanding biharmonic Poisson kernel: Conformal boundary operators $\mathscr{B}_{i}^{3}$ indeed act as a bridge between the biharmonic Poisson kernel and fundamental solution.
Definition 1.2 (biharmonic Green function).
Let $\bar{f}\in C^{\infty}(\overline{\mathbb{B}^{n+1}})$ and $i<j\in\{0,1,2,3\}$ with $i+j\neq 3$. A function $\bar{G}^{(i,j)}:(\xi,\eta)\in\overline{\mathbb{B}^{n+1}}\times\overline{\mathbb{B}^{n+1}}\mapsto\mathbb{R}\cup\{\infty\}$ for (1.8) is called a Green function, if $\bar{G}^{(i,j)}$ satisfies
  • $\xi\mapsto\bar{G}^{(i,j)}(\xi,\eta)-\Gamma(\xi-\eta)\in C^{\infty}(\overline{\mathbb{B}^{n+1}})$ for all $\eta\in\mathbb{B}^{n+1}$;
  • $\Delta_{\xi}^{2}\left(\bar{G}^{(i,j)}(\xi,\eta)-\Gamma(\xi-\eta)\right)=0$ for all $(\xi,\eta)\in\mathbb{B}^{n+1}\times\mathbb{B}^{n+1}$;
  • $\mathscr{B}_{i}^{3}\bar{G}^{(i,j)}=0,\mathscr{B}_{j}^{3}\bar{G}^{(i,j)}=0$ on $\mathbb{S}^{n}$.
Formally, the above Green function enables us to express a unique solution to (1.8) as
$U(\xi)=\bar{G}^{(i,j)}\ast\bar{f}(\xi)=\int_{\mathbb{B}^{n+1}}\bar{G}^{(i,j)}(\xi,\eta)\bar{f}(\eta)\mathrm{d}\eta.$
We first notice that $\bar{G}^{(0,1)}$ coincides with the Boggio formula (see, e.g., Boggio [2] or Gazzola-Grunau-Sweers [15, Lemma 2.27]), next combine this with biharmonic Poisson kernel to derive explicit formulae of Green functions.
Theorem 1.3.
For $n\neq 3$, the representation formula of Green function $\bar{G}^{(i,j)}$ to (1.8) is
$\bar{G}^{(0,1)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\frac{1}{2}\Gamma(|\xi|\cdot|\xi^{\ast}-\eta|)\left[(n-3)\frac{|\xi-\eta|^{2}}{|\xi|^{2}\cdot|\xi^{\ast}-\eta|^{2}}-(n-1)\right];$
$\bar{G}^{(2,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)-\frac{1}{2}\Gamma(|\xi|\cdot|\xi^{\ast}-\eta|)\left[(n-3)\frac{|\xi-\eta|^{2}}{|\xi|^{2}\cdot|\xi^{\ast}-\eta|^{2}}-(n-1)\right];$
$\bar{G}^{(0,2)}(\xi,\eta)=$ $\Gamma(\xi-\eta)-\Gamma(|\xi|\cdot|\xi^{*}-\eta|);$
$\bar{G}^{(1,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\Gamma(|\xi|\cdot|\xi^{*}-\eta|);$
where $\xi^{\ast}=|\xi|^{-2}\xi$ for $0\neq\xi\in\overline{\mathbb{B}^{n+1}}$. Moreover, there holds
$0\leq\bar{G}^{(0,1)}(\xi,\eta)\leq\bar{G}^{(0,2)}(\xi,\eta)\leq\bar{G}^{(1,3)}(\xi,\eta)\leq\bar{G}^{(2,3)}(\xi,\eta),\quad\forall~{}\xi\neq\eta\in\overline{\mathbb{B}^{n+1}},$
with equality if and only if either $\xi\in\mathbb{S}^{n}$ or $\eta\in\mathbb{S}^{n}$.
In the critical dimension that $n=3$, we have
Theorem 1.4.
Suppose $i,j\in\{0,1,2,3\}$ and $i<j$. The following are explicit formulae of biharmonic Poisson kernels on $\mathbb{R}_{+}^{4}$ and $\mathbb{B}^{4}$.
  • (1)
    Under the same assumptions as in Theorem 1.1 when $n=3$, then the same formulae of $P_{i}^{3}$ are true except for
    $P_{3}^{3}(X)=-\frac{1}{4|\mathbb{S}^{3}|}\log(t^{2}+|x|^{2})+C,\qquad C\in\mathbb{R}$
    for $X\in\mathbb{R}_{+}^{4}$, and the biharmonic Poisson kernel for (1.6) is $P_{i}^{3}\oplus P_{j}^{3}$ when $i+j\neq 3$.
  • (2)
    Under the same assumptions as in Theorem 1.2 when $n=3$, then the same formulae of $\bar{P}_{i}^{3}$ are true except for
    $\bar{P}_{3}^{3}(\xi,\eta)=-\frac{1}{2|\mathbb{S}^{3}|}\log|\xi-\eta|+C,\qquad C\in\mathbb{R}$
    for $\xi\in\mathbb{B}^{4},\eta\in\mathbb{S}^{3}$. Moreover, we have
    (i) $i=1,2$ and $j=3$, the solution of (1.7) is
    $U(\xi)-\frac{1}{|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}U\mathrm{d}V_{{\mathbb{S}^{3}}}$
    $=$ $\int_{\mathbb{S}^{3}}\bar{P}^{3}_{i}(\xi-\eta)\bar{f}_{i}(\eta)\mathrm{d}V_{{\mathbb{S}^{3}}}(\eta)-\frac{1}{2|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}\log|\xi-\eta|\bar{f}_{3}(\eta)\mathrm{d}V_{\mathbb{S}^{3}}(\eta),\quad\forall~{}\xi\in\overline{\mathbb{B}^{4}}.$
    (ii) $(i,j)=(0,1)\mathrm{~{}or~{}}(0,2)$, the unique solution of (1.7) is
    $U(\xi)=\int_{\mathbb{S}^{3}}\bar{P}^{3}_{i}(\xi-\eta)\bar{f}_{i}(\eta)+\bar{P}^{3}_{j}(\xi-\eta)\bar{f}_{j}(\eta)\mathrm{d}V_{{\mathbb{S}^{3}}}(\eta),\quad\forall~{}\xi\in\overline{\mathbb{B}^{4}}.$
Theorem 1.5.
The representation formula of Green function $\bar{G}^{(i,j)}$ to (1.8) is
$\bar{G}^{(0,1)}(\xi,\eta)=$ $\Gamma(\xi-\eta)-\Gamma(|\xi|\cdot|\xi^{*}-\eta|)+\frac{1}{8|\mathbb{S}^{3}|}\left(\frac{|\xi-\eta|^{2}}{|\xi|^{2}|\xi^{*}-\eta|^{2}}-1\right);$
$\bar{G}^{(2,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\Gamma(|\xi|\cdot|\xi^{*}-\eta|)-\frac{1}{8|\mathbb{S}^{3}|}\left(\frac{|\xi-\eta|^{2}}{|\xi|^{2}|\xi^{*}-\eta|^{2}}-1\right)+C,\quad C\in\mathbb{R};$
$\bar{G}^{(0,2)}(\xi,\eta)=$ $\Gamma(\xi-\eta)-\Gamma(|\xi|\cdot|\xi^{*}-\eta|);$
$\bar{G}^{(1,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\Gamma(|\xi|\cdot|\xi^{*}-\eta|)+C,\quad C\in\mathbb{R};$
for $\xi,\eta\in\overline{\mathbb{B}^{4}}$, where $\xi^{\ast}=|\xi|^{-2}\xi$ for $0\neq\xi\in\overline{\mathbb{B}^{4}}$.
We shall assemble the biharmonic Poisson kernel part in Section 3 and the biharmonic Green function part in Section 4.
We next turn to prove classification theorems of nonnegative solutions to (1.5) in $\mathbb{R}_{+}^{n+1}$. Before presenting the proof of our results, in Section 2 we first need to study the nontrivial solutions of an associated homogeneous problem to (1.6) under suitable assumptions on the decay rate at infinity of the initial data, and also describe asymptotic behavior of two typical singular integrals originating from the biharmonic Poisson integrals for $\mathbb{R}_{+}^{n+1}$.
For simplicity, we introduce a ‘geometric bubble function’ associated to above biharmonic boundary value problem by
$U_{\varepsilon,x_{0}}(x,t)=\left(\frac{2\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}},\qquad(x,t)\in\mathbb{R}_{+}^{n+1},$
and its boundary function by $u_{\varepsilon,x_{0}}(x)=U_{\varepsilon,x_{0}}(x,0)$, where $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$.
For the pair $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{3})$, under appropriate decay rate at infinity of initial data, the intrinsic nature of $\mathscr{B}^{3}_{3}$ boundary condition enables us to determine the explicit representation of $u$ on $\partial\mathbb{R}_{+}^{n+1}$, see Lemma 5.1. Using basic properties of $U_{\varepsilon,x_{0}}$ and biharmonic Poisson kernel, we establish that every nonnegative smooth solution is a sum of a bubble function and a polynomial. When $p_{3}=p_{3}^{\ast}$, a striking feature is that the bubble function (in general, without geometric meaning) does not vanish even if $p_{1}\neq p_{1}^{\ast}$, and is exact $U_{\varepsilon,x_{0}}$ when $p_{1}=p_{1}^{\ast}$; whereas for $p_{1}<p_{1}^{\ast}$, the bubble function vanishes.
Theorem 1.6.
For $n\geq 4$ and $p_{1},p_{3}>0$, let $u$ be a nonnegative smooth solution of
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}_{1}^{3}(u)=\mathbb{T}^{3}_{1}u^{p_{1}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}(u)=\mathbb{T}^{3}_{3}u^{p_{3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
Assume $u(x,0)=O(|x|^{-c})$ as $|x|\to\infty$ for some constant $c>\max\{3/p_{3},1/p_{1}\}$. Then we have
  • (1)
    If $p_{3}=p_{3}^{\ast}=\frac{n+3}{n-3}$ and $p_{1}>\frac{1}{n-3}$, then there exist some $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$ and nonnegative constant $c_{2}$ such that
    $u(x,t)=M_{1}(x,t)+c_{2}t^{2},$
    where
    $M_{1}(x,t)=\left[\mathbb{T}_{1}^{3}P_{1}^{3}\ast u_{\varepsilon,x_{0}}^{p_{1}}+\mathbb{T}_{3}^{3}P_{3}^{3}\ast u_{\varepsilon,x_{0}}^{p_{3}^{\ast}}\right](x,t).$
    In particular, if $p_{1}=p_{1}^{\ast}=\frac{n-1}{n-3}$, then
    $u(x,t)=U_{\varepsilon,x_{0}}(x,t)+c_{2}t^{2}.$
  • (2)
    If $0<p_{3}<p_{3}^{\ast}$, then $u(x,t)=c_{2}t^{2}$ for every nonnegative constant $c_{2}$.
For a nonlinear boundary conditions involving the exceptional pair $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{2})$, there still exists a surprising classification theorem for problem (1.5) relying not on $p_{1},p_{2}$ but on $p_{2}/p_{1}$.
Theorem 1.7.
For $n\geq 4$ and $p_{1},p_{2}>0$, let $u$ be a nonnegative smooth solution of
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{1}(u)=\mathbb{T}^{3}_{1}u^{p_{1}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}^{3}_{2}(u)=\mathbb{T}^{3}_{2}u^{p_{2}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
Assume $u(x,0)=O(|x|^{-c})$ as $|x|\to\infty$ for some constant $c>2/p_{2}$. Then we have
  • (1)
    If $\frac{p_{2}}{p_{1}}=\frac{n+1}{n-1}$ and $p_{1}>0$, then there exist some $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$ and nonnegative constant $c_{3}$ such that
    $u(x,t)=M_{2}(x,t)+c_{3}t^{3},$
    where
    $M_{2}(x,t)=\left[P_{0}^{3}\ast u_{\varepsilon,x_{0}}^{\frac{p_{1}^{\ast}}{p_{1}}}+\mathbb{T}_{2}^{3}P_{2}^{3}\ast u_{\varepsilon,x_{0}}^{p_{2}^{\ast}}\right](x,t).$
    In particular, $p_{1}=p_{1}^{\ast}=\frac{n-1}{n-3}$, then
    $u(x,t)=U_{\varepsilon,x_{0}}(x,t)+c_{3}t^{3}.$
  • (2)
    If $\frac{p_{2}}{p_{1}}<\frac{n+1}{n-1}$ , then $u(x,t)=c_{3}t^{3}$ for every nonnegative constant $c_{3}$.
The pair $(\mathscr{B}^{3}_{2},\mathscr{B}^{3}_{3})$ becomes more challenging and $p_{3}$ is still the dominant exponent. Especially, when $p_{i}=p_{i}^{\ast}$ for $i=1,2,3$, the bubble function happens to be the same $U_{\varepsilon,x_{0}}$; meanwhile, the polynomial part varies according to distinct circumstances.
Theorem 1.8.
For $n\geq 4$ and $p_{2},p_{3}>0$, let $u$ be a nonnegative smooth solution of
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{2}(u)=\mathbb{T}^{3}_{2}u^{p_{2}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}(u)=\mathbb{T}^{3}_{3}u^{p_{3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
Assume $u(x,0)=O(|x|^{-c})$ as $|X|\to\infty$ for some constant $c>\max\{2/p_{2},3/p_{3}\}$. Then we have
  • (1)
    If $p_{3}=p_{3}^{\ast}=\frac{n+3}{n-3}$ and $p_{2}>\frac{2}{n-3}$, then there exist some $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$ and nonnegative constant $c_{1}$ such that
    $u(x,t)=M_{3}(x,t)+c_{1}t,$
    where
    $M_{3}(x,t)=\left[P_{0}^{3}\ast u_{\varepsilon,x_{0}}+\mathbb{T}_{2}^{3}P_{2}^{3}\ast u_{\varepsilon,x_{0}}^{p_{2}}\right](x,t).$
    In particular, $p_{2}=p_{2}^{\ast}=\frac{n+1}{n-3}$, then
    $u(x,t)=U_{\varepsilon,x_{0}}(x,t)+c_{1}t.$
  • (2)
    If $0<p_{3}<\frac{n+3}{n-3}$, then $u(x,t)=c_{1}t$ for every nonnegative constant $c_{1}$.
These classification results demonstrate that the dominant impact of $\mathscr{B}^{3}_{i}$ is different, accordingly, $\mathscr{B}^{3}_{1}\preceq\mathscr{B}^{3}_{2}\preceq\mathscr{B}^{3}_{3}$. In Section 5 we shall present the proof of Theorems 1.6-1.8.
The above classification results in $\mathbb{R}_{+}^{n+1}$ enable us to classify nonnegative solutions to biharmonic boundary value problems on $\mathbb{B}^{n+1}$ with one boundary point singularity.
Theorem 1.9.
Suppose $n\geq 4$ and $U\in C^{4}(\overline{\mathbb{B}^{n+1}}\backslash\{-\mathbf{e}_{n+1}\})$ is a nonnegative solution to
$\begin{cases}\displaystyle\Delta^{2}U=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(U)=\mathbb{T}^{3}_{i}U^{\frac{n+2i-3}{n-3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\mathbb{S}^{n}\backslash\{-\mathbf{e}_{n+1}\},\\ \displaystyle\mathscr{B}^{3}_{j}(U)=\mathbb{T}^{3}_{j}U^{\frac{n+2j-3}{n-3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\mathbb{S}^{n}\backslash\{-\mathbf{e}_{n+1}\},\end{cases}$ (1.9)
where $i,j\in\{1,2,3\}$ with $i<j$. Assume that there exist $\alpha\in(0,\frac{n(n-3)}{n+3})$ and $C\in\mathbb{R}_{+}$ such that
$\sup_{\xi\in\mathbb{S}^{n}\backslash\{-\mathbf{e}_{n+1}\}}U(\xi)|\xi+\mathbf{e}_{n+1}|^{\alpha}\leq C.$ (1.10)
Then there exist some $\xi_{0}\in\mathbb{B}^{n+1}$ and nonnegative constant $\bar{c}$ such that
$U(\xi)=\left(\frac{1-|\xi_{0}|^{2}}{|\xi|^{2}|\xi_{0}|^{2}-2\xi_{0}\cdot\xi+1}\right)^{\frac{n-3}{2}}+\bar{c}|\xi+\mathbf{e}_{n+1}|^{3-n}\left(\frac{1-|\xi|^{2}}{|\xi+\mathbf{e}_{n+1}|}\right)^{6-i-j}.$
Remark 1.4.
If $i=1,j=2$, then the assumption of $\alpha$ can be relaxed to $\alpha\in(0,\frac{(n-1)(n-3)}{n+1})$, which ensures that $n-3-\alpha>\frac{2}{p_{2}^{\ast}}$ by virtue of Theorem 1.7.
A direct consequence of Theorem 1.9 is to prove the uniqueness of positive smooth solutions to a certain type of geometric partial differential equations on $\mathbb{B}^{n+1}$.
Corollary 1.1.
Suppose $n\geq 4$ and $U$ is a positive smooth solution in $\overline{\mathbb{B}^{n+1}}$ to
$\begin{cases}\displaystyle\Delta^{2}U=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(U)=\mathbb{T}^{3}_{i}U^{\frac{n+2i-3}{n-3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\mathbb{S}^{n},\\ \displaystyle\mathscr{B}^{3}_{j}(U)=\mathbb{T}^{3}_{j}U^{\frac{n+2j-3}{n-3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\mathbb{S}^{n},\end{cases}$
where $i<j\in\{1,2,3\}$. Then
$U(\xi)=\left(\frac{1-|\xi_{0}|^{2}}{|\xi|^{2}|\xi_{0}|^{2}-2\xi_{0}\cdot\xi+1}\right)^{\frac{n-3}{2}}$
for some $\xi_{0}\in\mathbb{B}^{n+1}$.
We now pause to comment on the decay assumptions at infinity of initial data that $u(x,0)=O(|x|^{-c})$ as $|x|\to\infty$ in Theorems 1.6-1.8. For critical exponents $p_{i}^{\ast}$, such assumptions are to ensure that involved Poisson integrals are well-defined. As a result, this deduces that $c>\frac{3(n-3)}{n+3}$ in Theorems 1.6 and 1.8; $c>\frac{2(n-3)}{n+1}$ in Theorem 1.7. As will show in the proof of Theorem 1.9, we know $c=n-3-\alpha$, then the assumption (1.10) leads to $\alpha<\frac{n(n-3)}{n+1}$ and $\alpha<\frac{(n-1)(n-3)}{n+1}$, respectively. Another interpretation of assumption of $\alpha$ is as follows: If $\mathscr{B}^{3}_{3}(U)=\frac{n-3}{2}\tilde{T}^{3}_{3}U^{\frac{n+3}{n-3}}\in L^{1}(\mathbb{S}^{n})$ for a smooth conformal metric $\tilde{g}=U^{4/(n-3)}|\mathrm{d}\xi|^{2}$, then near the singular point $-\mathbf{e}_{n+1}$,
$U^{\frac{n+3}{n-3}}(\xi)\lesssim|\xi+\mathbf{e}_{n+1}|^{-\frac{\alpha(n+3)}{n-3}}\in L^{1}(\mathbb{S}^{n}),$
which requires $0<\alpha<\frac{n(n-3)}{n+3}$. Intuitively, $\mathscr{B}^{3}_{3}(U)\in L^{1}(\mathbb{S}^{n})$ means $U\in W^{3,1}(\mathbb{S}^{n})\hookrightarrow W^{2,\frac{n}{n-1}}(\mathbb{S}^{n})$. This roughly gives $\mathscr{B}^{3}_{2}(U)\in L^{\frac{n}{n-1}}(\mathbb{S}^{n})$, and also
$\left(U^{\frac{n+1}{n-3}}(\xi)\right)^{\frac{n}{n-1}}\lesssim|\xi+\mathbf{e}_{n+1}|^{\frac{\alpha n(n+1)}{(n-1)(n-3)}}\in L^{1}(\mathbb{S}^{n}),$
which requires $\alpha<\frac{(n-1)(n-3)}{n+1}$.
In Section 6, we give the proof of Theorem 1.9, and as applications we also establish two sharp geometric inequalities over some $Q$-curvature flat conformal class of $(\mathbb{B}^{n+1},|\mathrm{d}\xi|^{2})$ for $n\geq 4$, including a sharp inequality of isoperimetric type.
Acknowledgments: The authors thank Professor Qianqiao Guo for bringing [17] into their sight.
2 Preliminaries
We give asymptotic behavior of two prototypes of singular integrals, coming from biharmonic Poisson integrals on $\mathbb{R}_{+}^{n+1}$.
Lemma 2.1.
Suppose $f\in L_{\mathrm{loc}}^{1}(\mathbb{R}^{n})$ with $f(x)=O(|x|^{-a})$ as $|x|\to\infty$ for some $a\in\mathbb{R}_{+}$. Assume $20<\alpha<n/$ 2 and $a+2\alpha>n$. We define
$v(x,t)=\int_{\mathbb{R}^{n}}\frac{f(y)}{(t^{2}+|x-y|^{2})^{\alpha}}\mathrm{d}y.$
Then there holds
$\left|v(X)\right|\lesssim$$\begin{cases}|X|^{n-2\alpha-\min\{a,n\}}\quad&\mathrm{if~{}~{}}a\neq n\\ |X|^{-2\alpha}\log|X|&\mathrm{if~{}~{}}a=n\end{cases},\qquad\quad\mathrm{~{}~{}as~{}~{}}\quad|X|\to\infty.$
Proof..
We decompose $\mathbb{R}^{n}=A_{1}\cup A_{2}\cup A_{3}$ with
$A_{1}=$$\left\{y||y|<|x|/2\right\},\qquad\qquad\qquad\qquad A_{2}=\left\{y||x-y|<|x|/2\right\},$
$A_{3}=$$\left\{y||y|\geq|x|/2,|x-y|\geq|x|/2\right\}.$
Case 1. $|X|\gg 1$ and $t>\frac{|x|}{2}$.
There hold $t\gg 1$ and $|X|\sim t$ due to
$1\leq\frac{|X|^{2}}{t^{2}}=\frac{t^{2}+|x|^{2}}{t^{2}}\leq 5$
In $A_{1}$, we have $|x-y|\sim|x|$ and $t^{2}+|x-y|^{2}\sim t^{2}$, hence
$|v(x,t)|\lesssim t^{-2\alpha}\int_{A_{1}}|f(y)|\mathrm{d}y\lesssim$$\begin{cases}t^{-2\alpha}|x|^{n-\min\{a,n\}}&\mathrm{if~{}~{}}a\neq n\\ t^{-2\alpha}\log(|x|+1)&\mathrm{if~{}~{}}a=n\end{cases}$
$\lesssim$ $\begin{cases}|X|^{n-2\alpha-\min\{a,n\}}&\mathrm{if~{}~{}}a\neq n\\ |X|^{-2\alpha}\log|X|&\mathrm{if~{}~{}}a=n\end{cases}.$
In $A_{2}$, we have $|x-y|<|x|/2<t$, $|y|\sim|x|$ and $t^{2}+|x-y|^{2}\sim t^{2}$, then
$|v(x,t)|\lesssim t^{-2\alpha}\int_{A_{2}}|f(y)|\mathrm{d}y\lesssim$$\begin{cases}|X|^{n-2\alpha-\min\{a,n\}}&\mathrm{if~{}~{}}a\neq n\\ |X|^{-2\alpha}\log|X|&\mathrm{if~{}~{}}a=n\end{cases}.$
In $A_{3}$, we have $|x-y|\leq|y|+|x|\leq 3|y|$ and $|y|\leq|x-y|+|x|\leq 3|x-y|$, whence $|x-y|\sim|y|$ We further decompose $A_{3}=A_{31}\cup A_{32}$ with
$A_{31}=A_{3}\cap\{y||y|>t\},\qquad\qquad A_{32}=A_{3}\cap\{y||y|\leq t\}.$
In $A_{31}$, we have $t^{2}+|x-y|^{2}\sim|y|^{2}$ and
[…]
In $A_{32}$, we have $t^{2}+|x-y|^{2}\sim t^{2}$ and
$|v(x,t)|\lesssim t^{-2\alpha}\int_{A_{32}}|f(y)|\mathrm{d}y\lesssim$$\begin{cases}t^{n-2\alpha-\min\{a,n\}}&\mathrm{if~{}~{}}a\neq n\\ t^{-2\alpha}\log t&\mathrm{if~{}~{}}a=n\end{cases}$
$\lesssim$ $\begin{cases}|X|^{n-2\alpha-\min\{a,n\}}&\mathrm{if~{}~{}}a\neq n\\ |X|^{-2\alpha}\log|X|&\mathrm{if~{}~{}}a=n\end{cases}.$
Hence we combine the estimates above to obtain
$\left|v(x,t)\right|\lesssim\begin{cases}|X|^{n-2\alpha-\min\{a,n\}}\qquad&\mathrm{if~{}~{}}a\neq n\\ |X|^{-2\alpha}\log|X|&\mathrm{if~{}~{}}a=n\end{cases}.$
Case 2. $|X|\gg 1$ and $t\leq\frac{|x|}{2}$.
We have $|x|\gg 1$ and $|X|\sim|x|$ due to
$|x|^{2}\leq|X|^{2}=t^{2}+|x|^{2}\leq\frac{5}{4}|x|^{2}.$
In $A_{1}$, we have $\frac{|x|}{2}\leq|x-y|\leq\frac{3|x|}{2}$ and $t^{2}+|x-y|^{2}\sim|x|^{2}$, then,
$|v(x,t)|\lesssim|x|^{-2\alpha}\int_{A_{1}}|f(y)|\mathrm{d}y\lesssim$$\begin{cases}|x|^{-2\alpha}|x|^{n-\min\{a,n\}}&\mathrm{if~{}~{}}a\neq n\\ |x|^{-2\alpha}\log(|x|+1)&\mathrm{if~{}~{}}a=n\end{cases}$
$\lesssim$ $\begin{cases}|X|^{n-2\alpha-\min\{a,n\}}&\mathrm{if~{}~{}}a\neq n\\ |X|^{-2\alpha}\log|X|&\mathrm{if~{}~{}}a=n\end{cases}.$
In $A_{2}$, we have $|x|/2<|y|<3|x|/2$. Further decompose $A_{2}=A_{21}\cup A_{22}$ with
$A_{21}=A_{2}\cap\{y||x-y|>t\}\qquad\mathrm{and}\qquad A_{22}=A_{2}\cap\{y||x-y|\leq t\}.$
For $A_{21}$, we have $t^{2}+|x-y|^{2}\sim|x-y|^{2}$ and
$|v(x,t)|\lesssim\int_{A_{21}}\frac{|f(y)|}{|x-y|^{2\alpha}}\mathrm{d}y\lesssim|x|^{-a}\int_{|x-y|<|x|/2}\frac{1}{|x-y|^{2\alpha}}\mathrm{d}y\lesssim|X|^{n-2\alpha-a}.$
For $A_{22}$, $t^{2}+|x-y|^{2}\sim t^{2}$, we know
$|v(x,t)|\lesssim t^{-2\alpha}\int_{A_{22}}|f(y)|\mathrm{d}y\lesssim t^{-2\alpha}|x|^{-a}\int_{|x-y|\leq t}\mathrm{d}y\lesssim|X|^{n-2\alpha-a}.$
For $A_{3}$, we know $|y|/3<|x-y|<3|y|$, $|y|>|x|/2>t$, whence, $t^{2}+|x-y|^{2}\sim|y|^{2}$. Then,
$|v(x,t)|\lesssim\int_{A_{3}}\frac{|f(y)|}{|y|^{2\alpha}}\mathrm{d}y\lesssim\int_{|y|>|x|/2}|y|^{-(2\alpha+a)}\mathrm{d}y\lesssim|X|^{n-2\alpha-a}.$
Hence, putting these facts above together we have
$\left|v(x,t)\right|\lesssim\begin{cases}|X|^{n-2\alpha-\min\{a,n\}}\qquad&\mathrm{if~{}~{}}a\neq n\\ |X|^{-2\alpha}\log|X|&\mathrm{if~{}~{}}a=n\end{cases}.$
$\Box$
Lemma 2.2.
Suppose $f\in L_{\mathrm{loc}}^{1}(\mathbb{R}^{n})$ with the decay that $f(x)=O(|x|^{-b})$ as $|x|\to\infty$ for some positive constant $b$. We define
$w(x,t)=\int_{\mathbb{R}^{n}}\frac{t^{\beta}}{(t^{2}+|x-y|^{2})^{\frac{n+\beta}{2}}}f(y)\mathrm{d}y,\quad\beta\in\mathbb{R}_{+}.$
Then there holds
$\left|w(x,t)\right|\lesssim$$\begin{cases}|X|^{-\min\{b,n\}}\qquad&\mathrm{if~{}~{}}b\neq n\\ |X|^{-n}\log|X|&\mathrm{if~{}~{}}b=n\end{cases},\qquad\quad\mathrm{~{}~{}as~{}~{}}\quad|X|\to\infty.$
Since the same trick works for the above second type singular integral, we omit the proof.
For $r\in\mathbb{R}_{+}$, let $B_{r}^{+}=B_{r}(0)\cap\mathbb{R}^{n+1}_{+}$ and $\partial^{+}B_{r}^{+}=\partial B_{r}^{+}\cap\mathbb{R}^{n+1}_{+}$. Let $P_{k}(x)$ denote a polynomial with degree $k$ in the variable $x$, where $k\in\mathbb{N}$. Given $u\in C^{\infty}(\overline{\mathbb{R}^{n+1}_{+}})$, we introduce a Kelvin transform for $\Delta^{2}$ by
$u^{\ast}(X)=|X|^{3-n}u\left(\frac{X}{|X|^{2}}\right),\qquad X\in\overline{\mathbb{R}_{+}^{n+1}}\backslash\{0\}.$
For our purpose, we need to refine Sun-Xiong’s results [30, Proposition 3.1 and Theorem 3.2] on homogeneous biharmonic boundary value problems.
Lemma 2.3.
Let $u\in C^{\infty}(\overline{\mathbb{R}^{n+1}_{+}})$ be a solution of
$\begin{cases}\displaystyle\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{R}_{+}^{n+1},\\ \displaystyle\partial_{t}u(x,0)=0&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta u(x,0)=0\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (2.1)
Suppose (i) $\lim_{|x|\to\infty}u(x,0)=0$; (ii) there exist some $\delta,C\in\mathbb{R}_{+}$ such that
$u(X)\geq-C|X|^{-\delta}\qquad\mathrm{as}\quad|X|\to\infty.$ (2.2)
Then, $u(x,t)=c_{2}t^{2}$ for some nonnegative constant $c_{2}$.
Proof..
It is notable that the nonnegativity of solutions is removed compared with that of Sun-Xiong [30].
The proof consists of three steps.
Step 1. $u^{\ast}\in L^{1}(B_{1}^{+})$.
First, rewrite the boundary conditions as $\mathscr{B}_{1}^{3}u=0,\mathscr{B}_{3}^{3}u=0$, then it follows from conformal covariance that $u^{\ast}$ satisfies the biharmonic boundary value problem (2.1) in $\overline{\mathbb{R}_{+}^{n+1}}\backslash\{0\}$.
Define
$\eta_{\varepsilon}(t)=\begin{cases}\displaystyle\frac{1}{24}(t-\varepsilon)^{4}\qquad&\mathrm{for}\qquad t\geq\varepsilon,\\ \displaystyle 0\qquad&\mathrm{for}\qquad 0\leq t<\varepsilon.\end{cases}$
Notice that $\eta_{\varepsilon}\in C^{3,1}(\mathbb{R}^{n+1}_{+})$ and $u^{*}$ is biharmonic on +$\overline{\mathbb{R}^{n+1}_{+}}\backslash\{0\}$, then integrating by parts gives
$\int_{B_{1}^{+}}u^{*}(X)\Delta^{2}\eta_{\varepsilon}\mathrm{d}X$
$=$ $\int_{\partial^{+}B_{1}^{+}}\left(\Delta u^{*}(X)\frac{\partial}{\partial r}\eta_{\varepsilon}-\frac{\partial}{\partial r}\Delta u^{*}(X)\eta_{\varepsilon}+u^{*}(X)\frac{\partial}{\partial r}\Delta\eta_{\varepsilon}-\frac{\partial}{\partial r}u^{*}(X)\Delta\eta_{\varepsilon}\right)\mathrm{d}\sigma.$
Letting $\varepsilon\to 0$ we obtain
$\int_{B_{1}^{+}}u^{*}(X)\mathrm{d}X\leq C.$ (2.3)
Write $u^{*}(X)=u_{+}^{*}(X)-u_{-}^{*}(X)$. It follows from the assumption (2.2) that $u_{-}^{*}(X)\leq C|X|^{\delta+3-n}$. Using (2.3) we obtain
$\int_{B_{1}^{+}}u_{+}^{*}(X)\mathrm{d}X\leq C+\int_{B_{1}^{+}}u_{-}^{*}(X)\mathrm{d}X\leq C$
and thus $u^{\ast}\in L^{1}(B_{1}^{+})$.
Step 2. $u(X)=o(|X|^{4})$ as $|X|\to\infty$.
For each $r\in\mathbb{R}_{+}$, we set $v_{r}(X)=u^{*}(rX)$, which is biharmonic in +$\overline{\mathbb{R}^{n+1}_{+}}\backslash\{0\}$. Now we extend $v_{r}$ by an even reflection to $\mathbb{R}^{n+1}\backslash\{0\}$ and apply the local estimates for biharmonic functions in [27, Proposition 4] to obtain
$\|v_{r}\|_{L^{\infty}(B^{+}_{3/4}\backslash B^{+}_{1/2})}\lesssim\|v_{r}\|_{L^{1}(B^{+}_{5/4}\backslash B^{+}_{1/4})}\lesssim r^{-(n+1)}\|u^{*}\|_{L^{1}(B^{+}_{5r/4}\backslash B^{+}_{r/4})}.$
Since $u^{\ast}\in L^{1}(B_{1}^{+})$, we may take $r$ sufficiently small such that
$\|u^{*}\|_{L^{\infty}(B^{+}_{3r/4}\backslash B^{+}_{r/2})}=o(r^{-(n+1)}).$
This in turn implies
$u(X)=o(|X|^{4})\qquad\mathrm{~{}~{}as~{}~{}}|X|\to\infty.$
Step 3. Completion of the proof.
Extend $u$ by an even reflection
$\tilde{u}(x,t)=\begin{cases}\displaystyle u(x,t)\qquad&\mathrm{for}\qquad t\geq 0,\\ \displaystyle u(x,-t)\qquad&\mathrm{for}\qquad t<0.\end{cases}$
Notice that $\partial_{t}\tilde{u}(x,0)=0$ and
$\partial_{t}^{3}u(x,0)=\partial_{t}\Delta u(x,0)-\partial_{t}\overline{\Delta}u(x,0)=0,$
so $\tilde{u}\in C^{4}(\mathbb{R}^{n+1})$ and is biharmonic in $\mathbb{R}^{n+1}$. By [27, Proposition 4] we obtain
$\|\nabla^{4}\tilde{u}\|_{L^{\infty}(B_{R})}\leq\frac{C}{R^{4}}\frac{1}{|B_{2R}|}\int_{B_{2R}}|\tilde{u}|\mathrm{d}X\leq\frac{o(R^{4})}{R^{4}}\to 0\qquad\mathrm{as}\qquad R\to+\infty.$
Thus, we have
$u(x,t)=\sum_{i=0}^{3}t^{i}P_{3-i}(x).$
The homogeneous boundary conditions on $\partial\mathbb{R}_{+}^{n+1}$ give
$0=\partial_{t}u(x,0)=\sum_{i=1}^{3}it^{i-1}P_{3-i}(x)=P_{2}(x)$
and
$0=\partial^{3}_{t}u(x,0)=\sum_{i=1}^{3}i(i-1)(i-2)t^{i-3}P_{3-i}(x)=6P_{0}(x).$
This yields
$u(x,t)=P_{3}(x)+t^{2}P_{1}(x).$
Since $\liminf_{|X|\to+\infty}u(X)\geq 0$ by virtue of (2.2) and $u(x,0)=o(1)$ as $|x|\to\infty$, we conclude that $P_{3}(x)=0$ and $P_{1}(x)=c_{2}$ for some nonnegative constant $c_{2}$. Hence, the desired assertion follows. $\Box$
Remark 2.1.
If we relax the above assumption (2.2) to $u(X)\geq-C|X|^{\delta}$ for some constant $40<\delta<$ 4, which ensures that $u^{\ast}_{-}\in L^{1}(B_{1}^{+})$, then we apply the argument above to show that
$u(x,t)=t^{2}P_{1}(x).$
Lemma 2.4.
Let $u\in C^{\infty}(\overline{\mathbb{R}^{n+1}_{+}})$ be a solution of
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle u(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\Delta u(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (2.4)
Suppose there exist some $\delta,C\in\mathbb{R}_{+}$ such that
$u(X)\geq-C|X|^{-\delta}\qquad\mathrm{as}\quad|X|\to\infty.$ (2.5)
Then,
$u(x,t)=tP_{2}(x)+c_{3}t^{3},$
where $c_{3}$ is a nonnegative constant and
$\liminf_{|x|\to\infty}P_{2}(x)\geq 0.$
Proof..
Similarly, the boundary conditions become $\mathscr{B}_{0}^{3}u=0,\mathscr{B}_{3}^{3}u=0$ on $\partial\mathbb{R}_{+}^{n+1}$. By conformal covariance, $u^{\ast}$ satisfies the biharmonic boundary value problem (2.4) on +$\overline{\mathbb{R}^{n+1}_{+}}\backslash\{0\}$. Following the same lines as before, we have $u^{\ast}\in L^{1}(B_{1}^{+})$ and $u(x,t)=o(|X|^{4})$ as $|X|\to\infty$. Finally, we extend $u$ by an odd reflection
$\tilde{u}(x,t)=\begin{cases}\displaystyle u(x,t)\qquad&\mathrm{for}\qquad t\geq 0,\\ \displaystyle-u(x,-t)\qquad&\mathrm{for}\qquad t<0.\end{cases}$
A similar argument as in Lemma 2.3 combined with boundary conditions shows
$u(x,t)=tP_{2}(x)+c_{3}t^{3}.$
This together with the decay assumption at infinity yields the desired assertion. $\Box$
Remark 2.2.
If we relax the above assumption (2.5) to $u(X)\geq-C|X|^{\delta}$ as $|X|\to\infty$ for some $30<\delta<$ 3, then the same result in Lemma 2.4 still holds.
3 Biharmonic Poisson kernel
Our purpose is to derive explicit formulae of biharmonic Poisson kernels for $\mathbb{R}_{+}^{n+1}$ and $\mathbb{B}^{n+1}$.
3.1 Noncritical dimension: The upper half-space
We start with the standard Poisson kernel for $-\Delta$ in $\mathbb{R}_{+}^{n+1}$:
$P(x,t)=\frac{2}{|\mathbb{S}^{n}|}\frac{t}{(t^{2}+|x|^{2})^{\frac{n+1}{2}}}.$
Moreover, $P$ enjoys the property that let $f\in C(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$, then (a) the Poisson integral $P\ast f\in C^{\infty}(\mathbb{R}_{+}^{n+1})\cap L^{\infty}(\mathbb{R}_{+}^{n+1})$ and $\Delta(P\ast f)=0$ in $\mathbb{R}_{+}^{n+1}$; (b) for each fixed $x\in\mathbb{R}^{n}$, $\lim_{t\to 0}P\ast f(x,t)=f(x)$.
Lemma 3.1.
Suppose $f\in C^{2}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$ and $\alpha>1$, then for every fixed $x\in\mathbb{R}^{n}$,
$\lim_{t\searrow 0}\partial_{t}\int_{\mathbb{R}^{n}}\frac{t^{\alpha}}{\left(t^{2}+|x-y|^{2}\right)^{\frac{n+\alpha}{2}}}f(y)\mathrm{d}y=0.$
It is not hard to see that the above result is true provided both $f$ and $|\nabla f|$ are bounded.
Proof..
A direct calculation yields
$\partial_{t}\int_{\mathbb{R}^{n}}\frac{t^{\alpha}}{\left(t^{2}+|x-y|^{2}\right)^{\frac{n+\alpha}{2}}}f(y)\mathrm{d}y$
$=$ $\int_{\mathbb{R}^{n}}\left[\frac{\alpha t^{\alpha-1}}{{\left(t^{2}+|x-y|^{2}\right)^{\frac{n+\alpha}{2}}}}-\frac{(n+\alpha)t^{\alpha+1}}{{\left(t^{2}+|x-y|^{2}\right)^{\frac{n+\alpha+2}{2}}}}\right]f(y)\mathrm{d}y$
$=$ $\frac{\alpha}{t}\int_{\mathbb{R}^{n}}\frac{f(x-ty)}{(1+|y|^{2})^{\frac{n+\alpha}{2}}}\mathrm{d}y-\frac{n+\alpha}{t}\int_{\mathbb{R}^{n}}\frac{f(x-ty)}{(1+|y|^{2})^{\frac{n+\alpha+2}{2}}}\mathrm{d}y.$
For the first term, with some $\sigma\in(0,1)$ we have
$\int_{\mathbb{R}^{n}}\frac{f(x-ty)}{(1+|y|^{2})^{\frac{n+\alpha}{2}}}\mathrm{d}y=\int_{|y|>t^{-\sigma}}+\int_{|y|\leq t^{-\sigma}}\frac{f(x-ty)}{(1+|y|^{2})^{\frac{n+\alpha}{2}}}\mathrm{d}y:=I_{1}+I_{2}.$
For $I_{1}$, it is not hard to see that
$|I_{1}|\lesssim\int_{|y|>t^{-\sigma}}\frac{1}{(1+|y|^{2})^{\frac{n+\alpha}{2}}}\mathrm{d}y\lesssim t^{\sigma\alpha}.$
For $I_{2}$, as $t|y|<t^{1-\sigma}$ for all $|y|<t^{-\sigma}$, the Taylor expansion together with symmetry enables us to estimate
$I_{2}=$ $\int_{|y|\leq t^{-\sigma}}\frac{f(x)-t\nabla f(x)\cdot y+O(t^{2}|y|^{2})}{(1+|y|^{2})^{\frac{n+\alpha}{2}}}\mathrm{d}y$
$=$ $f(x)\int_{\mathbb{R}^{n}}\frac{1}{(1+|y|^{2})^{\frac{n+\alpha}{2}}}\mathrm{d}y-f(x)\int_{|y|>t^{-\sigma}}\frac{1}{(1+|y|^{2})^{\frac{n+\alpha}{2}}}\mathrm{d}y+\int_{|y|\leq t^{-\sigma}}\frac{O(t^{2}|y|^{2})}{(1+|y|^{2})^{\frac{n+\alpha}{2}}}\mathrm{d}y.$
As $\alpha>1$, we now take $\sigma\in\left(\frac{1}{\alpha},\frac{2}{\alpha}\right)$ for $\alpha\geq 2$ and $\sigma\in(\frac{1}{\alpha},1)$ for $\alpha<2$. Then we have
$I_{2}=\frac{|\mathbb{S}^{n-1}|}{2}B(\frac{n}{2},\frac{\alpha}{2})f(x)+O(t^{\sigma\alpha}).$
Thus, we obtain
$\int_{\mathbb{R}^{n}}\frac{f(x-ty)}{(1+|y|^{2})^{\frac{n+\alpha}{2}}}\mathrm{d}y=\frac{|\mathbb{S}^{n-1}|}{2}B(\frac{n}{2},\frac{\alpha}{2})f(x)+O(t^{\sigma\alpha}).$
For the second term, with the same $\sigma$ as above a similar argument yields
$\int_{\mathbb{R}^{n}}\frac{f(x-ty)}{(1+|y|^{2})^{\frac{n+\alpha+2}{2}}}\mathrm{d}y=\frac{|\mathbb{S}^{n-1}|}{2}B(\frac{n}{2},\frac{\alpha+2}{2})f(x)+O(t^{\sigma(\alpha+2)})+O(t^{2}).$
Therefore, with the above choice of $\sigma$ we combine these estimates to conclude that for $t\ll 1$,
$\partial_{t}\int_{\mathbb{R}^{n}}\frac{t^{\alpha}}{\left(t^{2}+|x-y|^{2}\right)^{\frac{n+\alpha}{2}}}f(y)\mathrm{d}y$
$=$ $\frac{|\mathbb{S}^{n-1}|}{2t}\left[\alpha B(\frac{n}{2},\frac{\alpha}{2})-(n+\alpha)B(\frac{n}{2},\frac{\alpha+2}{2})\right]f(x)+O(t^{\sigma\alpha-1})$
$=$ $\ O(t^{\sigma\alpha-1}),$
from which the desired conclusion follows. $\Box$
Remark 3.1.
The result in Proposition 3.1 is sharp. Otherwise, if $\alpha=1$ and let $f$ be a nonnegative function in $\mathbb{R}^{n}$ with compact support, then the classical Poisson integral
$u(x,t)=P\ast f(x,t)$
solves $\Delta u=0$ in $\mathbb{R}_{+}^{n+1}$ and $u=f(x)$ on $\partial\mathbb{R}_{+}^{n+1}$. However, for each $x\notin\mathrm{supp}(f)$, the Hopf boundary point lemma yields $\partial_{t}u(x,0)>0$.
Proof of Theorem 1.1. It suffices to show that fix $i,k\in\{0,1,2,3\}$, we define
$[(\mathscr{B}_{i}^{3}P_{k}^{3})\ast f_{k}](x,t)=\int_{\mathbb{R}^{n}}(\mathscr{B}_{i}^{3}P_{k}^{3})(x-y,t)f_{k}(y)\mathrm{d}y,$
then we have
  • (i)
    $\Delta^{2}P_{i}^{3}=0$ in $\mathbb{R}_{+}^{n+1}$.
  • (ii)
    For each pair $(i,k)$ other than $(0,2),(1,3)$, and fixed $x\in\mathbb{R}^{n}$ there hold
    $\lim_{t\to 0}[(\mathscr{B}_{i}^{3}P_{i}^{3})\ast f_{i}](x,t)=f_{i}(x);\qquad\lim_{t\to 0}[(\mathscr{B}_{i}^{3}P_{k}^{3})\ast f_{k}](x,t)=0;$
    $\lim_{t\to 0}[(\mathscr{B}_{k}^{3}P_{k}^{3})\ast f_{k}](x,t)=f_{k}(x);\qquad\lim_{t\to 0}[(\mathscr{B}_{k}^{3}P_{i}^{3})\ast f_{i}](x,t)=0.$
For (i), notice that
$\Delta^{2}|X|^{3-n}=0,\qquad\partial_{t}|X|^{3-n}=(3-n)|X|^{1-n}t,$
this yields that $P_{3}^{3}$ is biharmonic in $\mathbb{R}_{+}^{n+1}$, so is $P_{2}^{3}=\partial_{t}P_{3}^{3}$. Observe that
$\Delta|X|^{1-n}=0,\qquad\partial_{t}|X|^{1-n}=(1-n)|X|^{-1-n}t,$
then
$\Delta^{2}(t\cdot t|X|^{-1-n})=\Delta\left(2\partial_{t}(t|X|^{-1-n})\right)=\frac{2}{1-n}\partial_{t}^{2}\Delta|X|^{1-n}=0,$
from which we know that $P_{1}^{3}$ is biharmonic in $\mathbb{R}_{+}^{n+1}$. Finally, we use
$\partial_{t}(t^{2}|X|^{-1-n})=2t|X|^{-1-n}-(1+n)t^{3}|X|^{-3-n}$ (3.1)
to conclude that $\Delta^{2}P_{0}^{3}=0$ in $\mathbb{R}_{+}^{n+1}$.
In the following, we prove assertion (ii) case by case.
Case 1. Type $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{1})$ boundary condition.
The kernel $P_{0}^{3}$ shares the common property as $P$ that
$\lim_{t\to 0}P_{0}^{3}\ast f_{0}(x,t)=f_{0}(x).$ (3.2)
This is due to
$\lim_{t\to 0}\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}f_{0}(y)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y$
$=$ $\lim_{t\to 0}\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{f_{0}(x-ty)}{(1+|y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y=f_{0}(x),$
where the last identity follows from
$\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{1}{(1+|y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y=$ $\frac{2(n+1)|\mathbb{S}^{n-1}|}{|\mathbb{S}^{n}|}\int_{0}^{+\infty}\frac{r^{n-1}}{(1+r^{2})^{\frac{n+3}{2}}}\mathrm{d}r$
$=$ $\frac{(n+1)|\mathbb{S}^{n-1}|}{|\mathbb{S}^{n}|}B\left(\frac{n}{2},\frac{3}{2}\right)=1.$
Here, we emphasize $f_{0}\in C(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$ is enough to ensure (3.2), same as $P$.
Notice that
$\lim_{t\to 0}P_{1}^{3}\ast f_{1}(x,t)=-\lim_{t\to 0}t[P\ast f_{1}](x,t)=0.$
By Lemma 3.1 we have
$\lim_{t\to 0}[(\mathscr{B}_{1}^{3}P_{0}^{3})\ast f_{0}](x,t)=-\frac{2(n+1)}{|\mathbb{S}^{n}|}\lim_{t\to 0}\partial_{t}\int_{\mathbb{R}^{n}}\frac{t^{3}}{\left(t^{2}+|x-y|^{2}\right)^{\frac{n+3}{2}}}f_{0}(y)\mathrm{d}y=0.$
Again by (3.1) we obtain
$\lim_{t\to 0}[(\mathscr{B}_{1}^{3}P_{1}^{3})\ast f_{1}](x,t)=-\lim_{t\to 0}[(\partial_{t}P_{1}^{3})\ast f_{1}](x,t)=\lim_{t\to 0}[(2P-P_{0}^{3})\ast f_{1}](x,t)=f_{1}(x).$
Case 2. Type $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{2})$ boundary condition.
Rewrite
$\mathscr{B}^{3}_{2}=\partial_{t}^{2}-\overline{\Delta}=\Delta-2\overline{\Delta}.$
Recall that
$P_{0}^{3}(X)=\frac{2(n+1)}{|\mathbb{S}^{n}|}t^{3}|X|^{-3-n},\quad P_{2}^{3}(X)=-\frac{1}{(n-1)|\mathbb{S}^{n}|}t|X|^{1-n}.$
It has been known in (3.2) that
$\lim_{t\to 0}(P_{0}^{3}\ast f_{0})(x,t)=f_{0}(x).$
Also we have
$\lim_{t\to 0}(P_{2}^{3}\ast f_{2})(x,t)=$ $-\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}f_{2}(y)\mathrm{d}y$
$=$ $-\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}f_{2}(x-ty)}{(1+|y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y=0,$ (3.3)
where the last identity follows from the following estimate: For each fixed $x$, we choose $M=M(x)>2|x|$ sufficiently large such that $f_{2}(x-ty)=O(|x-ty|^{-2-\delta_{2}})$ for all $|y|>M/t$, then
$\left|\int_{\mathbb{R}^{n}}\frac{t^{2}f_{2}(x-ty)}{(1+|y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y\right|\lesssim$ $\int_{|y|<\frac{M}{t}}\frac{t^{2}}{(1+|y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y+\int_{|y|>\frac{M}{t}}\frac{t^{2}}{|y|^{n-1}|x-ty|^{2+\delta_{2}}}\mathrm{d}y$
$\lesssim$ $t+t^{-\delta_{2}}\int_{|y|>\frac{M}{t}}\frac{1}{|y|^{n+1+\delta_{2}}}\mathrm{d}y\lesssim t.$
Taking $\Delta$ to Equation (3.1) to show
$(n+1)\Delta(t^{3}|X|^{-3-n})=-\partial_{t}\Delta(t\cdot t|X|^{-1-n})=-2\partial_{t}^{2}(t|X|^{-1-n})=2\overline{\Delta}(t|X|^{-1-n}).$
Then we obtain
$\mathscr{B}^{3}_{2}P_{0}^{3}=2\overline{\Delta}(P-P_{0}^{3}).$
This implies
$\lim_{t\to 0}[(\mathscr{B}^{3}_{2}P_{0}^{3})\ast f_{0}](x,t)=2\lim_{t\to 0}\overline{\Delta}[(P-P_{0}^{3})\ast f_{0}](x,t)=0.$
Next we combine
$\Delta(t|X|^{1-n})=2\partial_{t}|X|^{1-n}=2(1-n)t|X|^{-1-n}$
and
$-2\overline{\Delta}(t|X|^{1-n})=2t\partial_{t}^{2}(|X|^{1-n})=2(1-n)[t|X|^{-1-n}-(n+1)t^{3}|X|^{-3-n}]$ (3.4)
to show
$\mathscr{B}^{3}_{2}P_{2}^{3}=P+(P-P_{0}^{3}),$
from which we have
$\lim_{t\to 0}[(\mathscr{B}^{3}_{2}P_{2}^{3})\ast f_{2}](x,t)=f_{2}(x).$ (3.5)
Case 3. Type $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{3})$ boundary condition.
A direct computation yields
$\mathscr{B}^{3}_{1}P^{3}_{1}=\partial_{t}\left(\frac{2}{|\mathbb{S}^{n}|}t^{2}|X|^{-1-n}\right)=$ $\frac{2}{|\mathbb{S}^{n}|}\left(\frac{2t}{(t^{2}+|x|^{2})^{\frac{n+1}{2}}}-\frac{(n+1)t^{3}}{(t^{2}+|x|^{2})^{\frac{n+3}{2}}}\right)$
::=:= $2P-P_{0}^{3},$
Then, it is not hard to see that $\lim_{t\to 0}[(\mathscr{B}^{3}_{1}P^{3}_{1})\ast f_{1}](x,t)=f_{1}(x)$.
Since
$\mathscr{B}^{3}_{1}P^{3}_{3}=\frac{1}{(n-1)|\mathbb{S}^{n}|}t|X|^{1-n},$
similar to (3.1) we have
$\lim_{t\to 0}[(\mathscr{B}_{1}^{3}P_{3}^{3})\ast f_{3}](x,t)=$ $\lim_{t\to 0}\int_{\mathbb{R}^{n}}\frac{1}{(n-1)|\mathbb{S}^{n}|}\frac{tf_{3}(y)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y$
$=$ $\lim_{t\to 0}\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}f_{3}(x-ty)}{(1+|y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y=0.$
Recall that
$\mathscr{B}^{3}_{3}=\partial_{t}\Delta+2\overline{\Delta}\partial_{t}$
and
$P_{1}^{3}(X)=-\frac{2}{|\mathbb{S}^{n}|}t^{2}|X|^{-1-n},\quad P_{3}^{3}(X)=\frac{1}{(n-1)(n-3)|\mathbb{S}^{n}|}|X|^{3-n}.$
A direct computation yields
$\partial_{t}\Delta(t\cdot t|X|^{-1-n})=$ $2\partial_{t}^{2}(t|X|^{-1-n});$
$2\overline{\Delta}\partial_{t}(t\cdot t|X|^{-1-n})=$ $2\overline{\Delta}\left[2t|X|^{-1-n}-(1+n)t^{3}|X|^{-3-n}\right].$
Since $t|X|^{-1-n}$ is harmonic in $\mathbb{R}_{+}^{n+1}$, we have
$\mathscr{B}^{3}_{3}P_{1}^{3}=-2\overline{\Delta}(P-P_{0}^{3}).$
This directly implies
$\lim_{t\to 0}[(\mathscr{B}^{3}_{3}P_{1}^{3})\ast f_{1}](x,t)=-2\lim_{t\to 0}\overline{\Delta}[(P-P_{0}^{3})\ast f_{1}](x,t)=0.$
A direct computation yields
$\partial_{t}\Delta|X|^{3-n}=$ $(3-n)\Delta(t|X|^{1-n})=2(n-1)(n-3)t|X|^{-1-n};$
$2\overline{\Delta}\partial_{t}|X|^{3-n}=$ $2(3-n)\overline{\Delta}(t|X|^{1-n}).$
These together with (3.4) give
$\mathscr{B}^{3}_{3}P_{3}^{3}=P-\overline{\Delta}(P-P_{0}^{3}).$
This directly implies
$\lim_{t\to 0}[(\mathscr{B}^{3}_{3}P_{3}^{3})\ast f_{3}](x,t)=$ $f_{3}(x)-\overline{\Delta}\lim_{t\to 0}[(P-P_{0}^{3})\ast f_{3}](x,t)=f_{3}(x).$ (3.6)
Case 4. Type $(\mathscr{B}^{3}_{2},\mathscr{B}^{3}_{3})$ boundary condition.
Since we already have (3.5) and (3.6), it remains to show
$\lim_{t\to 0}[(\mathscr{B}^{3}_{2}P_{3}^{3})\ast f_{3}](x,t)=0,\qquad\lim_{t\to 0}[(\mathscr{B}^{3}_{3}P_{2}^{3})\ast f_{2}](x,t)=0.$
To this end, notice that
$\mathscr{B}^{3}_{2}|X|^{3-n}=$ $\Delta|X|^{3-n}-2\overline{\Delta}|X|^{3-n}$
$=$ $2(3-n)|X|^{1-n}-2(3-n)[(n-1)|X|^{-1-n}t^{2}+|X|^{1-n}]$
$=$ $2(n-1)(n-3)t^{2}|X|^{-1-n},$
whence,
$\mathscr{B}^{3}_{2}P_{3}^{3}=\frac{2}{|\mathbb{S}^{n}|}t^{2}|X|^{-1-n}.$
This implies
$\lim_{t\to 0}[(\mathscr{B}^{3}_{2}P_{3}^{3})\ast f_{3}](x,t)=\lim_{t\to 0}[tP\ast f_{3}(x,t)]=0.$
Since
$\mathscr{B}^{3}_{3}(t|X|^{1-n})=$ $\partial_{t}\Delta(t|X|^{1-n})+2\overline{\Delta}\partial_{t}(t|X|^{1-n})$
$=$ $2\partial_{t}^{2}|X|^{1-n}+2\overline{\Delta}|X|^{1-n}+2t\overline{\Delta}\partial_{t}|X|^{1-n}$
$=$ $2(1-n)t^{2}\overline{\Delta}|X|^{-1-n},$
we have
$\mathscr{B}^{3}_{3}P_{2}^{3}=\frac{2}{|\mathbb{S}^{n}|}\overline{\Delta}(t^{2}|X|^{-1-n}).$
This gives
$\lim_{t\to 0}[(\mathscr{B}^{3}_{3}P_{2}^{3})\ast f_{2}](x,t)=\overline{\Delta}\lim_{t\to 0}[tP\ast f_{2}(x,t)]=0.$
$\Box$
3.2 Noncritical dimension: The unit ball
Our purpose is to give explicit representation formulae of biharmonic Poisson kernels for $\mathbb{B}^{n+1}$ with the help of conformal equivalence between $\mathbb{B}^{n+1}$ and $\mathbb{R}_{+}^{n+1}$,
Let $F:(\mathbb{R}^{n+1}_{+},|\mathrm{d}X|^{2})\to(\mathbb{B}^{n+1},|\mathrm{d}\xi|^{2})$
$\xi=F(X)=-\mathbf{e}_{n+1}+\frac{2(X+\mathbf{e}_{n+1})}{|X+\mathbf{e}_{n+1}|^{2}}$ (3.7)
be a conformal map with the property that
$F^{\ast}(|\mathrm{d}\xi|^{2})=\left(\frac{2}{|X+\mathbf{e}_{n+1}|^{2}}\right)^{2}|\mathrm{d}X|^{2}:=U_{0}(X)^{\frac{4}{n-3}}|\mathrm{d}X|^{2}.$
Clearly, $F$ is exact an inversion of the sphere centered at $-\mathbf{e}_{n+1}$ of radius $\sqrt{2}$ and thus $F=F^{-1}$.
The following important but elementary identity will be frequently used:
$2|\xi+\mathbf{e}_{n+1}|^{-1}|\eta+\mathbf{e}_{n+1}|^{-1}|\xi-\eta|=|F^{-1}(\xi)-F^{-1}(\eta)|,\quad\forall~{}\xi,\eta\in\overline{\mathbb{B}^{n+1}}\setminus\{-\mathbf{e}_{n+1}\}.$ (3.8)
This can be verified by a straightforward calculation.
Given $U\in C^{\infty}(\overline{\mathbb{B}^{n+1}})$, we set
$u=U\circ FU_{0}.$ (3.9)
Clearly, $u(X)=O^{(k)}(|X|^{3-n})$ as $|X|\to\infty$, where $O^{(k)}(|X|^{3-n})$ means the $k$-th derivatives of $v$ belong to $O(|X|^{3-n-k})$ for every $k\in\mathbb{N}$. Moreover, by conformal covariance of the Paneitz operator, we have
$\Delta^{2}u=P_{4}^{|\mathrm{d}X|^{2}}(u)=P_{4}^{|\mathrm{d}X|^{2}}(U\circ FU_{0})=$ $U_{0}^{\frac{n+5}{n-3}}P_{4}^{U_{0}^{\frac{4}{n-3}}|\mathrm{d}X|^{2}}(U\circ F)$
$=$ $U_{0}^{\frac{n+5}{n-3}}(P_{4}^{|\mathrm{d}\xi|^{2}}U)\circ F=U_{0}^{\frac{n+5}{n-3}}(\Delta^{2}U)\circ F.$
Similarly, by (1.4) we have
$\left(\mathscr{B}_{k}^{3}\right)_{|\mathrm{d}X|^{2}}(u)=U_{0}^{\frac{n+2k-3}{n-3}}\left[\left(\mathscr{B}_{k}^{3}\right)_{|\mathrm{d}\xi|^{2}}U\right]\circ F\qquad\mathrm{~{}~{}for~{}~{}}\quad 0\leq k\leq 3.$ (3.10)
Proof of Theorem 1.2. For assertion $(i)$, we first handle $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{2})$ type boundary condition.
Let $u$ define as in (3.9), equivalently,
$U(\xi)=\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}u\circ F^{-1}(\xi).$
By virtue of (3.10) we set
$f_{0}(x)=\left(\frac{2}{1+|x|^{2}}\right)^{\frac{n-3}{2}}\bar{f}_{0}\circ F\left((x,0)\right)\quad\mathrm{and}\quad f_{2}(x)=\left(\frac{2}{1+|x|^{2}}\right)^{\frac{n+1}{2}}\bar{f}_{2}\circ F((x,0)).$
Then for $|X|,|x|\gg 1$, we know
$u(X)=O(|X|^{3-n}),\quad f_{0}(x)=O(|x|^{3-n}),\quad f_{2}(x)=O(|x|^{-1-n}),$
and $u$ solves
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle u(x,0)=f_{0}\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial^{2}_{t}u(x,0)-\overline{\Delta}u(x,0)=f_{2}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (3.11)
As a direct consequence of Remark 2.2, the solution to the above problem (3.11) is unique. This together with Theorem 1.1 gives
$u(X)=\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}f_{0}(y)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y-\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{tf_{2}(y)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y.$
If we let
$\xi=$ $F(X)=\left(\frac{2x}{(1+t)^{2}+|x|^{2}},\frac{1-t^{2}-|x|^{2}}{(1+t)^{2}+|x|^{2}}\right),$
$\eta=$ $F((y,0))=\left(\frac{2y}{1+|y|^{2}},\frac{1-|y|^{2}}{1+|y|^{2}}\right),$
then
$t\circ F^{-1}=\frac{1-|\xi|^{2}}{|\xi+\mathbf{e}_{n+1}|^{2}}$
and from (3.8) that
$2|\xi+\mathbf{e}_{n+1}|^{-1}|\eta+\mathbf{e}_{n+1}|^{-1}|\xi-\eta|=|(x,t)-(y,0)|.$
Therefore, putting these facts together we obtain
$U(\xi)=$ $\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{(t\circ F^{-1})^{3}f_{0}\circ F^{-1}(\eta)\left(\frac{2}{|\eta+\mathbf{e}_{n+1}|^{2}}\right)^{n}}{((t\circ F^{-1})^{2}+|(x-y)\circ F^{-1}|^{2})^{\frac{n+3}{2}}}\mathrm{d}y$
$-\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{t\circ F^{-1}f_{2}\circ F^{-1}(\eta)\left(\frac{2}{|\eta+\mathbf{e}_{n+1}|^{2}}\right)^{n}}{((t\circ F^{-1})^{2}+|(x-y)\circ F^{-1}|^{2})^{\frac{n-1}{2}}}\mathrm{d}y$
$=$ $\frac{(n+1)\left(1-|\xi|^{2}\right)^{3}}{4|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\bar{f}_{0}(\eta)}{|\xi-\eta|^{n+3}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta)-\frac{(1-|\xi|^{2})}{2(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\bar{f}_{2}(\eta)}{|\xi-\eta|^{n-1}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta).$
As indicated in the preceding discussion, for other pairs of conformal boundary operators it suffices to show the uniqueness.
The uniqueness for $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{1})$ is trivial. For $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{3})$, the uniqueness follows from a similar argument above combined with Lemma 2.3 or Remark 2.1.
The remaining boundary condition is $(\mathscr{B}^{3}_{2},\mathscr{B}^{3}_{3})$, the uniqueness becomes a little subtle. Suppose $U$ is a smooth solution of
$\begin{cases}\displaystyle\Delta^{2}U=0\qquad\qquad&\mathrm{in}\qquad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{2}U=0&\mathrm{on}\qquad\mathbb{S}^{n},\\ \displaystyle\mathscr{B}^{3}_{3}U=0&\mathrm{on}\qquad\mathbb{S}^{n},\end{cases}$
and let $u$ define as in (3.9), then $u$ solves
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\partial_{t}^{2}u(x,0)-\overline{\Delta}u(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta u(x,0)+2\overline{\Delta}\partial_{t}u(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\end{cases}$
with $u(X)=O(|X|^{3-n})$ as $|X|\to\infty$.
By Lemma 2.3 and Theorem 1.1, the unique solution for $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{3})$ is
$u(x,t)=\frac{2}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}\partial_{t}u(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\mathrm{d}y.$
Since $\partial_{t}u(y,0)=O(|y|^{1-n})$ as $|y|\to\infty$ by definition (3.9) of $u$, we have
$\lim_{t\to 0}\frac{2t}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t\partial_{t}u(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\mathrm{d}y=0.$
So, by the uniqueness for $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{2})$ due to Lemma 2.4 we obtain $u=0$ and $U=0$ as well.
For assertion $(ii)$, the multiplicity of smooth solutions for the remaining two boundary conditions is a direct consequence of the nex Lemma 3.2. $\Box$
Lemma 3.2.
The following homogeneous biharmonic boundary value problems:
$\begin{cases}\displaystyle\Delta^{2}U=0\qquad&\mathrm{in}\qquad\mathbb{B}^{n+1}\\ \displaystyle\mathscr{B}^{3}_{0}(U)=0\qquad&\mathrm{on}\qquad\mathbb{S}^{n}\\ \displaystyle\mathscr{B}^{3}_{3}(U)=0\qquad&\mathrm{on}\qquad\mathbb{S}^{n}\end{cases}$ (3.12)
and
$\begin{cases}\displaystyle\Delta^{2}U=0\qquad&\mathrm{in}\qquad\mathbb{B}^{n+1}\\ \displaystyle\mathscr{B}^{3}_{1}(U)=0\qquad&\mathrm{on}\qquad\mathbb{S}^{n}\\ \displaystyle\mathscr{B}^{3}_{2}(U)=0\qquad&\mathrm{on}\qquad\mathbb{S}^{n}\end{cases}$ (3.13)
have nontrivial smooth solutions.
Proof..
For every $\varphi\in C_{\mathrm{c}}^{\infty}(\mathbb{R}^{n})$, we introduce
$u_{\varphi}(x,t)=-\frac{2}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}\varphi(y)}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\mathrm{d}y$
for the boundary condition $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{3})$; and
$u_{\varphi}(x,t)=\frac{1}{(n-1)(n-3)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{\varphi(y)}{(t^{2}+|x-y|^{2})^{\frac{n-3}{2}}}\mathrm{d}y$
for the boundary condition $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{2})$.
As in the proof of Theorem 1.1, we know that $u=u_{\varphi}$ solves
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1}\\ \displaystyle u(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\\ \displaystyle\partial_{t}\Delta u(x,0)+2\overline{\Delta}\partial_{t}u(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\end{cases}$
and
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1}\\ \displaystyle\partial_{t}u(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\\ \displaystyle\partial^{2}_{t}u(x,0)-\overline{\Delta}u(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\end{cases},$
respectively. Finally, we may obtain nontrivial smooth solutions of (3.12) and (3.13) via
$U_{\varphi}(\xi)=\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}u_{\varphi}\circ F^{-1}(\xi).$
$\Box$
This also reflects the loss of ‘the complementing condition’ for $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{3})$ and $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{2})$, under which boundary conditions problem (3.12) or (3.13) fails to have the uniqueness of the solution as explained in [1].
3.3 Critical dimension
Motivated by Chen-Shi [10, Theorem 1(1)] we are able to derive explicit formulae of biharmonic Poisson kernels on $\mathbb{R}_{+}^{4}$ and $\mathbb{B}^{4}$.
Proof of Theorem 1.4. Based on the proof of Theorems 1.1 and 1.2, it suffices to point out several significant differences.
(1) Observe that
$\mathscr{B}^{3}_{3}P^{3}_{3}(X)=(\partial_{t}\Delta+2\overline{\Delta}\partial_{t})P^{3}_{3}(X)=\frac{8}{|\mathbb{S}^{3}|}\frac{t^{3}}{(t^{2}+|x|^{2})^{3}}=P^{3}_{0}(X),$
so, $\lim_{t\to 0}[(\mathscr{B}^{3}_{3}P^{3}_{3})\ast f_{3}](x,t)=f_{3}(x)$. Also, we can apply
$\mathscr{B}^{3}_{1}P^{3}_{3}(X)=$ $\frac{t}{2|\mathbb{S}^{3}|(t^{2}+|x|^{2})}=-P_{2}^{3}(X),$
$\mathscr{B}^{3}_{2}P^{3}_{3}(X)=$ $\frac{2}{|\mathbb{S}^{3}|}\frac{t^{2}}{(t^{2}+|x|^{2})^{2}}=-P_{1}^{3}(X)$
to show
$\lim_{t\to 0}[(\mathscr{B}^{3}_{1}P^{3}_{3})\ast f_{3}](x,t)=0\qquad\mathrm{and}\qquad\lim_{t\to 0}[(\mathscr{B}^{3}_{2}P^{3}_{3})\ast f_{3}](x,t)=0.$
(2) Suppose $U$ is a solution of (1.7). In critical dimension four, we introduce
$u(x,t)=U\circ F(x,t)+\log\frac{2}{(1+t)^{2}+|x|^{2}},\qquad\hat{U}_{0}(x,t):=\log\frac{2}{(1+t)^{2}+|x|^{2}}$
and
$f_{0}(x)=$ $\bar{f}_{0}\circ F(x,0)+\hat{U}(x,0),$
$f_{k}(x)=$ $\left[\bar{f}_{k}\circ F(x,0)+(T_{k}^{3})_{g_{\mathbb{S}^{3}}}\right]e^{k\hat{U}_{0}(x,0)},k\in\{1,2,3\}.$
This essentially follows from conformal covariance of $P_{4}^{g}$ and $\mathscr{B}_{k}^{3}$ that
$e^{2u}|\mathrm{d}X|^{2}=F^{\ast}(e^{2U}|\mathrm{d}\xi|^{2}),\quad e^{2\hat{U}_{0}}|\mathrm{d}X|^{2}=F^{\ast}(|\mathrm{d}\xi|^{2})$
and
$f_{k}=(\mathscr{B}_{k}^{3})_{|\mathrm{d}X|^{2}}(u)=$ $(\mathscr{B}_{k}^{3})_{|\mathrm{d}X|^{2}}(U\circ F)+(\mathscr{B}_{k}^{3})_{|\mathrm{d}X|^{2}}(\hat{U}_{0})$
[…]
$=$ $e^{k\hat{U}_{0}}(\mathscr{B}_{k}^{3})_{F^{\ast}(|\mathrm{d}\xi|^{2})}(U\circ F)+(T_{k}^{3})_{g_{\mathbb{S}^{3}}}e^{k\hat{U}_{0}}$
$=$ $e^{k\hat{U}_{0}}[(\mathscr{B}_{3}^{3})_{|\mathrm{d}\xi|^{2}}(U)]\circ F+(T_{k}^{3})_{g_{\mathbb{S}^{3}}}e^{k\hat{U}_{0}}$
$=$ $e^{k\hat{U}_{0}}(\bar{f}_{k}\circ F+(T_{k}^{3})_{g_{\mathbb{S}^{3}}}).$
In particular, $(T_{1}^{3})_{g_{\mathbb{S}^{3}}}=1,(T_{2}^{3})_{g_{\mathbb{S}^{3}}}=2,(T_{3}^{3})_{g_{\mathbb{S}^{3}}}=4$.
Suppose $u$ is a solution of (1.6). By the biharmonic Poisson kernel on $\mathbb{R}_{+}^{4}$ we have
$u(x,t)=$ $\int_{\partial\mathbb{R}_{+}^{4}}[P_{i}^{3}(x-y,t)f_{i}(y)+P_{j}^{3}(x-y,t)f_{j}(y)]\mathrm{d}y$
$=$ $\int_{\partial\mathbb{R}_{+}^{4}}[P_{i}^{3}(x-y,t)e^{i\hat{U}_{0}(y,0)}f_{i}\circ F(y)+P_{j}^{3}(x-y,t)e^{j\hat{U}_{0}(y,0)}f_{j}\circ F(y)]\mathrm{d}y+I(x,t),$
where
$I(x,t)=\begin{cases}\int_{\partial\mathbb{R}_{+}^{4}}[P_{0}^{3}(x-y,t)\hat{U}_{0}(y,0)+P_{j}^{3}(x-y,t)e^{j\hat{U}_{0}(y,0)}(T_{j}^{3})_{g_{\mathbb{S}^{3}}}]\mathrm{d}y,\\ \int_{\partial\mathbb{R}_{+}^{4}}[P_{i}^{3}(x-y,t)e^{i\hat{U}_{0}(y,0)}f_{i}\circ F(y)+P_{j}^{3}(x-y,t)e^{j\hat{U}_{0}(y,0)}f_{j}\circ F(y)]\mathrm{d}y\quad\mathrm{~{}~{}if~{}~{}}i\neq 0.\end{cases}$
Then it is not hard to see that $I=\hat{U}_{0}$ when $(i,j)=(0,1),(0,2)$; $I=\hat{U}_{0}+C,C\in\mathbb{R}$ when $(i,j)=(1,3),(2,3)$.
Hence, using variable changes as in the proof of Theorem 1.2 we obtain
$U(\xi)=\int_{\mathbb{S}^{3}}\bar{P}^{3}_{i}(\xi-\eta)\bar{f}_{i}(\eta)+\bar{P}^{3}_{j}(\xi-\eta)\bar{f}_{j}(\eta)\mathrm{d}V_{{\mathbb{S}^{3}}}(\eta)$
when $(i,j)=(0,1),(0,2)$; and
$U(\xi)=\int_{\mathbb{S}^{3}}\bar{P}^{3}_{i}(\xi-\eta)\bar{f}_{i}(\eta)\mathrm{d}V_{{\mathbb{S}^{3}}}(\eta)-\frac{1}{2|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}\log|\xi-\eta|\bar{f}_{3}(\eta)\mathrm{d}V_{\mathbb{S}^{3}}(\eta)+C,$ (3.14)
when $(i,j)=(1,3),(2,3)$.
It remains to determine the constant $C$ in the following way: On $\mathbb{S}^{3}$, a compatible condition of $\bar{f}_{3}$ arises from the fact that $\mathrm{ker}P_{3}=\mathbb{R}$ by virtue of [10, Theorem 1(1)], more explicitly,
$\int_{\mathbb{S}^{3}}\bar{f}_{3}\mathrm{d}V_{\mathbb{S}^{3}}=\int_{\mathbb{S}^{3}}(-\frac{\partial\Delta U}{\partial r}-2\Delta_{\mathbb{S}^{3}}\frac{\partial U}{\partial r}-2\Delta_{\mathbb{S}^{3}}U)\mathrm{d}V_{\mathbb{S}^{3}}=0.$ (3.15)
Restricting $\xi$ to $\mathbb{S}^{3}$ in (3.14) we find
$U(\xi)=-\frac{1}{2|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}\log|\xi-\eta|\bar{f}_{3}(\eta)\mathrm{d}V_{\mathbb{S}^{3}}+C.$
Integrating the above equation over $\mathbb{S}^{3}$ to give
$\int_{\mathbb{S}^{3}}U\mathrm{d}V_{\mathbb{S}^{3}}=-\frac{1}{2|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}\bar{f}_{3}(\eta)\mathrm{d}V_{\mathbb{S}^{3}}(\eta)\int_{\mathbb{S}^{3}}\log|\xi-\eta|\mathrm{d}V_{\mathbb{S}^{3}}(\xi)+C|\mathbb{S}^{3}|=C|\mathbb{S}^{3}|,$
where we have used (3.15) and the fact that $\int_{\mathbb{S}^{3}}\log|\xi-\eta|\mathrm{d}V_{\mathbb{S}^{3}}(\xi),\forall~{}\eta\in\mathbb{S}^{3}$ is constant. This follows that
$C=\frac{1}{|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}U\mathrm{d}V_{{\mathbb{S}^{3}}}.$
Then the desired assertion follows by inserting the above constant into (3.14). $\Box$
It is not hard to see that Theorem 1.4 could give a direct proof of [6, Proposition 2.2, Lemma 2.3]).
We point out that the proof of Theorems 1.1 and 1.2 is still valid for the case of $n=2$. However, we shall not go deep into this, since it has been well studied by authors [11].
Thanks to biharmonic Poisson kernels on $\mathbb{R}^{n+1}$, we can recover a special case in [7, Theorem 1.4], which clarifies the relationship between extrinsic and intrinsic GJMS operators, and leave its proof to Appendix A for clarity. Something about $P_{1}$ and $P_{3}$.
Proposition 3.1 (cf. [7, Theorem 1.4] and [11, Porposition 2.2]).
Suppose $U$ is a biharmonic function in $\mathbb{B}^{n+1}$ for $n\geq 2$, then there hold
$\mathscr{B}^{3}_{3}U=2P_{3}U\qquad\mathrm{and}\qquad\mathscr{B}^{3}_{2}\left(U\right)=2P_{1}(\mathscr{B}^{3}_{1}U)\qquad\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n},$
where $P_{1},P_{3}$ are the fractional GJMS operators with respect to the round metric, see [7, 10].
Proposition 3.1 indicates that a Dirichlet datum is equivalent to a $\mathscr{B}_{3}^{3}$ datum since $P_{3}$ is invertible; a $\mathscr{B}_{1}^{3}$ datum is equivalent to a $\mathscr{B}_{2}^{3}$ datum since $P_{1}$ is invertible. That is exact the reason why we can construct multiple smooth solutions for a biharmonic equation on $\mathbb{B}^{n+1}$ with these two types of boundary conditions.
4 Biharmonic Green function
For $n\geq 2$, we let $X=(x,t)\in\mathbb{R}_{+}^{n+1},Y=(y,s)\in\mathbb{R}_{+}^{n+1}$. We define $\bar{X}=(x,-t)$, and $\xi=F(X),\eta=F(Y)$ with $F$ being the conformal map as in (3.7).
4.1 Noncritical dimension
Recall that the fundamental solution of $\Delta^{2}$ in $\mathbb{R}^{n+1}$ is given by
$\Gamma(X-Y)=\Gamma(|X-Y|)=\frac{1}{2(n-1)(n-3)|\mathbb{S}^{n}|}|X-Y|^{3-n}.$
Proof of Theorem 1.3. We denote by $G^{(i,j)}$ and $\bar{G}^{(i,j)}$ the Green functions in +$\mathbb{R}^{n+1}_{+}$ and $\mathbb{B}^{n+1}$, respectively.
Fix $X\in\mathbb{R}_{+}^{n+1}$, then for all $Y\in\partial\mathbb{R}_{+}^{n+1}$, direct computations (cf. calculations in the proof of Theorem 1.1) show
$\begin{split}\mathscr{B}^{3}_{0}(\Gamma(X-Y))=&\frac{1}{2}P^{3}_{3}(x-y,t);\\ \mathscr{B}^{3}_{1}(\Gamma(X-Y))=&-\frac{t}{2(n-1)|\mathbb{S}^{n}|(|x-y|^{2}+t^{2})^{\frac{n-1}{2}}}=\frac{1}{2}P^{3}_{2}(x-y,t);\\ \mathscr{B}^{3}_{2}(\Gamma(X-Y))=&\frac{1}{|\mathbb{S}^{n}|}\frac{t^{2}}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}=-\frac{1}{2}P^{3}_{1}(x-y,t);\\ \mathscr{B}^{3}_{3}(\Gamma(X-Y))=&-\frac{(n+1)}{|\mathbb{S}^{n}|}\frac{t^{3}}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}=-\frac{1}{2}P^{3}_{0}(x-y,t).\end{split}$ (4.1)
The above formulae demonstrate an intimate connection between the Poisson kernel and fundamental function.
By definition of Green functions on both +$\mathbb{R}^{n+1}_{+}$ and $\mathbb{B}^{n+1}$, we invoke the conformal invariance of $P_{4}^{g}$ to show
$\bar{G}^{(i,j)}(\xi,\eta)=G^{(i,j)}(X,Y)\left(\frac{2}{|\eta+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}$ (4.2)
by setting $u=U\circ FU_{0}$ and $f=\bar{f}\circ FU_{0}^{(n+5)/(n-3)}$. For brevity, we let
$H^{(i,j)}(X,Y)=G^{(i,j)}(X,Y)-\Gamma(X-Y)$
and
$\bar{H}^{(i,j)}(\xi,\eta)=\bar{G}^{(i,j)}(\xi,\eta)-\Gamma(\xi-\eta).$
Then $H^{(i,j)}$ solves
$\begin{cases}\displaystyle\Delta^{2}H^{(i,j)}=0\qquad&\mathrm{in}\qquad\mathbb{R}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(H^{(i,j)})=-\mathscr{B}^{3}_{i}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{j}(H^{(i,j)})=-\mathscr{B}^{3}_{j}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}^{n+1}.\end{cases}$
For $(i,j)=(0,2)$, we find
$H^{(0,2)}(X,Y)=-\Gamma(\bar{X}-Y).$
This gives
$G^{(0,2)}(X,Y)=\Gamma(X-Y)-\Gamma(\bar{X}-Y).$
Let $\xi^{\ast}=|\xi|^{-2}\xi$ whenever $\xi\neq 0$, then it is not hard to see that $F(\bar{X})=\xi^{\ast}$. Using
$|X+\mathbf{e}_{n+1}||\xi+\mathbf{e}_{n+1}|=2,\quad|\bar{X}+\mathbf{e}_{n+1}||\xi^{\ast}+\mathbf{e}_{n+1}|=2,$
we obtain
$|\xi|^{2}=\frac{|\bar{X}+\mathbf{e}_{n+1}|^{2}}{|X+\mathbf{e}_{n+1}|^{2}}=\frac{|\xi+\mathbf{e}_{n+1}|^{2}}{|\xi^{\ast}+\mathbf{e}_{n+1}|^{2}}.$
Also, the following elementary identity
$2|\xi^{*}-\eta|=|\xi^{\ast}+\mathbf{e}_{n+1}||\eta+\mathbf{e}_{n+1}||\bar{X}-Y|$
gives
$2|\xi||\xi^{*}-\eta|=|\xi+\mathbf{e}_{n+1}||\eta+\mathbf{e}_{n+1}||\bar{X}-Y|.$ (4.3)
Consequently, by (4.2), (3.8) and (4.3) we obtain
$\bar{G}^{(0,2)}(\xi,\eta)=$ $\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\left(\frac{2}{|\eta+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}G^{(0,2)}(X,Y)$
$=$ $\Gamma(\xi-\eta)-\Gamma(|\xi|\cdot|\xi^{*}-\eta|).$
For $(i,j)=(1,3)$, similarly we have
$G^{(1,3)}(X,Y)=\Gamma(X-Y)+\Gamma(\bar{X}-Y)$
and
$\bar{G}^{(1,3)}(\xi,\eta)=\Gamma(\xi-\eta)+\Gamma(|\xi|\cdot|\xi^{*}-\eta|).$
For $(i,j)=(0,1),(2,3)$, the Boggio formula (see, Boggio [2] or Gazzola-Grunau-Sweers [15, Lemma 2.27]) gives
$\bar{G}^{(0,1)}(\xi,\eta)=$ $\frac{1}{4|\mathbb{S}^{n}|}|\xi-\eta|^{3-n}\int_{1}^{\frac{|\xi||\xi^{\ast}-\eta|}{|\xi-\eta|}}(s^{2}-1)s^{-n}\mathrm{d}s$
$=$ $\Gamma(\xi-\eta)+\frac{1}{2}\Gamma(|\xi|\cdot|\xi^{\ast}-\eta|)\left[(n-3)\frac{|\xi-\eta|^{2}}{|\xi|^{2}\cdot|\xi^{\ast}-\eta|^{2}}-(n-1)\right].$
Set
$f_{2}(y)=-\mathscr{B}^{3}_{2}(\Gamma(X-Y))=\frac{1}{2}P^{3}_{1}(x-y,t)$
and
$f_{3}(y)=-\mathscr{B}^{3}_{3}(\Gamma(X-Y))=\frac{1}{2}P^{3}_{0}(x-y,t).$
We apply Theorem 1.1 together with (4.1) to solve
$H^{(2,3)}(X,Y)=$ $\int_{\mathbb{R}^{n}}P^{3}_{2}(y-z,s)f_{2}(z)\mathrm{d}z+\int_{\mathbb{R}^{n}}P^{3}_{3}(y-z,s)f_{3}(z)\mathrm{d}z$
$=$ $\frac{1}{2}\int_{\mathbb{R}^{n}}P^{3}_{2}(y-z,s)P^{3}_{1}(x-z,t)\mathrm{d}z+\frac{1}{2}\int_{\mathbb{R}^{n}}P^{3}_{3}(y-z,s)P^{3}_{0}(x-z,t)\mathrm{d}z.$
A similar trick gives
$H^{(0,1)}(X,Y)=-H^{(2,3)}(X,Y).$
Hence, by (3.8) and (4.2) we have
$\bar{H}^{(2,3)}(\xi,\eta)=$ $\left(\frac{2}{|\eta+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}H^{(2,3)}(X,Y)$
$=$ $-\left(\frac{2}{|\eta+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}H^{(0,1)}(X,Y)=-\bar{H}^{(0,1)}(\xi,\eta),$
from which we conclude that
$\bar{G}^{(2,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)-\frac{1}{2}\Gamma(|\xi|\cdot|\xi^{\ast}-\eta|)\left[(n-3)\frac{|\xi-\eta|^{2}}{|\xi|^{2}\cdot|\xi^{\ast}-\eta|^{2}}-(n-1)\right].$
Until now, we have already obtained biharmonic Green functions on $\mathbb{R}_{+}^{n+1}$ at the same time.
Observe that
$|\xi|^{2}|\xi^{\ast}-\eta|^{2}=|\xi-\eta|^{2}+(1-|\xi|^{2})(1-|\eta|^{2}),\qquad\forall~{}\xi,\eta\in\overline{\mathbb{B}^{n+1}},$
then it follows that
$\bar{G}^{(2,3)}(\xi,\eta)\geq$ $\Gamma(\xi-\eta)+\Gamma(|\xi|\cdot|\xi^{*}-\eta|)$
$=$ $\bar{G}^{(1,3)}(\xi,\eta)\geq\bar{G}^{(0,2)}(\xi,\eta)\geq 0.$
On the other hand, we have
$[\bar{G}^{(0,2)}-\bar{G}^{(0,1)}](\xi,\eta)=\frac{n-3}{2}\Gamma(|\xi|\cdot|\xi^{\ast}-\eta|)\left(1-\frac{|\xi-\eta|^{2}}{|\xi|^{2}\cdot|\xi^{\ast}-\eta|^{2}}\right)\geq 0.$
Consequently, we combine these facts together to conclude the second assertion. $\Box$
4.2 Critical dimension
The fundamental solution for $\Delta^{2}$ on $\mathbb{R}_{+}^{4}$ is given by
$\Gamma(X-Y)=-\frac{1}{4|\mathbb{S}^{3}|}\log|X-Y|.$
Proof of Theorem 1.5. Fix $X\in\mathbb{R}^{4}_{+}$, then for all $Y\in\partial\mathbb{R}^{4}_{+}$, it is straightforward to show
$\mathscr{B}^{3}_{0}(\Gamma(X-Y))=$ $\frac{1}{2}P^{3}_{3}(x-y,t);$
$\mathscr{B}^{3}_{1}(\Gamma(X-Y))=$ $\frac{1}{2}P^{3}_{2}(x-y,t);$
$\mathscr{B}^{3}_{2}(\Gamma(X-Y))=$ $-\frac{1}{2}P^{3}_{1}(x-y,t);$
$\mathscr{B}^{3}_{3}(\Gamma(X-Y))=$ $-\frac{1}{2}P^{3}_{0}(x-y,t).$
We next derive the Green function on $\mathbb{B}^{4}$ associated with the unique smooth solution of
$\begin{cases}\displaystyle\Delta^{2}U=f\qquad\quad&\mathrm{in}\qquad\mathbb{B}^{4},\\ \displaystyle\mathscr{B}^{3}_{i}U=0\qquad&\mathrm{on}\qquad\mathbb{S}^{3},\\ \displaystyle\mathscr{B}^{3}_{j}U=0\qquad&\mathrm{on}\qquad\mathbb{S}^{3},\end{cases}$
where $\bar{f}\in C^{\infty}(\overline{\mathbb{B}^{4}})$. We claim that
$\bar{G}^{(i,j)}(\xi,\eta)=G^{(i,j)}(X,Y).$
To this end, by conformal covariance (1.3) and (1.4) of $P_{4}^{g}$ and $\mathscr{B}_{k}^{3}$ we have
$\Delta^{2}(U\circ F)=e^{4\hat{U}_{0}}P_{4}^{e^{2\hat{U}_{0}}|\mathrm{d}X|^{2}}(U\circ F)=e^{4\hat{U}_{0}}[P_{4}^{|\mathrm{d}\xi|^{2}}(U)]\circ F=e^{4\hat{U}_{0}}\bar{f}\circ F$
and for $k\in\{0,1,2,3\}$,
$(\mathscr{B}_{k}^{3})_{|\mathrm{d}X|^{2}}(U\circ F)=e^{k\hat{U}_{0}}(\mathscr{B}_{k}^{3})_{e^{2\hat{U}_{0}}|\mathrm{d}X|^{2}}(U\circ F)=e^{k\hat{U}_{0}}[(\mathscr{B}_{k}^{3})_{|\mathrm{d}\xi|^{2}}U]\circ F=0.$
Hence, we are able to apply the Green formula on +$\mathbb{R}^{4}_{+}$ for $U\circ F$ to obtain
$U(\xi)=U\circ F(X)=$ $\int_{\mathbb{R}_{+}^{4}}G^{(i,j)}(X,Y)e^{4\hat{U}_{0}(Y)}\bar{f}\circ F(Y)\mathrm{d}Y$
$=$ $\int_{\mathbb{B}^{4}}G^{(i,j)}(F^{-1}(\xi),F^{-1}(\eta))\bar{f}(\eta)\mathrm{d}\eta,$
which in turn gives
$\bar{G}^{(i,j)}(\xi,\eta)=G^{(i,j)}(F^{-1}(\xi),F^{-1}(\eta)).$
As before, we let
$H^{(i,j)}(X,Y)=G^{(i,j)}(X,Y)-\Gamma(X-Y)$
satisfy
$\begin{cases}\displaystyle\Delta^{2}H^{(i,j)}=0\qquad&\mathrm{in}\qquad\mathbb{R}^{4}_{+},\\ \displaystyle\mathscr{B}^{3}_{i}(H^{(i,j)})=-\mathscr{B}^{3}_{i}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}^{4}_{+},\\ \displaystyle\mathscr{B}^{3}_{j}(H^{(i,j)})=-\mathscr{B}^{3}_{j}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}^{4}_{+}.\end{cases}$
For $(i,j)=(0,2)$, we find
$H^{(0,2)}(X,Y)=-\Gamma(\bar{X}-Y)$
whence,
$G^{(0,2)}(X,Y)=\Gamma(X-Y)-\Gamma(\bar{X}-Y).$
This together with (3.8), (4.3) gives
$\bar{G}^{(0,2)}(\xi,\eta)=\Gamma(\xi-\eta)-\Gamma(|\xi||\xi^{*}-\eta|).$
We introduce
$\bar{G}^{(1,3)}(\xi,\eta)=\Gamma(\xi-\eta)+\Gamma(|\xi||\xi^{*}-\eta|)+C,\quad C\in\mathbb{R}.$
For $(i,j)=(0,1),(2,3)$, again the Boggio formula gives
$\bar{G}^{(0,1)}(\xi,\eta)=$ $\frac{1}{4|\mathbb{S}^{3}|}\int_{1}^{\frac{|\xi||\xi^{*}-\eta|}{|\xi-\eta|}}(s^{2}-1)s^{-3}\mathrm{d}s$
$=$ $\Gamma(\xi-\eta)-\Gamma(|\xi||\xi^{*}-\eta|)+\frac{1}{8|\mathbb{S}^{3}|}\left(\frac{|\xi-\eta|^{2}}{|\xi|^{2}|\xi^{*}-\eta|^{2}}-1\right).$
Notice that
$H^{(2,3)}(X,Y)=$ $\frac{1}{2}\int_{\mathbb{R}^{3}}P^{3}_{2}(y-z,s)P^{3}_{1}(x-z,t)\mathrm{d}z+\frac{1}{2}\int_{\mathbb{R}^{3}}(P^{3}_{3}(y-z,s)+c)P^{3}_{0}(x-z,t)\mathrm{d}z$
$=$ $\frac{1}{2}\int_{\mathbb{R}^{3}}P^{3}_{2}(y-z,s)P^{3}_{1}(x-z,t)\mathrm{d}z+\frac{1}{2}\int_{\mathbb{R}^{3}}P^{3}_{3}(y-z,s)P^{3}_{0}(x-z,t)\mathrm{d}z+\frac{C}{2},$
then
$H^{(2,3)}(X,Y)=-H^{(0,1)}(X,Y)+\frac{C}{2}.$
Consequently, we have
$\bar{G}^{(2,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\Gamma(|\xi||\xi^{*}-\eta|)-\frac{1}{8|\mathbb{S}^{3}|}\left[\frac{|\xi-\eta|^{2}}{|\xi|^{2}|\xi^{*}-\eta|^{2}}-1\right]+C.$
$\Box$
Eventually, combining the biharmonic Poisson kernel and Green function, we are in a position to present a Green formula for an nonhomogeneous biharmonic boundary value problem on $\mathbb{B}^{n+1}$.
Corollary 4.1 (Green formula for a biharmonic boundary value problem).
For $i,j\in\{0,1,2,3\}$ and $i+j\neq 3$, suppose $\bar{f}\in C^{\infty}(\overline{\mathbb{B}^{n+1}}),\bar{f}_{i},\bar{f}_{j}\in C^{\infty}(\mathbb{S}^{n})$. Then the unique solution of
$\begin{cases}\Delta^{2}U=\bar{f}\qquad&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{B}^{n+1},\\ \mathscr{B}_{i}^{3}U=\bar{f}_{i}&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{S}^{n},\\ \mathscr{B}_{j}^{3}U=\bar{f}_{j}&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{S}^{n},\end{cases}$
has the form of
$U(\xi)=[\bar{G}^{(i,j)}\ast\bar{f}+\bar{P}_{i}^{3}\ast\bar{f}_{i}+\bar{P}_{j}^{3}\ast\bar{f}_{j}](\xi).$
As an application of the Green formula, we are able to obtain the following comparison principles for biharmonic boundary value problems on $\mathbb{B}^{n+1}$.
Proposition 4.1.
Suppose $n\geq 4$ and $U$ is a smooth function in $\overline{\mathbb{B}^{n+1}}$ satisfying
$\begin{cases}\displaystyle\quad\Delta^{2}U\geq 0\qquad&\mathrm{~{}~{}in~{}~{}}\qquad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(U)\geq 0\qquad&\mathrm{~{}~{}on~{}~{}}\qquad\mathbb{S}^{n},\\ \displaystyle\mathscr{B}^{3}_{j}(U)\leq 0\qquad&\mathrm{~{}~{}on~{}~{}}\qquad\mathbb{S}^{n},\end{cases}$
for $i\in\{0,3\},j\in\{1,2\}$, then $U\geq 0$ in $\mathbb{B}^{n+1}$.
Proof..
This is a direct consequent of Corollary 4.1, Theorem 1.2 and Theorem 1.3. $\Box$
5 Classification theorems
As above, let $P_{k}(X)$ and $P_{k}(x)$ denote a polynomial with degree $k$ in the variables $X$ and $x$, where $k\in\mathbb{N}$.
5.1 Proof of Theorem 1.6
Lemma 5.1.
Suppose $u$ is a nonnegative smooth function in $\overline{\mathbb{R}_{+}^{n+1}}$ satisfying
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}_{3}^{3}(u)=\mathbb{T}^{3}_{3}u^{p_{3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\end{cases}$
for $0<p_{3}\leq p_{3}^{\ast}=\frac{n+3}{n-3}$. Assume
$u(x,0)=O(|x|^{-c}),\qquad\partial_{t}u(x,0)=O(|x|^{-d})\qquad\mathrm{~{}~{}as~{}~{}}|x|\to\infty$
for some constants $c>3/p_{3}$ and $d>1$. Then we have
  • (i)
    If $p_{3}=p_{3}^{\ast}$, then there exist $\varepsilon>0$ and $x_{0}\in\mathbb{R}^{n}$ such that
    $u(x,0)=\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}}.$
  • (ii)
    If $0<p_{3}<p_{3}^{\ast}$, then $u(x,0)=0$.
Proof..
Notice that $f_{1}(x):=-\partial_{t}u(x,0)=O(|x|^{-d})$ and $f_{3}(x):=\mathbb{T}^{3}_{3}u^{p_{3}}(x,0)=O(|x|^{-p_{3}c})$ as $|x|\to\infty$. We let
$v(x,t)=-\frac{2}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}f_{1}(y)}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\mathrm{d}y+\frac{1}{(n-3)(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{f_{3}(y)}{(t^{2}+|x-y|^{2})^{\frac{n-3}{2}}}.$
By Lemmas 2.1 and 2.2, as $|X|\to\infty$ we have
$|v(x,t)|\lesssim t\cdot\begin{cases}|X|^{-\min\{d,n\}}&\mathrm{if~{}~{}}d\neq n\\ |X|^{-n}\log|X|&\mathrm{if~{}~{}}d=n\end{cases}+\begin{cases}|X|^{3-\min\{p_{3}c,n\}}&\mathrm{if~{}~{}}p_{3}c\neq n\\ |X|^{3-n}\log|X|&\mathrm{if~{}~{}}p_{3}c=n\end{cases}\to 0.$ (5.1)
By Theorem 1.1, $w:=u-v$ solves
$\begin{cases}\displaystyle\Delta^{2}w=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\partial_{t}w(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta w(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
Moreover, since $u$ is nonnegative, by (5.1) we have
$w(x,t)\geq-C|X|^{-\delta}\qquad\mathrm{~{}~{}as~{}~{}}|X|\to\infty,$
where $\delta=\min\{\min\{d,n\}-1,\min\{p_{3}c,n\}-3\}/2>0$, and
$w(x,0)=$ $u(x,0)-\frac{1}{(n-3)(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{f_{3}(y)}{|x-y|^{n-3}}\mathrm{d}y$
$=$ $O(|x|^{\frac{3-\min\{p_{3}c,n\}}{2}})+O(|x|^{-c})\to 0,\qquad\quad\mathrm{~{}~{}as~{}~{}}|x|\to\infty.$
Then it follows from Lemma 2.3 that $w(x,t)=c_{2}^{\prime}t^{2}$ for some nonnegative constant $c_{2}^{\prime}$. This means
$u(x,t)=\frac{n+1}{4|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u^{p_{3}}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n-3}{2}}}\mathrm{d}y+\frac{2}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}\partial_{t}u(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\mathrm{d}y+c_{2}^{\prime}t^{2}.$ (5.2)
Letting $t\to 0$ we obtain
$u(x,0)=\frac{n+1}{4|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u^{p_{3}}(y,0)}{|x-y|^{n-3}}\mathrm{d}y.$
The classification theorem (see Chen-Li-Ou [9] and Y. Y. Li [24]) for the integral equation above gives
$u(x,0)=\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}},\quad\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}\qquad\qquad\mathrm{~{}~{}for~{}~{}}p_{3}=p_{3}^{\ast}$
and
$u(x,0)=0\qquad\qquad\mathrm{~{}~{}for~{}~{}}0<p_{3}<p_{3}^{\ast}.$
$\Box$
Proof of Theorem 1.6. For $c>\max\{3/p_{3},1/p_{1}\}$, $u(x,0)=O(|x|^{-c})$ and $\partial_{t}u(x,0)=-\frac{n-3}{2}u^{p_{1}}(x,0)=O(|x|^{-p_{1}c})$ as $|x|\to\infty$.
(1) If $p_{3}=p_{3}^{\ast}=\frac{n+3}{n-3}$, then by Lemma 5.1 we have
$u(x,0)=\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}}.$
Now $u$ solves
$\begin{cases}\displaystyle\Delta^{2}u=0&\mathrm{in}\quad\mathbb{R}_{+}^{n+1},\\ \displaystyle\partial_{t}u(x,0)=-\frac{n-3}{2}\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{(n-3)p_{1}}{2}}&\mathrm{on}\quad\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}u=\frac{(n^{2}-1)(n-3)}{4}\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{n+3}{2}}\quad&\mathrm{on}\quad\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (5.3)
By (5.2) we know
$u(x,t)=M_{1}(x,t)+c_{2}^{\prime}t^{2}.$ (5.4)
where $c_{2}^{\prime}$ is a nonnegative constant and
$M_{1}(x,t)=$ $\frac{n+1}{4|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{1}{(t^{2}+|x-y|^{2})^{\frac{n-3}{2}}}\left(\frac{2\varepsilon}{\varepsilon^{2}+|y-x_{0}|^{2}}\right)^{\frac{n+3}{2}}\mathrm{d}y$
$-\frac{n-3}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\left(\frac{2\varepsilon}{\varepsilon^{2}+|y-x_{0}|^{2}}\right)^{\frac{(n-3)p_{1}}{2}}\mathrm{d}y.$
Especially, assume $p_{1}=p_{1}^{\ast}=\frac{n-1}{n-3}$. On one hand, by Lemmas 2.1 and 2.2 we have
$M_{1}(x,t)=O(|X|^{3-n})\qquad\mathrm{~{}~{}as~{}~{}}|X|\to\infty.$ (5.5)
On the other hand, we consider
$w(x,t)=u(x,t)-\left(\frac{2\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}},$
then $w$ satisfies
$\begin{cases}\displaystyle\Delta^{2}w=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\partial_{t}w(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta w(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\end{cases}$ (5.6)
and for $|X|,|x|\gg 1$,
$w(X)\geq-C|X|^{3-n}\qquad\mathrm{and}\qquad w(x,0)=O(|x|^{-\min\{c,n-3\}}).$
Hence, it follows from Lemma 2.3 that the solution of (5.3) has the form of
$u(x,t)=\left(\frac{2\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}}+c_{2}t^{2}.$ (5.7)
Eventually, we combine (5.4), (5.5) and (5.7) to conclude that $c_{2}^{\prime}=c_{2}$.
(2) If $p_{3}<p_{3}^{\ast}$, then again by Lemma 5.1 we obtain $u(x,0)=0$. In this case, $u$ also solves the same homogeneous biharmonic boundary value problem (5.6), thanks to Lemma 2.3, we conclude that $u(x,t)=c_{2}t^{2}$ for every nonnegative constant $c_{2}$. $\Box$
5.2 Proof of Theorem 1.7
Proof of Theorem 1.7. For $cp_{2}>2$, $f_{0}(x)=u(x,0)=O(|x|^{-c})$ and $f_{2}(x)=\mathbb{T}^{3}_{2}u^{p_{2}}=O(|x|^{-p_{2}c})$ as $|x|\to\infty$. Let
$v(x,t)=\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}f_{0}(y)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y-\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{tf_{2}(y)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y.$
with
$|v(x,t)|\lesssim\begin{cases}|X|^{2-\min\{p_{2}c,n\}}&\mathrm{if~{}~{}}p_{2}c\neq n\\ |X|^{2-n}\log|X|&\mathrm{if~{}~{}}p_{2}c=n\end{cases}+\begin{cases}|X|^{-\min\{c,n\}}&\mathrm{if~{}~{}}p_{2}c\neq n\\ |X|^{-n}\log|X|&\mathrm{if~{}~{}}p_{2}c=n\end{cases}\to 0$ (5.8)
as $|X|\to\infty$ by virtue of Lemmas 2.1 and 2.2.
By Theorem 1.1, $w:=u-v$ satisfies
$\begin{cases}\displaystyle\Delta^{2}w=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle w(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\Delta w(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
Moreover, by (5.8) we have
$w(x,t)\geq-C|X|^{-\delta}\qquad\mathrm{as}\qquad|X|\to\infty,$
where $\delta=\min\left\{\min\{p_{2}c,n\}-2,\min\{c,n\}\right\}/2>0$. By Lemma 2.4 we have
$u(x,t)=$ $\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}u(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y-\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{tu^{p_{2}}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y$
$+tP_{2}(x)+c_{3}t^{3},$ (5.9)
where $c_{3}$ is a nonnegative constant.
From Lemma 3.1 that
$\lim_{t\to 0}\partial_{t}\int_{\mathbb{R}^{n}}\frac{t^{3}}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}u(y,0)\mathrm{d}y=0,$
and the first boundary condition we obtain
$\frac{n-3}{2}u^{p_{1}}(x,0)=-\partial_{t}u(x,0)=\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u^{p_{2}}(y,0)}{|x-y|^{n-1}}\mathrm{d}y-P_{2}(x).$ (5.10)
Since $u(x,0)=o(1)$ as $|x|\to\infty$, by Lemma 2.1 we have
$P_{2}(x)=o(1)\qquad\mathrm{as}\qquad|x|\to+\infty,$
so $P_{2}(x)=0$.
Write $k(x)=u^{p_{1}}(x,0)$, the integral equation (5.10) becomes
$k(x)=\frac{1}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{k^{\frac{p_{2}}{p_{1}}}(y)}{|x-y|^{n-1}}\mathrm{d}y.$
(1) If $\frac{p_{2}}{p_{1}}=\frac{n+1}{n-1}$, then we apply the classification theorem (see [9] and [24]) to know
$k(x)=\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-1}{2}},\qquad\varepsilon\in\mathbb{R}_{+},~{}x_{0}\in\mathbb{R}^{n}.$
Meanwhile, the second boundary condition becomes
$\mathscr{B}^{3}_{2}u=\frac{(n-1)(n-3)}{2}\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{n+1}{2}}.$
By (5.2) we have
$u(x,t)=M_{2}(x,t)+c_{3}t^{3}$
and
$M_{2}(x,t)=$ $\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\left(\frac{2\varepsilon}{\varepsilon^{2}+|y-x_{0}|^{2}}\right)^{\frac{n-1}{2p_{1}}}\mathrm{d}y$
$-\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\left(\frac{2\varepsilon}{\varepsilon^{2}+|y-x_{0}|^{2}}\right)^{\frac{n+1}{2}}\mathrm{d}y.$
Especially, when $p_{1}=p_{1}^{\ast}=\frac{n-1}{n-3}$, we consider
$w(x,t)=u(x,t)-\left(\frac{2\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}},$
which satisfies
$\begin{cases}\displaystyle\Delta^{2}w=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle w(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\Delta w(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\end{cases}$ (5.11)
with
$w(X)\geq-C|X|^{3-n}\qquad\mathrm{as~{}~{}}|X|\to\infty.$
Moreover, by Lemmas 2.1 and 2.2 we have
$M_{2}(x,t)=O(|X|^{-\min\{\frac{n-1}{p_{1}},n\}})+O(|X|^{2-n})\to 0\qquad\mathrm{as~{}~{}}|X|\to\infty.$
Hence, by Lemma 2.4 we obtain
$u(x,t)=\left(\frac{2\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}}+c_{3}t^{3}.$
(2) If $\frac{p_{2}}{p_{1}}<\frac{n-1}{n-3}$, then the uniqueness theorem in Y. Y. Li [24] yields $u(x,0)=0$ in $\mathbb{R}^{n}$, thus $u(x,t)$ satisfies the homogeneous boundary value problem (5.11). This together with Lemma 2.4 yields $u(x,t)=c_{3}t^{3}$ in $\mathbb{R}_{+}^{n+1}$ for every nonnegative constant $c_{3}$. $\Box$
5.3 Proof of Theorem 1.8
Proof of Theorem 1.8. We now consider
$\begin{cases}\displaystyle\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}_{2}^{3}(u)=\frac{(n-1)(n-3)}{2}u^{p_{2}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}(u)=\frac{(n^{2}-1)(n-3)}{4}u^{p_{3}}&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (5.12)
As $cp_{2}>2$, a similar argument as in Theorem 1.7 yields
$u(x,t)=$ $\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}u(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y-\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{tu^{p_{2}}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y$
$+tP_{2}(x)+c_{3}t^{3},$ (5.13)
where $c_{3}$ is a nonnegative constant and
$\liminf_{|x|\to\infty}P_{2}(x)\geq 0.$ (5.14)
Moreover,
$-\partial_{t}u(x,0)=\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u^{p_{2}}(y,0)}{|x-y|^{n-1}}\mathrm{d}y-P_{2}(x).$
We first claim that
$P_{2}(x)\equiv c\geq 0\qquad\mathrm{and}\qquad c_{3}=0.$
It is known from (5.14) that $\overline{\Delta}P_{2}(x)$ is a nonnegative constant. With $\tilde{c}_{3}:=c_{3}+\frac{1}{2}\overline{\Delta}P_{2}(x)$ we define
$\tilde{u}(x,t)=$ $u(x,t)-P_{2}(x)t-c_{3}t^{3}+\tilde{c}_{3}t^{3}$
$=$ $\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}u(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y-\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{tu^{p_{2}}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y+\tilde{c}_{3}t^{3}.$ (5.15)
Clearly, $\tilde{u}$ is biharmonic in $\mathbb{R}_{+}^{n+1}$ and $\tilde{u}(x,0)=u(x,0)$. Recall that $\mathscr{B}^{3}_{2}=\partial_{t}^{2}-\overline{\Delta}$ and $\mathscr{B}^{3}_{3}=\partial_{t}\Delta+2\overline{\Delta}\partial_{t}$, then
$\mathscr{B}^{3}_{3}(\tilde{u})=\mathscr{B}^{3}_{3}(u)+6(\tilde{c}_{3}-c_{3})-3\overline{\Delta}P_{2}(x)=\mathscr{B}^{3}_{3}(u)$
and $\mathscr{B}^{3}_{2}(\tilde{u})=\mathscr{B}^{3}_{2}(u)$ on $\partial\mathbb{R}_{+}^{n+1}$. This means that $\tilde{u}$ also satisfies the biharmonic boundary value problem (5.12), though it might change sign.
On the other hand, notice that $\tilde{u}(x,t)\geq-C|X|^{3}$ as $|X|\to\infty$. By Lemma 2.1 as $|x|\to\infty$ we have
$-\partial_{t}\tilde{u}(x,0)=\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{\tilde{u}^{p_{2}}(y,0)}{|x-y|^{n-1}}\mathrm{d}y=$$\begin{cases}O(|x|^{1-\min\{p_{2}c,n\}})&\qquad\mathrm{if~{}~{}}p_{2}c\neq n\\ O(|x|^{1-n}\log|x|)&\qquad\mathrm{if~{}~{}}p_{2}c=n\end{cases}$
$=$ $O(|x|^{-d}),$
where $d=\min\{cp_{2}-1,(n-1)/2\}>1$. For $c>\max\{2/p_{2},3/p_{3}\}$, we apply Remark 2.1 to obtain
$\tilde{u}(x,t)=\frac{2}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}\partial_{t}\tilde{u}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\mathrm{d}y+\frac{n+1}{4|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u^{p_{3}}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n-3}{2}}}\mathrm{d}y+t^{2}P_{1}(x).$ (5.16)
By (5.3) and (5.16), we can apply Lemmas 2.1 and 2.2 to conclude that
$\tilde{c}_{3}t^{3}-t^{2}P_{1}(x)=o(1)\qquad\mathrm{~{}~{}as~{}~{}}|X|\to\infty.$
This implies $\tilde{c}_{3}=0$ and $P_{1}(x)\equiv 0$. Then, (5.16) becomes
$\tilde{u}(x,t)=\frac{2}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}\partial_{t}\tilde{u}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\mathrm{d}y+\frac{n+1}{4|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u^{p_{3}}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n-3}{2}}}\mathrm{d}y.$ (5.17)
Keep in mind that
$0=\tilde{c}_{3}=c_{3}+\frac{\overline{\Delta}P_{2}(x)}{2},$
so, $c_{3}=0$ and $\overline{\Delta}P_{2}(x)=0$. This together with (5.14) implies $P_{2}(x)\equiv c_{1}\geq 0$. Thus, we prove the above claim.
Going back to (5.3) and (5.3) we now have
$u(x,t)=$ $\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}u(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y-\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{tu^{p_{2}}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y$
$+c_{1}t$ (5.18)
and
$\tilde{u}(x,t)=u(x,t)-c_{1}t\geq-C|X|\qquad\mathrm{as~{}~{}}|X|\to\infty.$ (5.19)
Furthermore, letting $t\to 0$ in (5.17) we arrive at
$u(x,0)=\tilde{u}(x,0)=\frac{n+1}{4|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u^{p_{3}}(y,0)}{|x-y|^{n-3}}\mathrm{d}y.$
(1) If $p_{3}=\frac{n+3}{n-3}$ and $p_{2}>\frac{2}{n-3}$, then the classification theorem in [9] and [24] gives
$u(x,0)=\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}},\quad x_{0}\in\mathbb{R}^{n},\varepsilon\in\mathbb{R}_{+}.$
This together with (5.3) yields
$u(x,t)=M_{3}(x,t)+c_{1}t,$
where
$M_{3}(x,t)=$ $\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\left(\frac{2\varepsilon}{\varepsilon^{2}+|y-x_{0}|^{2}}\right)^{\frac{n-3}{2}}\mathrm{d}y$
$-\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\left(\frac{2\varepsilon}{\varepsilon^{2}+|y-x_{0}|^{2}}\right)^{\frac{(n-3)p_{2}}{2}}\mathrm{d}y.$
In particular, when $p_{2}=p_{2}^{\ast}$, as before, we notice that $(\tilde{u}-U_{\varepsilon,x_{0}})(X)\geq-C|X|$ as $|X|\to\infty$, and apply Remark 2.2 to conclude that $M_{3}=U_{\varepsilon,x_{0}}$ and then
$u(x,t)=U_{\varepsilon,x_{0}}(x,t)+c_{1}t.$
(2) If $p_{3}<\frac{n+3}{n-3}$, then the uniqueness theorem in [24] implies $u(x,0)=0$. This together with (5.3) implies $u(x,t)=c_{1}t$ in $\mathbb{R}_{+}^{n+1}$ for every nonnegative constant $c_{1}$. $\Box$
6 Applications to biharmonic boundary value problems on balls
Although we focus on the conformal geometry nature for biharmonic boundary value problems, we would like to mention that from a physical point of view, the boundary conditions $(\mathscr{B}_{0}^{3},\mathscr{B}_{1}^{3})$ and $(\mathscr{B}_{0}^{3},\mathscr{B}_{2}^{3})$ have a close connection with the clamped plate and hinged plate model (cf. [15]), respectively. However, the study of other boundary conditions is far less developed.
6.1 One boundary point singularity
Proof of Theorem 1.9. It suffices to give a proof for $i=1$ and $j=3$, since the others are the same in the spirit due to Theorems 1.7 and 1.8.
Let $U$ be a nonnegative solution of (1.9). As before, we introduce
$u(X)=\left(\frac{2}{|X+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}U\circ F.$
Then $u$ solves
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(u)=\mathbb{T}^{3}_{1}u^{\frac{n-1}{n-3}}&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{j}(u)=\mathbb{T}^{3}_{3}u^{\frac{n+3}{n-3}}&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}_{+}^{n+1}.\end{cases}$
By definition of the spherical inversion $F$, we know that $|X+\mathbf{e}_{n+1}||\xi+\mathbf{e}_{n+1}|=2$ and $F(\partial\mathbb{R}_{+}^{n+1})=\partial\mathbb{B}^{n+1}$. By the assumption (1.10) for $U$, we obtain
$u(x,0)=\left(\frac{2}{1+|x|^{2}}\right)^{\frac{n-3}{2}}U\circ F(x,0)\lesssim|x|^{3-n}|X+\mathbf{e}_{n+1}|^{\alpha}\lesssim|x|^{-(n-3-\alpha)},\quad\mathrm{as~{}~{}}|x|\to\infty$
with $n-3-\alpha>\frac{3(n-3)}{n+3}$. By Theorem 1.6, we have
$u(x,t)=U_{x_{0},\varepsilon}(x,t)+c_{2}t^{2}$
for some $x_{0}\in\mathbb{R}^{n},\varepsilon\in\mathbb{R}_{+}$ and nonnegative constant $c_{2}$.
If we let $\xi_{0}=F(x_{0},\varepsilon)\in\mathbb{B}^{n+1}$, then it is not hard to verify that
$F^{\ast}(U_{\xi_{0}}(\xi)^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2})=U_{x_{0},\varepsilon}(X)^{\frac{4}{n-3}}|\mathrm{d}X|^{2},$
where
$U_{\xi_{0}}(\xi)=\left(\frac{1-|\xi_{0}|^{2}}{|\xi|^{2}|\xi_{0}|^{2}-2\xi_{0}\cdot\xi+1}\right)^{\frac{n-3}{2}},\qquad\xi\in\mathbb{B}^{n+1}.$
In other words,
$U_{\xi_{0}}(\xi)=\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}U_{x_{0},\varepsilon}\circ F^{-1}(\xi).$
Using
$t\circ F^{-1}=\frac{1-|\xi|^{2}}{|\xi+\mathbf{e}_{n+1}|^{2}}$
and (6.1) we obtain
$U(\xi)=$ $\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}u\circ F^{-1}(\xi)$
$=$ $U_{\xi_{0}}(\xi)+c_{2}\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\left(\frac{1-|\xi|^{2}}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{2}.$
This follows the desired assertion. $\Box$
6.2 Sharp geometric inequalities
On a compact manifold with boundary, the maximizing problem of the isometric ratio over a scalar-flat conformal class was studied in [22, 21]. With comparison principles in Proposition 4.1, we are motivated to study of the maximizers of the isometric ratio and another sharp geometric inequality over some $Q$-curvature flat conformal class on $(\overline{\mathbb{B}^{n+1}},|\mathrm{d}\xi|^{2})$.
More general, for $n\geq 4$ we define
$\mathcal{M}_{1}=$$\{g\in[|\mathrm{d}\xi|^{2}]~{}|~{}Q_{g}\leq 0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}(T_{2}^{3})_{g}\geq 0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\};$
$\mathcal{M}_{2}=$$\{g\in[|\mathrm{d}\xi|^{2}]~{}|~{}Q_{g}\leq 0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}h_{g}\geq 0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\}$
and
$\mathring{\mathcal{M}}:=$$\{g\in[|\mathrm{d}\xi|^{2}]~{}|~{}Q_{g}=0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}(T_{2}^{3})_{g}=0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\}$
$=$ $\{g\in[|\mathrm{d}\xi|^{2}]~{}|~{}Q_{g}=0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}h_{g}=0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\}$
by virtue of Proposition 3.1.
Lemma 6.1.
Let $g=U^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2}$ be a smooth metric in $\overline{\mathbb{B}^{n+1}}$. Suppose $g\in\mathcal{M}_{1}\cup\mathcal{M}_{2}$. Then we have
  • (1)
    there exists a conformal metric $\tilde{g}=\tilde{U}^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2}\in\mathring{\mathcal{M}}$ such that $\tilde{U}\geq U$ in $\mathbb{B}^{n+1}$ with $\tilde{U}=U$ on $\mathbb{S}^{n}$;
  • (2)
    there exists a conformal metric $\tilde{g}=\tilde{U}^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2}\in\mathring{\mathcal{M}}$ such that $\tilde{U}\geq U$ in $\mathbb{B}^{n+1}$ with $\mathscr{B}^{3}_{3}(U)=\mathscr{B}^{3}_{3}(\tilde{U})$ on $\mathbb{S}^{n}$.
Proof..
For (i), we introduce an auxiliary function by
$\tilde{U}(\xi)=$ $\frac{(n+1)\left(1-|\xi|^{2}\right)^{3}}{4|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{U(\eta)}{|\xi-\eta|^{n+3}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta).$
By Theorem 1.2 we have
$\begin{cases}\displaystyle\quad\Delta^{2}U\leq 0=\Delta^{2}\tilde{U}\qquad&\mathrm{in}\qquad\mathbb{B}^{n+1},\\ \displaystyle\qquad~{}U=\tilde{U}\qquad&\mathrm{on}\qquad\mathbb{S}^{n},\\ \displaystyle\mathscr{B}^{3}_{i}(U)\geq 0=\mathscr{B}^{3}_{i}(\tilde{U})\qquad&\mathrm{on}\qquad\mathbb{S}^{n},\end{cases}$
for $i\in\{1,2\}$. Then the desired assertion follows by Proposition 4.1.
For (ii), we consider
$\tilde{U}(\xi)=\frac{1}{(n-1)(n-3)|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\mathscr{B}^{3}_{3}(U)(\eta)}{|\xi-\eta|^{n-3}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta).$
By Theorem 1.2 we have
$\begin{cases}\displaystyle\quad\Delta^{2}U\leq 0=\Delta^{2}\tilde{U}\qquad&\mathrm{in}\qquad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(U)\geq 0=\mathscr{B}^{3}_{i}(\tilde{U})\qquad&\mathrm{on}\qquad\mathbb{S}^{n},\\ \displaystyle\mathscr{B}^{3}_{3}(U)=\mathscr{B}^{3}_{3}(\tilde{U})\qquad&\mathrm{on}\qquad\mathbb{S}^{n},\end{cases}$
for $i\in\{1,2\}$. Then the desired assertion follows from Proposition 4.1. $\Box$
In the following, we consider the isoperimetric ratio over $\mathcal{M}_{1}\cup\mathcal{M}_{2}$:
$\sup_{g\in\mathcal{M}_{1}\cup\mathcal{M}_{2}}\frac{|\mathbb{B}^{n+1}|_{g}}{|\mathbb{S}^{n}|_{g}^{\frac{n+1}{n}}}=\sup_{g\in\mathring{\mathcal{M}}}\frac{|\mathbb{B}^{n+1}|_{g}}{|\mathbb{S}^{n}|_{g}^{\frac{n+1}{n}}}$ (6.1)
by virtue of Lemma 6.1. Given each $g=U^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2}\in\mathring{\mathcal{M}}$, notice that
$\int_{\mathbb{S}^{n}}|(T_{3}^{3})_{g}|^{\frac{2n}{n+3}}\mathrm{d}V_{g}=\left(\frac{2}{n-3}\right)^{\frac{2n}{n+3}}\int_{\mathbb{S}^{n}}|\mathscr{B}^{3}_{3}(u)|^{\frac{2n}{n+3}}\mathrm{d}V_{\mathbb{S}^{n}}.$
then it follows from Theorem 1.2 that $(T_{3}^{3})_{g}$ does not vanish identically. So, again by Lemma 6.1 it is reasonable to study
$\sup_{g\in\mathcal{M}_{1}\cup\mathcal{M}_{2}}\frac{|\mathbb{B}^{n+1}|^{\frac{n-3}{2(n+1)}}_{g}}{\|(T_{3}^{3})_{g}\|_{L^{\frac{2n}{n+3}}(\mathbb{S}^{n},g)}}=\sup_{g\in\mathring{\mathcal{M}}}\frac{|\mathbb{B}^{n+1}|^{\frac{n-3}{2(n+1)}}_{g}}{\|(T_{3}^{3})_{g}\|_{L^{\frac{2n}{n+3}}(\mathbb{S}^{n},g)}}.$ (6.2)
Theorem 6.1.
Let $n\geq 4$. Then for all $g\in\mathcal{M}_{1}\cup\mathcal{M}_{2}$, there hold
$|\mathbb{B}^{n+1}|_{g}\leq d_{n}|\mathbb{S}^{n}|_{g}^{\frac{n+1}{n}}\quad\mathrm{~{}~{}and~{}~{}}\quad|\mathbb{B}^{n+1}|^{\frac{n-3}{2(n+1)}}_{g}\leq e_{n}\|(T_{3}^{3})_{g}\|_{L^{\frac{2n}{n+3}}(\mathbb{S}^{n},g)},$
where
$d_{n}=|\mathbb{S}^{n}|^{-\frac{1}{n}}\int_{0}^{1}r^{n}\big{(}1+\frac{n-3}{4}(1-r^{2})\big{)}^{\frac{2(n+1)}{n-3}}\mathrm{d}r$
and
$e_{n}=|\mathbb{S}^{n}|^{-\frac{7n+3}{2n(n+1)}}\frac{4}{(n^{2}-1)(n-3)}\left[\int_{0}^{1}r^{n}\big{(}1+\frac{n-3}{4}(1-r^{2})\big{)}^{\frac{2(n+1)}{n-3}}\mathrm{d}r\right]^{\frac{n-3}{2(n+1)}}.$
Moreover, the above two equalities hold if and only if $g=cU_{a}^{4/(n-3)}|\mathrm{d}\xi|^{2}\in\mathring{\mathcal{M}}$ with $c\in\mathbb{R}_{+}$ and
$U_{a}(\xi)=\Big{(}\frac{1-|a|^{2}}{|a|^{2}|\xi|^{2}-2a\cdot\xi+1}\Big{)}^{\frac{n-3}{2}}+\frac{n-3}{4}(1-|\xi|^{2})\Big{(}\frac{1-|a|^{2}}{|a|^{2}|\xi|^{2}-2a\cdot\xi+1}\Big{)}^{\frac{n-1}{2}}$ (6.3)
for every $a\in\mathbb{B}^{n+1}$.
Proof..
Given a conformal metric $g=U^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2}\in\mathring{\mathcal{M}}$, if we let
$u(x)=\left(\frac{2}{1+|x|^{2}}\right)^{\frac{n-3}{2}}U\circ F(x,0),$
then a similar argument in the proof of Theorem 1.9 yields
$\left(\frac{2}{(1+t)^{2}+|x|^{2}}\right)^{\frac{n-3}{2}}U\circ F(x,t)=\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}u(y)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y.$
To be consistent with notation in Gluck [17], we define
$E_{-2,3}(u)(x,t):=\left(\frac{2}{(1+t)^{2}+|x|^{2}}\right)^{\frac{n-3}{2}}U\circ F(x,t).$
So the extremal problem (6.1) is equivalent to
$d_{n}:=\sup_{\begin{subarray}{c}0<u\in C^{\infty}(\mathbb{R}^{n})\\ u=O(|x|^{3-n})\mathrm{~{}as~{}}|x|\to\infty\end{subarray}}\frac{\|E_{-2,3}(u)\|_{L^{\frac{2(n+1)}{n-3}}(\mathbb{R}^{n+1}_{+})}}{\|u\|_{L^{\frac{2n}{n-3}}(\mathbb{R}^{n})}}.$ (6.4)
Gluck [17, Theorem 1.1] states that the optimal constant $d_{n}$ for the extremal problem (6.4) is attained by
$c(|x-x_{0}|^{2}+d^{2})^{\frac{3-n}{2}},\qquad c,d\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n},$
equivalently, modulo of a positive constant, the extremal metric on $\mathbb{S}^{n}$ is a conformal metric $u_{a}^{4/(n-3)}g_{\mathbb{S}^{n}}$ induced by some conformal transformation on $\mathbb{S}^{n}$, which can be expressed as
$u_{a}(\xi)=\Big{(}\frac{1-|a|^{2}}{|a|^{2}|-2a\cdot\xi+1}\Big{)}^{\frac{n-3}{2}}\qquad\mathrm{on~{}~{}}\quad\mathbb{S}^{n}$ (6.5)
for every $a\in\mathbb{B}^{n+1}$. Meanwhile, the uniqueness Theorem 1.2 for both $(\mathscr{B}_{0}^{3},\mathscr{B}_{1}^{3})$ and $(\mathscr{B}_{0}^{3},\mathscr{B}_{2}^{3})$ shows that
$U_{a}(\xi)=\frac{(n+1)\left(1-|\xi|^{2}\right)^{3}}{4|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{u_{a}(\eta)}{|\xi-\eta|^{n+3}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta)$
such that $U_{a}^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2}\in\mathring{\mathcal{M}}$ is an extremal metric with $U_{a}=u_{a}$ on $\mathbb{S}^{n}$. Indeed, the trick in authors [11, Section 6] can be applied to derive an explicit representation formula of
$U_{a}(\xi)=\Big{(}\frac{1-|a|^{2}}{|a|^{2}|\xi|^{2}-2a\cdot\xi+1}\Big{)}^{\frac{n-3}{2}}+\frac{n-3}{4}(1-|\xi|^{2})\Big{(}\frac{1-|a|^{2}}{|a|^{2}|\xi|^{2}-2a\cdot\xi+1}\Big{)}^{\frac{n-1}{2}}.$
Moreover, we prefer to compute the exact value of $d_{n}$ in a geometric way: Take $1+\frac{n-3}{4}(1-|\xi|^{2})$ as an extremal function, and then
$d_{n}=$ $\frac{\int_{\mathbb{B}^{n+1}}\left(1+\frac{n-3}{4}(1-|\xi|^{2})\right)^{\frac{2(n+1)}{n-3}}\mathrm{d}\xi}{|\mathbb{S}^{n}|^{\frac{n+1}{n}}}$
$=$ $|\mathbb{S}^{n}|^{-\frac{1}{n}}\int_{0}^{1}r^{n}\left(1+\frac{n-3}{4}(1-r^{2})\right)^{\frac{2(n+1)}{n-3}}\mathrm{d}r.$
Similarly, given a conformal metric $g=U^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2}\in\mathring{\mathcal{M}}$, if we let
$u(x)=\left(\frac{2}{1+|x|^{2}}\right)^{\frac{n+3}{2}}\mathscr{B}^{3}_{3}(U)\circ F(x,0),$ (6.6)
then
$\left(\frac{2}{(1+t)^{2}+|x|^{2}}\right)^{\frac{n-3}{2}}U\circ F(x,t)$
$=$ $\frac{1}{(n-3)(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u(y)}{(t^{2}+|x-y|^{2})^{\frac{n-3}{2}}}\mathrm{d}y.$
To be consistent with the notation in Dou-Zhu [12], we define
$E_{4}(u)(x,t):=\left(\frac{2}{(1+t)^{2}+|x|^{2}}\right)^{\frac{n-3}{2}}U\circ F(x,t).$
So the extremal problem (6.2) is equivalent to
$\sup_{\begin{subarray}{c}0\not\equiv u\in C^{\infty}(\mathbb{R}^{n})\\ u=O(|x|^{3-n})\mathrm{~{}as~{}}|x|\to\infty\end{subarray}}\frac{\|E_{4}(u)\|_{L^{\frac{2(n+1)}{n-3}}(\mathbb{R}^{n+1}_{+})}}{\|u\|_{L^{\frac{2n}{n+3}}(\mathbb{R}^{n})}}.$ (6.7)
Dou-Zhu [12, Theorem 1.4] states that every maximizer for (6.7) has the form of
$c(|x-x_{0}|^{2}+d^{2})^{-\frac{n+3}{2}},\qquad c,d\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}.$
Combining the above with authors [11, Section 6], we conclude from (6.6) that, modulo a positive constant,
$2P_{3}U=\mathscr{B}^{3}_{3}(U)=\frac{n(n-2)(n-4)}{4}u_{a}^{\frac{n+3}{n-3}}\quad\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n},$
where $u_{a}$ is a conformal factor induced by a conformal transformation on $\mathbb{S}^{n}$ as in (6.5). Since $\mathrm{ker}P_{3}=\{0\}$ implies $U=u_{a}$ on $\mathbb{S}^{n}$, the remaining part of proof is identical to that of (1). $\Box$
Appendix A Appendix: Proof of Proposition 3.1
The combination of Theorem 1.2 and [10, Theorem 1(3)] demonstrates that the Green functions of fractional GJMS operators $P_{1}$ and $P_{3}$ are
$G_{P_{1}}(\xi,\eta)=\frac{2}{(n-1)|\mathbb{S}^{n}|}|\xi-\eta|^{1-n},\quad G_{P_{3}}(\xi,\eta)=2\bar{P}_{3}^{3}(\xi,\eta),\qquad\xi,\eta\in\mathbb{S}^{n}.$
Proof of Proposition 3.1. By (3.10) we set $\bar{f}_{i}=\mathscr{B}^{3}_{i}U$ and
$f_{i}(x)=\left(\frac{2}{1+|x|^{2}}\right)^{\frac{n+2i-3}{2}}\bar{f}_{i}\circ F\left((x,0)\right)$
for $i\in\{0,1,2,3\}$.
When $n\neq 3$, by Theorem 1.2 we have
$U(\xi)=-$ $\frac{(1-|\xi|^{2})^{2}}{2|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\bar{f}_{1}(\eta)}{|\xi-\eta|^{n+1}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta)$
$+$ $\frac{1}{(n-1)(n-3)|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\bar{f}_{3}(\eta)}{|\xi-\eta|^{n-3}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta),\qquad\xi\in\mathbb{B}^{n+1}.$
When $n=4$, by Theorem 1.4 we have
$U(\xi)-\frac{1}{|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}U\mathrm{d}V_{{\mathbb{S}^{3}}}$
$=$ $\int_{\mathbb{S}^{3}}\bar{P}^{3}_{1}(\xi-\eta)\bar{f}_{1}(\eta)\mathrm{d}V_{{\mathbb{S}^{3}}}(\eta)-\frac{1}{2|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}\log|\xi-\eta|\bar{f}_{3}(\eta)\mathrm{d}V_{\mathbb{S}^{3}}(\eta),\quad\forall~{}\xi\in\overline{\mathbb{B}^{4}}.$
Using the elementary formula (3.8) and conformal changes we obtain
$\frac{(1-|\xi|^{2})^{2}}{2|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\bar{f}_{1}(\eta)}{|\xi-\eta|^{n+1}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta)=\frac{2}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}f_{1}(y)}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\mathrm{d}y.$
Since $|\xi|\to 1$ means $t\to 0$, we have
$\lim_{|\xi|\to 1}\frac{(1-|\xi|^{2})^{2}}{2|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\bar{f}_{1}(\eta)}{|\xi-\eta|^{n+1}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta)=0.$
Consequently, we arrive at
$\forall~{}\xi\in\mathbb{S}^{n},\quad U(\xi)=\bar{P}_{3}^{3}\ast\bar{f}_{3}(\xi)=\frac{1}{2}G_{P_{3}}\ast\bar{f}_{3}(\xi),\qquad\mathrm{when~{}~{}}n\neq 3$
and
$\forall~{}\xi\in\mathbb{S}^{n},\quad U(\xi)-\frac{1}{|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}U\mathrm{d}V_{\mathbb{S}^{3}}=\bar{P}_{3}^{3}\ast\bar{f}_{3}(\xi)=\frac{1}{2}G_{P_{3}}\ast\bar{f}_{3}(\xi),\qquad\mathrm{when~{}~{}}n=3.$
Consequently, we immediately obtain
$2P_{3}U=\bar{f}_{3}=\mathscr{B}^{3}_{3}U\qquad\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}.$
Similarly, again by Theorem 1.2 we have
$U(\xi)=$ $\frac{(n+1)\left(1-|\xi|^{2}\right)^{3}}{4|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\bar{f}_{0}(\eta)}{|\xi-\eta|^{n+3}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta)$
$-\frac{(1-|\xi|^{2})}{2(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\bar{f}_{2}(\eta)}{|\xi-\eta|^{n-1}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta),\qquad\xi\in\mathbb{B}^{n+1}.$
Connecting a function $u$ with $U$ as in (3.9) we employ (3.8) and conformal changes to show
$u(x,t)=$ $\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}f_{0}(y)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y-\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{tf_{2}(y)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y$
$=$ $[P^{3}_{0}\ast f_{0}+P^{3}_{2}\ast f_{2}](x).$
Then it is not hard to see that
$\lim_{t\to 0}\mathscr{B}^{3}_{1}(u)(x,t)=\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{f_{2}(y)}{|x-y|^{n-1}}\mathrm{d}y.$
This together with (3.10) and (3.8) in turn implies that $\forall~{}\xi\in\mathbb{S}^{n}$,
$\mathscr{B}^{3}_{1}U(\xi)=$ $\left[\lim_{t\to 0}\left(\frac{(1+t)^{2}+|x|^{2}}{2}\right)^{\frac{n-1}{2}}\mathscr{B}^{3}_{1}(u)\right]\circ F^{-1}(\xi)$
$=$ $\left[\frac{1}{(n-1)|\mathbb{S}^{n}|}\left(\frac{1+|x|^{2}}{2}\right)^{\frac{n-1}{2}}\int_{\mathbb{R}^{n}}\frac{f_{2}(y)}{|x-y|^{n-1}}\mathrm{d}y\right]\circ F^{-1}(\xi)$
$=$ $\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\bar{f}_{2}(\eta)}{|\xi-\eta|^{n-1}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta)=\frac{1}{2}G_{P_{1}}\ast\bar{f}_{2},$
which leads to
$2P_{1}\mathscr{B}^{3}_{1}U=\mathscr{B}^{3}_{2}U\qquad\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}.$
$\Box$

### v2

@@
 Xuezhang Chen  and Shihong Zhang
-Abstract: We introduce the notion of a biharmonic Poisson kernel associated with certain pair of conformal boundary operators and present its explicit formula.
-With this powerful tool, we next establish the classification theorems of nonnegative solutions to a biharmonic equation on the upper half-space and unit ball with proper pairs of conformal boundary operators.
-$\mathbf{2020}$ MSC: 35B53,53C18 (35G30,35J08)
-Keywords: Conformal boundary operators, biharmonic Poisson kernel, biharmonic Green function.
+Abstract: For conformal boundary operators associated with the Paneitz operator, we introduce a rigorous definition of the biharmonic Poisson kernel consisting of a pair of kernel functions and derive its explicit representation formula.
+With this powerful tool, we establish classification theorems of nonnegative solutions to the $Q$-flat and constant $T$-curvature equations on $\mathbb{R}_{+}^{n+1}$ and $\mathbb{B}^{n+1}$.
+$\mathbf{2020}$ MSC: 35J40, 53C18 (35J91, 53C24)
+Keywords: $Q$-curvature, conformal boundary operator, $T$-curvature, biharmonic Poisson kernel.
 1 Introduction
-Caffarelli-Gidas-Spruck [5] established the celebrated Liouville-type theorem: For $n\geq 2$, a $C^{2}$ solution of
-$-\Delta u=(n+1)(n-1)u^{\frac{n+3}{n-1}},~{}~{}u>0\qquad\mathrm{~{}~{}in~{}~{}}\mathbb{R}^{n+1}$
+In a seminal work [7], Caffarelli-Gidas-Spruck established a Liouville-type theorem of the second-order critical semilinear elliptic equation: For $n\geq 2$, every positive $C^{2}$ solution to
+$-\Delta u=(n+1)(n-1)u^{\frac{n+3}{n-1}},\qquad\quad\mathrm{~{}~{}in~{}~{}}\mathbb{R}^{n+1}$
 is of the form
-$u(X)=\left(\frac{\varepsilon}{\varepsilon^{2}+|X-X_{0}|^{2}}\right)^{\frac{n-1}{2}},\qquad\varepsilon\in\mathbb{R}_{+},X_{0}\in\mathbb{R}^{n+1}.$
+$u(X)=\left(\frac{\varepsilon}{\varepsilon^{2}+|X-X_{0}|^{2}}\right)^{\frac{n-1}{2}},\qquad\mathrm{~{}~{}for~{}~{}some~{}~{}}\varepsilon\in\mathbb{R}_{+},X_{0}\in\mathbb{R}^{n+1}.$
 Under an additional hypothesis $u(X)=O(|X|^{1-n})$ as $|X|\to\infty$, the same result was also proved earlier by Obata [34] and Gidas-Ni-Nirenberg [21].
-Let $\mathbb{R}_{+}^{n+1}=\{X=(x,t)|~{}x\in\mathbb{R}^{n},t\in\mathbb{R}_{+}\}$ denote the upper half-space with boundary $\partial\mathbb{R}_{+}^{n+1}=\mathbb{R}^{n}$ and $\mathbb{B}^{n+1}$ denote the unit ball in $\mathbb{R}^{n+1}$.
-Escobar [13] used the Obata-type argument to characterize minimizers of a sharp Sobolev trace inequality on $\mathbb{B}^{n+1}$, via a conformal map between $\mathbb{R}_{+}^{n+1}$ and $\mathbb{B}^{n+1}$, which is equivalent to seeking smooth solutions to
-$\begin{cases}\displaystyle\Delta u=0,\quad u>0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{R}_{+}^{n+1},\\ \displaystyle-\partial_{t}u(x,0)=(n-1)u^{\frac{n+1}{n-1}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\end{cases}$ (1.1)
-under an extra hypothesis that $u(X)=O(|X|^{1-n})$ as $|X|\to\infty$.
+Escobar [17] used the Obata-type argument to characterize minimizers of a sharp Sobolev trace inequality on the unit ball $\mathbb{B}^{n+1}$, via a conformal map between $\mathbb{R}_{+}^{n+1}:=\{X=(x,t)\in\mathbb{R}^{n}\times\mathbb{R}_{+}\}$ and $\mathbb{B}^{n+1}$, which is equivalent to seeking positive smooth solutions to
+$\begin{cases}\displaystyle\Delta u=0,&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{R}_{+}^{n+1},\\ \displaystyle-\partial_{t}u(x,0)=(n-1)u^{\frac{n+1}{n-1}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\end{cases}$ (1.1)
+under a decay hypothesis that $u(X)=O(|X|^{1-n})$ as $|X|\to\infty$.
 Y.
@@
 Y.
-Li-Zhang [25]) employed the moving sphere method to prove the following Liouville-type theorem without the above hypothesis: A $C^{2}$ solution of (1.1) is of the form
-$u(x,t)=\left(\frac{\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-1}{2}},\qquad\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}.$
-Such Liouville-type theorems of general critical growth second order semilinear elliptic equations on $\mathbb{R}_{+}^{n+1}$ are referred to Escobar [14], Y. Y. Li-Zhang [25] etc.
-Sun-Xiong [30] classified nonnegative smooth solutions of the following conformally invariant polyharmonic boundary value problem: For $m\geq 2,n+1>2m$ and $1<p\leq\frac{n+2m-1}{n+1-2m}$,
+Li-Zhang [31]) employed the moving sphere method to prove the following Liouville-type theorem without the decay hypothesis: Every positive $C^{2}$ solution to (1.1) is of the form
+$u(x,t)=\left(\frac{\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-1}{2}},\qquad\mathrm{~{}~{}for~{}~{}some~{}~{}}\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}.$
+Such Liouville-type theorems of critical semilinear second-order elliptic equations on $\mathbb{R}_{+}^{n+1}$ are referred to Escobar [18], Chipot-Shafrir-Fila [11], Y.
+Y.
+Li-Zhang [31] and references therein.
+Sun-Xiong [36] classified nonnegative smooth solutions of the following conformally invariant polyharmonic boundary value problem (BVP): For $m\geq 2,n+1>2m$ and $1<p\leq\frac{n+2m-1}{n+1-2m}$,
 $\begin{cases}\displaystyle\Delta^{m}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\partial_{t}\Delta^{k}u(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle(-1)^{m}\partial_{t}\Delta^{m-1}u(x,0)=u^{p}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\end{cases}$ (1.2)
-where $20\leq k\leq m-$ 2. When $m$ is even, assume in addition that $u(X)=o(|X|^{2m-1})$ as $|X|\to\infty$, then for the critical exponent $p=\frac{n+2m-1}{n-(2m-1)}$,
-$u(x,t)=$ $c(n,m)\int_{\mathbb{R}^{n}}\frac{t^{2m-1}}{(|x-y|^{2}+t^{2})^{\frac{n+2m-1}{2}}}\left(\frac{\lambda}{\lambda^{2}+|y|^{2}}\right)^{\frac{n-2m+1}{2}}\mathrm{d}y+\sum_{k=1}^{m-1}t^{2k}P_{2k}(x),$
-where $x_{0}\in\mathbb{R}^{n},\lambda\in\mathbb{R}_{+}$, $c(n,m)$ is a normalized constant and $P_{2k}(x)$ denotes a polynomial of degree $2k$ in the variable $x$; for $1<p<\frac{n+2m-1}{n+1-2m}$,
-$u(x,t)=\sum_{k=1}^{m-1}t^{2k}P_{2k}(x).$
-A key step of their approach is to improve the integrability of a Kevin transformation $u^{\ast}$ for $(-\Delta)^{m}$ via an iteration method, and eventually to show that the singularity of $u^{\ast}$ can be removed at $t=0$; in the next step, the integral equation of $u^{\ast}$ on $\partial\mathbb{R}_{+}^{n+1}$ is solved by the classification theorems due to Chen-Li-Ou [9] and Y.
+where $20\leq k\leq m-$ 2. When $m$ is even, assume in addition that $u(X)=o(|X|^{2m-1})$ as $|X|\to\infty$. Then for the critical Sobolev exponent $p=\frac{n+2m-1}{n-(2m-1)}$,
+$u(x,t)=$ $c(n,m)\int_{\mathbb{R}^{n}}\frac{t^{2m-1}}{(|x-y|^{2}+t^{2})^{\frac{n+2m-1}{2}}}\left(\frac{\lambda}{\lambda^{2}+|y|^{2}}\right)^{\frac{n-2m+1}{2}}\mathrm{d}y+\sum_{k=1}^{m-1}t^{2k}P_{2m-2-2k}(x),$
+where $x_{0}\in\mathbb{R}^{n},\lambda\in\mathbb{R}_{+}$, $c(n,m)$ is a normalized constant and $P_{2m-2-2k}(x)$ denotes a polynomial of degree $\leq 2m-2-2k$ satisfying $\liminf_{|x|\to\infty}P_{2m-2-2k}(x)\geq 0$; for $1<p<\frac{n+2m-1}{n+1-2m}$,
+$u(x,t)=\sum_{k=1}^{m-1}t^{2k}P_{2m-2-2k}(x),$
+where $P_{2m-2-2k}$ is a nonnegative polynomial of degree $\leq 2m-2-2k$.
+One of crucial steps in [36] is to improve the integrability of a Kevin transformation $u^{\ast}$ for $(-\Delta)^{m}$ through an iteration method, and to show that the singularity of $u^{\ast}$ can be eventually removed at $t=0$.
+Next the integral equation of $u^{\ast}$ on $\partial\mathbb{R}_{+}^{n+1}$ is solved by the classification theorems due to Y.
 Y.
-Li [24], then the standard method can give the classification of nonnegative solutions to (1.2).
-Conformal boundary operators $\mathscr{B}_{k}^{3}$ for $k\in\{0,1,2,3\}$ associated with the Paneitz operator have attracted a lot of attention, see [18, 4, 6, 11] etc.
-In the study of zeta functional determinants on a four-manifold with boundary, Chang-Qing [8] discovered a third order conformally covariant boundary operator $\mathscr{B}_{3}^{3}$.
-The discovery of conformal boundary operators $\mathscr{B}_{k}^{3}$ to an $(n+1)$-manifold $(X^{n+1},g)$ with boundary $\partial X^{n+1}=M^{n}$ is due to Branson-Gover [3], Juhl [23], Gover-Peterson [19], Grant [20], Stafford [29] etc.
-The Paneitz operator $P_{4}^{g}$ and $\mathscr{B}_{k}^{3}$ enjoy the property (cf.
-[7, Theorem 1.1]) that if $\tilde{g}=e^{2\tau}g$, then for all $\Psi\in C^{\infty}(\overline{X^{n+1}})$,
+Li [30] and Chen-Li-Ou [12], then the classification of solutions to the BVP (1.2) follows by the the standard procedure.
+Conformal boundary operators $\mathscr{B}_{k}^{3}$ for $k\in\{1,2,3\}$ associated with the Paneitz operator have attracted a lot of attention, see [23, 6, 8, 15] etc.
+On a smooth four-manifold with boundary, Chang-Qing [10] discovered a third-order conformal boundary operator $\mathscr{B}_{3}^{3}$ and its associated curvature $T_{3}^{3}$.
+To generalize the Chang-Qing boundary operator to an $(n+1)$-manifold $(X^{n+1},g)$ with boundary $\partial X:=M^{n}$ for all conformal boundary operators $\mathscr{B}_{k}^{3},k\in\{1,2,3\}$, Branson-Gover [5] initiated the Dirichlet-to-Neumann programme for all the higher order GJMS operators, and introduced the first- and second-order conformal boundary operators $\mathscr{B}_{k}^{3}$ associated with the Paneitz operator.
+Grant [25] and Stafford [35] made suitable adjustments of the approach in [5] to conduct $\mathscr{B}_{3}^{3}$.
+Readers are referred to Juhl [29], Gover-Peterson [24] for other treatments.
+The Paneitz operator $P_{4}^{g}$ and $\mathscr{B}_{k}^{3}$ enjoy the property that if $\tilde{g}=e^{2\tau}g$, then for all $\Psi\in C^{\infty}(\overline{X^{n+1}})$,
 $\tilde{P}_{4}(\Psi)=e^{-\frac{n+5}{2}\tau}P_{4}^{g}(e^{\frac{n-3}{2}\tau}\Psi)$ (1.3)
@@
 ~$\tilde{\mathscr{B}}_{k}^{3}(\Psi)=e^{-\frac{n+2k-3}{2}\tau}\mathscr{B}_{k}^{3}(e^{\frac{n-3}{2}\tau}\Psi).$ (1.4)
-Readers are referred to Case [7] or authors [11] for explicit formulae of $P_{4}^{g}$ and $\mathscr{B}_{k}^{3}$ on manifolds with boundary.
-We are concerned with two conformally equivalent model spaces: $\mathbb{R}_{+}^{n+1}$ and $\mathbb{B}^{n+1}$.
-On the upper half-space (+$(\mathbb{R}^{n+1}_{+},\partial\mathbb{R}^{n+1}_{+},|\mathrm{d}X|^{2})$, the explicit formulae of conformal boundary operators are $\mathscr{B}^{3}_{0}(u)=u$ and
-$\mathscr{B}^{3}_{1}(u)=$ $-\partial_{t}u;$
-$\mathscr{B}^{3}_{2}(u)=$ $\partial_{t}^{2}u-\overline{\Delta}u;$
-$\mathscr{B}^{3}_{3}(u)=$ $\partial_{t}\Delta u+2\overline{\Delta}\partial_{t}u$
-for all $u\in C^{\infty}(\overline{\mathbb{R}_{+}^{n+1}})$.
-On the unit ball $(\mathbb{B}^{n+1},\mathbb{S}^{n},|\mathrm{d}\xi|^{2})$, these conformal boundary operators are $\mathscr{B}^{3}_{0}(U)=U$ and
+For future reference, adopting the notations as in [10] and [24], we call $(T_{k})_{g}:=\frac{2}{n-3}\mathscr{B}_{k}^{3}(1),k\in\{1,2,3\}$ the $T$-curvatures on $M$, as the higher order generalization of the boundary mean curvature $h_{g}=\frac{2}{n-3}\mathscr{B}_{1}^{3}(1)$.
+Readers are referred to Case [9] or the authors [15] for explicit formulae of $P_{4}^{g}$ and $\mathscr{B}_{k}^{3}$ on manifolds with boundary.
+We are concerned with two models: $\mathbb{R}_{+}^{n+1}$ and $\mathbb{B}^{n+1}$.
+On (+$(\mathbb{R}^{n+1}_{+},\partial\mathbb{R}^{n+1}_{+},|\mathrm{d}X|^{2})$, the explicit formulae of conformal boundary operators are $\mathscr{B}^{3}_{0}(u)=u$ and
+$\mathscr{B}^{3}_{1}(u)=-\partial_{t}u;\qquad\mathscr{B}^{3}_{2}(u)=\partial_{t}^{2}u-\overline{\Delta}u;\qquad\mathscr{B}^{3}_{3}(u)=$ $\partial_{t}\Delta u+2\overline{\Delta}\partial_{t}u$
+for all $u\in C^{\infty}(\overline{\mathbb{R}_{+}^{n+1}})$. On $(\mathbb{B}^{n+1},\mathbb{S}^{n},|\mathrm{d}\xi|^{2})$, these conformal boundary operators are $\mathscr{B}^{3}_{0}(U)=U$ and
 $\mathscr{B}^{3}_{1}(U)=$ $\frac{\partial U}{\partial r}+\frac{n-3}{2}U,$
@@
 $+\frac{n-3}{2}\frac{\partial U}{\partial r}+\frac{(n^{2}-1)(n-3)}{4}U$
-for all $U\in C^{\infty}(\overline{\mathbb{B}^{n+1}})$, where $r=|\xi|$ for $\xi\in\overline{\mathbb{B}^{n+1}}$; the corresponding boundary curvatures $(T_{i}^{3})_{g_{\mathbb{S}^{n}}}$ are constant and set $\mathbb{T}^{3}_{i}=\frac{n-3}{2}(T_{i}^{3})_{g_{\mathbb{S}^{n}}}$ for $1\leq i\leq 3$, explicitly,
+for all $U\in C^{\infty}(\overline{\mathbb{B}^{n+1}})$, where $r=|\xi|$ for $\xi\in\overline{\mathbb{B}^{n+1}}$; the corresponding $T$-curvatures $(T_{i}^{3})_{\mathbb{S}^{n}}$ are constant and set $\mathbb{T}^{3}_{i}=\frac{n-3}{2}(T_{i}^{3})_{\mathbb{S}^{n}}$ for $i\in\{1,2,3\}$, explicitly,
 $\mathbb{T}^{3}_{1}=\frac{n-3}{2},\qquad\mathbb{T}^{3}_{2}=\frac{(n-1)(n-3)}{2}\qquad\mathrm{and}\qquad\mathbb{T}^{3}_{3}=\frac{(n^{2}-1)(n-3)}{4}$
@@
 Convention..
-Let $\Delta$ and $\overline{\Delta}$ denote the Euclidean Laplacians on $\mathbb{R}^{n+1}$ and $\mathbb{R}^{n}$, respectively.
-Let $p_{k}^{\ast}=\frac{n+2k-3}{n-3}$ for $k\in\{0,1,2,3\}$, as $p_{k}^{\ast}+1$ is the critical Sobolev exponent associated to $\mathscr{B}^{3}_{k}$.
-This motivates us to classify nonnegative smooth solutions to biharmonic boundary value problems on $\mathbb{R}_{+}^{n+1}$ associated with pairs of conformal boundary operators $(\mathscr{B}^{3}_{i},\mathscr{B}^{3}_{j})$ for $i<j\in\{1,2,3\}$: For $p_{i},p_{j}>0$, provided $u(x,0)$ has some suitable decay at infinity,
-$\begin{cases}\displaystyle\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\quad~{}~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(u)=\mathbb{T}^{3}_{i}u^{p_{i}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{j}^{3}(u)=\mathbb{T}^{3}_{j}u^{p_{j}}&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (1.5)
-The nonlinear interaction among $\mathscr{B}^{3}_{i}$ and $\mathscr{B}^{3}_{j}$ brings us new obstacles, compared with vanishing lower order part of derivatives in [30].
-So, it has great difficulty in applying the approach in [30] to (1.5), as well as some peculiar phenomena occur.
-An extremely powerful tool in our approach is so-called biharmonic Poisson kernel consisting of two kernel functions corresponding to a proper pair of conformal boundary operators $(\mathscr{B}_{i}^{3},\mathscr{B}_{j}^{3})$.
-In the following, we separate our discussion into two cases: $n\neq 3$ and $n=3$.
-In the latter case, it involves the critical dimension for both the Paneitz operator and $\mathscr{B}_{3}^{3}$.
-Definition 1.1 (Poisson kernel for a biharmonic boundary value problem).
-For $i,j\in\{0,1,2,3\}$ and $i<j$, a pair of kernel functions $P_{i}^{3}\oplus P_{j}^{3}(X)$ for $X\in\mathbb{R}_{+}^{n+1}$ is called a biharmonic Poisson kernel associated to a pair of conformal boundary operators $(\mathscr{B}_{i}^{3},\mathscr{B}_{j}^{3})$, if the following Poisson integral
+Let $\Delta$ and $\overline{\Delta}$ denote Euclidean Laplacians on $\mathbb{R}^{n+1}$ and $\mathbb{R}^{n}$, respectively.
+Let $p_{k}^{\ast}=\frac{n+2k-3}{n-3}$ for $k\in\{1,2,3\}$, as $p_{k}^{\ast}+1$ is the critical Sobolev exponent associated to $\mathscr{B}^{3}_{k}$.
+We are interested in the classification of nonnegative solutions to the $Q$-flat and constant $T$-curvature type equations on $\mathbb{R}_{+}^{n+1}$: Let $n\geq 4$ and $p_{i},p_{j}>0$ for $i,j\in\{1,2,3\},i<j$, consider nonnegative smooth solutions to
+$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\quad~{}~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(u)=\mathbb{T}^{3}_{i}u^{p_{i}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{j}^{3}(u)=\mathbb{T}^{3}_{j}u^{p_{j}}&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\end{cases}$ (1.5)
+under hypothesis that $u(x,0)$ has some suitable decay at infinity.
+The nonlinear interaction among $\mathscr{B}^{3}_{i}$ and $\mathscr{B}^{3}_{j}$ brings us new obstacles, compared with vanishing of lower order derivative boundary conditions in [36].
+So, it has great difficulty in applying the approach in [36] to (1.5), and peculiar phenomena should occur simultaneously.
+Although the Poisson kernel of a higher order elliptic BVP might appear elsewhere in the literature, a rigorous definition of the biharmonic Poisson kernel to the fourth-order elliptic BVP (1.5) is still in demand.
+Definition 1.1 (Biharmonic Poisson kernel).
+For $i,j\in\{0,1,2,3\},i<j$, a pair of kernel functions $P_{i}^{3}\oplus P_{j}^{3}(X)$ for $X\in\mathbb{R}_{+}^{n+1}$ is called a biharmonic Poisson kernel associated to a conformal boundary operator pair $(\mathscr{B}_{i}^{3},\mathscr{B}_{j}^{3})$, if the following Poisson integral
 $v(X)=$ $[P_{i}^{3}\ast f_{i}+P_{j}^{3}\ast f_{j}](X)$
@@
 makes sense, provided $f_{i},f_{j}\in C^{2}(\mathbb{R}^{n})$ have suitable decays at infinity, and gives a classical solution to
-$\begin{cases}\displaystyle\Delta^{2}v=0&\mathrm{~{}~{}in~{}~{}}\quad~{}~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(v)=f_{i}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{j}^{3}(v)=f_{j}\qquad\,\,\,\,\,\,\,&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (1.6)
-Obviously, the notion of biharmonic Poisson kernel could be naturally extended to $\mathbb{B}^{n+1}$ with boundary condition $(\mathscr{B}_{i}^{3},\mathscr{B}_{j}^{3})$.
+$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}v=0&\mathrm{~{}~{}in~{}~{}}\quad~{}~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(v)=f_{i}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{j}^{3}(v)=f_{j}\qquad\,\,\,\,\,\,\,&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (1.6)
+We are now ready to give the explicit formulae of biharmonic Poisson kernels.
 Theorem 1.1.
-For $n\neq 3$ and $X=(x,t)\in\mathbb{R}_{+}^{n+1}$, we let
+For $n\geq 2$ and $X=(x,t)\in\mathbb{R}_{+}^{n+1}$, we let111 A heuristic observation allows us to find these explicit kernel functions $P_{k}^{3},k\in\{0,1,2,3\}$.
 $P_{0}^{3}(X)=$ $\frac{2(n+1)}{|\mathbb{S}^{n}|}\frac{t^{3}}{(t^{2}+|x|^{2})^{\frac{n+3}{2}}};$
@@
 $P_{2}^{3}(X)=$ $-\frac{1}{(n-1)|\mathbb{S}^{n}|}\frac{t}{(t^{2}+|x|^{2})^{\frac{n-1}{2}}};$
-$P_{3}^{3}(X)=$ $\frac{1}{(n-1)(n-3)|\mathbb{S}^{n}|}\frac{1}{(t^{2}+|x|^{2})^{\frac{n-3}{2}}}.$
+and
+$P_{3}^{3}(X)=\frac{1}{(n-1)(n-3)|\mathbb{S}^{n}|}\frac{1}{(t^{2}+|x|^{2})^{\frac{n-3}{2}}},\quad$ $\qquad\mathrm{~{}~{}when~{}~{}}n\neq 3;$
+$P_{3}^{3}(X)=-\frac{1}{4|\mathbb{S}^{3}|}\log(t^{2}+|x|^{2})+C,\quad C\in\mathbb{R},$ $\qquad\mathrm{~{}~{}when~{}~{}}n=3.$
 Suppose (i) $f_{0}\in C^{2}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$ ; (ii) $k\in\{1,2,3\}$, $f_{k}\in C^{2}(\mathbb{R}^{n})$ and $f_{k}=O(|x|^{-k-\delta_{k}})$ as $|x|\to\infty$ for some $\delta_{k}\in\mathbb{R}_{+}$.
-Then for $i,j\in\{0,1,2,3\}$ and $i+j\neq 3$, the explicit formula of Poisson kernel corresponding to biharmonic boundary value problem (1.6) is $P_{i}^{3}\oplus P_{j}^{3}$.
+Then for $i,j\in\{0,1,2,3\},i+j\neq 3$, the explicit formula of Poisson kernel corresponding to problem (1.6) is $P_{i}^{3}\oplus P_{j}^{3}$.
 Remark 1.1.
-It is straightforward to verify that other than two pairs $(\mathscr{B}_{0}^{3},\mathscr{B}_{3}^{3})$ and $(\mathscr{B}_{1}^{3},\mathscr{B}_{2}^{3})$, each $(\mathscr{B}_{i}^{3},\mathscr{B}_{j}^{3})$ satisfies ‘the complementing condition’ introduced by Agmon-Douglis-Nirenberg [1, Section 10], see also [15, Section 2.3].
-For a smooth compact Riemannian manifold $(X^{n+1},g)$ with boundary $M^{n}$ and $n\geq 4$, under the assumption that a certain conformal invariant $\lambda_{1}(P_{4}^{g})>0$ (see [7]), the existence and uniqueness of smooth solutions to the following boundary value problem was studied in [7, Proposition 1.3]: For $\bar{f}_{0},\bar{f}_{1}\in C^{\infty}(M^{n})$,
-$\begin{cases}\displaystyle P_{4}^{g}U=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad X^{n+1},\\ \displaystyle\mathscr{B}^{3}_{0}U=\bar{f}_{0}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,M^{n},\\ \displaystyle\mathscr{B}^{3}_{1}U=\bar{f}_{1}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,M^{n}.\end{cases}$
-On the model space $(\mathbb{B}^{n+1},\mathbb{S}^{n},|\mathrm{d}\xi|^{2})$, for each proper pair of conformal boundary operators, such a biharmonic boundary value problem is expected to be better understood, serving as a first step to the prescribing boundary curvature problem on manifolds with boundary.
+It is direct to verify that other than $(\mathscr{B}_{0}^{3},\mathscr{B}_{3}^{3})$ and $(\mathscr{B}_{1}^{3},\mathscr{B}_{2}^{3})$, the BVP $(\Delta^{2};\mathscr{B}_{i}^{3},\mathscr{B}_{j}^{3})$ for problem (1.6) satisfies the Lopatinski-Shapiro condition (e.g., Branson-Gover [5, Section 6]; or [9, p.311]) or complementing condition (e.g., Agmon-Douglis-Nirenberg [1, Section 10] or [20, Section 2.3]).
+Let
+$\Gamma(X-Y)=\Gamma(|X-Y|)=\begin{cases}\displaystyle\frac{1}{2(n-1)(n-3)|\mathbb{S}^{n}|}|X-Y|^{3-n}\qquad&\mathrm{~{}~{}when~{}~{}}n\neq 3,\\ \displaystyle-\frac{1}{4|\mathbb{S}^{3}|}\log|X-Y|&\mathrm{~{}~{}when~{}~{}}n=3.\end{cases}$
+denote the fundamental solution of $\Delta^{2}$ in $\mathbb{R}^{n+1}$.
+Remark 1.2.
+As a traditional way of the classical Poisson kernel, conformal boundary operators $\mathscr{B}_{k}^{3}$ indeed act as a bridge between the biharmonic Poisson kernel and fundamental solution: Let $n\geq 2$, for every fixed $X=(x,t)\in\mathbb{R}_{+}^{n+1}$ and all $Y=(y,0)\in\partial\mathbb{R}_{+}^{n+1}$, a direct computation yields
+$\begin{split}\mathscr{B}^{3}_{0}(\Gamma(X-Y))=\frac{1}{2}P^{3}_{3}(x-y,t);&\qquad\mathscr{B}^{3}_{1}(\Gamma(X-Y))=\frac{1}{2}P^{3}_{2}(x-y,t);\\ \mathscr{B}^{3}_{2}(\Gamma(X-Y))=-\frac{1}{2}P^{3}_{1}(x-y,t);&\qquad\mathscr{B}^{3}_{3}(\Gamma(X-Y))=-\frac{1}{2}P^{3}_{0}(x-y,t).\end{split}$ (1.7)
+Consider
+$\begin{cases}\displaystyle\quad\Delta^{2}v=f&\mathrm{~{}~{}in~{}~{}}\quad~{}~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(v)=0\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{j}^{3}(v)=0\qquad\,\,\,\,\,\,\,&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (1.8)
+Definition 1.2 (Biharmonic Green function).
+Let $n\geq 3$, $i,j\in\{0,1,2,3\},i+j\neq 3$ and $f\in C^{\infty}(\mathbb{R}_{+}^{n+1})$.
+A function $G^{(i,j)}:(X,Y)\in\overline{\mathbb{R}_{+}^{n+1}}\times\overline{\mathbb{R}_{+}^{n+1}}\mapsto\mathbb{R}\cup\{\infty\}$ for (1.8) is called a Green function, if $G^{(i,j)}(X,Y)=\Gamma(X-Y)+H^{(i,j)}(X,Y)$, where $H^{(i,j)}$ is a solution of
+$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}H^{(i,j)}=0\qquad&\mathrm{in}\qquad~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(H^{(i,j)})=-\mathscr{B}^{3}_{i}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{j}(H^{(i,j)})=-\mathscr{B}^{3}_{j}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}_{+}^{n+1}.\end{cases}$ (1.9)
+Remark 1.3.
+As a direct consequence of Theorem 1.1 and (1.7), problem (1.9) admits a unique classical solution in $\mathbb{R}_{+}^{n+1}$ for $n\geq 3$.
+More importantly, we can derive the explicit formulae of $H^{(i,j)}$ with the help of the Boggio formula in Boggio [4] or Gazzola-Grunau-Sweers [20, Lemma 2.27].
+For clarity, we distinguish these representation formulae into two cases: noncritical and critical dimensions.
 Theorem 1.2.
-Suppose $n\neq 3$ and $\bar{f}_{i}\in C^{\infty}(\mathbb{S}^{n})$ for $i\in\{0,1,2,3\}$. For $\xi\in\mathbb{B}^{n+1}$ and $\eta\in\mathbb{S}^{n}$, we let
-$\bar{P}_{0}^{3}(\xi,\eta)=$ $\frac{n+1}{4|\mathbb{S}^{n}|}\frac{\left(1-|\xi|^{2}\right)^{3}}{|\xi-\eta|^{n+3}};$
-$\bar{P}_{1}^{3}(\xi,\eta)=$ $-\frac{1}{2|\mathbb{S}^{n}|}\frac{(1-|\xi|^{2})^{2}}{|\xi-\eta|^{n+1}};$
-$\bar{P}_{2}^{3}(\xi,\eta)=$ $-\frac{1}{2(n-1)|\mathbb{S}^{n}|}\frac{1-|\xi|^{2}}{|\xi-\eta|^{n-1}};$
-$\bar{P}_{3}^{3}(\xi,\eta)=$ $\frac{1}{(n-1)(n-3)|\mathbb{S}^{n}|}\frac{1}{|\xi-\eta|^{n-3}}.$
-Then, for $i<j\in\{0,1,2,3\}$ we have
-  • (i)
-    For $i+j\neq 3$, the solution of the follow biharmonic boundary value problem
-    $\begin{cases}\displaystyle\Delta^{2}U=0\qquad\quad&\mathrm{in}\qquad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}U=\bar{f}_{i}\qquad&\mathrm{on}\qquad\mathbb{S}^{n},\\ \displaystyle\mathscr{B}^{3}_{j}U=\bar{f}_{j}\qquad&\mathrm{on}\qquad\mathbb{S}^{n},\end{cases}$ (1.7)
-    is unique and has the form of
-    $U(\xi)=\int_{\mathbb{S}^{n}}\bar{P}_{i}^{3}(\xi,\eta)\bar{f}_{i}(\eta)\mathrm{d}V_{\mathbb{S}^{n}}(\eta)+\int_{\mathbb{S}^{n}}\bar{P}_{j}^{3}(\xi,\eta)\bar{f}_{j}(\eta)\mathrm{d}V_{\mathbb{S}^{n}}(\eta).$
-  • (ii)
-    For pairs $(0,3)$ and $(1,2)$, problem (1.7) has multiple smooth solutions.
-Remark 1.2.
-It is notable that restricting $\xi$ to $\mathbb{S}^{n}$, $2\bar{P}_{3}^{3}(\xi,\eta)$ coincides with the Green function of fractional GJMS operator $P_{3}$ on $\mathbb{S}^{n}$ (cf.
-the first author and Shi [10, Theorem 1(3)]).
-As an application, we can recover a special case in Case [7, Theorem 1.4] about the relationship between extrinsic and intrinsic GJMS operators, see Proposition 3.1 and its biharmonic Poisson kernel proof in Appendix A.
-Moreover, the non-uniqueness in Theorem 1.2(ii) actually arises from the invertibility of fractional GJMS operators $P_{3}$ and $P_{1}$, see [7, 10] for the definition.
-We are ready to introduce a biharmonic Green function on $\mathbb{B}^{n+1}$ with each of four proper pairs of conformal boundary operators. Let
-$\Gamma(\xi-\eta)=\Gamma(|\xi-\eta|)=\frac{1}{2(n-1)(n-3)|\mathbb{S}^{n}|}|\xi-\eta|^{3-n}\qquad\mathrm{~{}~{}when~{}~{}}n\neq 3$
-and
-$\Gamma(\xi-\eta)=\Gamma(|\xi-\eta|)=-\frac{1}{4|\mathbb{S}^{3}|}\log|\xi-\eta|\qquad\qquad\mathrm{~{}~{}when~{}~{}}n=3$
-denote the fundamental solution of $\Delta^{2}$ in $\mathbb{R}^{n+1}$. We consider an nonhomogeneous biharmonic boundary value problem with vanishing boundary conditions
-$\begin{cases}\Delta^{2}U=\bar{f}\qquad&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{B}^{n+1},\\ \mathscr{B}_{i}^{3}U=0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{S}^{n},\\ \mathscr{B}_{j}^{3}U=0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{S}^{n}.\end{cases}$ (1.8)
-Remark 1.3.
-A clever observation in (4.1) provides a traditional way in understanding biharmonic Poisson kernel: Conformal boundary operators $\mathscr{B}_{i}^{3}$ indeed act as a bridge between the biharmonic Poisson kernel and fundamental solution.
-Definition 1.2 (biharmonic Green function).
-Let $\bar{f}\in C^{\infty}(\overline{\mathbb{B}^{n+1}})$ and $i<j\in\{0,1,2,3\}$ with $i+j\neq 3$.
-A function $\bar{G}^{(i,j)}:(\xi,\eta)\in\overline{\mathbb{B}^{n+1}}\times\overline{\mathbb{B}^{n+1}}\mapsto\mathbb{R}\cup\{\infty\}$ for (1.8) is called a Green function, if $\bar{G}^{(i,j)}$ satisfies
-  • $\xi\mapsto\bar{G}^{(i,j)}(\xi,\eta)-\Gamma(\xi-\eta)\in C^{\infty}(\overline{\mathbb{B}^{n+1}})$ for all $\eta\in\mathbb{B}^{n+1}$;
-  • $\Delta_{\xi}^{2}\left(\bar{G}^{(i,j)}(\xi,\eta)-\Gamma(\xi-\eta)\right)=0$ for all $(\xi,\eta)\in\mathbb{B}^{n+1}\times\mathbb{B}^{n+1}$;
-  • $\mathscr{B}_{i}^{3}\bar{G}^{(i,j)}=0,\mathscr{B}_{j}^{3}\bar{G}^{(i,j)}=0$ on $\mathbb{S}^{n}$.
-Formally, the above Green function enables us to express a unique solution to (1.8) as
-$U(\xi)=\bar{G}^{(i,j)}\ast\bar{f}(\xi)=\int_{\mathbb{B}^{n+1}}\bar{G}^{(i,j)}(\xi,\eta)\bar{f}(\eta)\mathrm{d}\eta.$
-We first notice that $\bar{G}^{(0,1)}$ coincides with the Boggio formula (see, e.g., Boggio [2] or Gazzola-Grunau-Sweers [15, Lemma 2.27]), next combine this with biharmonic Poisson kernel to derive explicit formulae of Green functions.
+For $n\geq 2$ and $n\neq 3$, the biharmonic Green functions for $\mathbb{R}_{+}^{n+1}$ are
+$G^{(0,2)}(X,Y)=$ $\Gamma(X-Y)-\Gamma(\bar{X}-Y);$
+$G^{(1,3)}(X,Y)=$ $\Gamma(X-Y)+\Gamma(\bar{X}-Y);$
+$G^{(0,1)}(X,Y)=$ $\Gamma(X-Y)+\frac{1}{2}\Gamma(|\bar{X}-Y|)\left[(n-3)\frac{|X-Y|^{2}}{|\bar{X}-Y|^{2}}-(n-1)\right];$
+$G^{(2,3)}(X,Y)=$ $\Gamma(X-Y)-\frac{1}{2}\Gamma(|\bar{X}-Y|)\left[(n-3)\frac{|X-Y|^{2}}{|\bar{X}-Y|^{2}}-(n-1)\right];$
+where $\bar{X}=(x,-t)$ for $X\in\mathbb{R}_{+}^{n+1}$. Moreover, there holds
+$0\leq G^{(0,1)}(X,Y)\leq G^{(0,2)}(X,Y)\leq G^{(1,3)}(X,Y)\leq G^{(2,3)}(X,Y),\quad\forall~{}X,Y\in\overline{\mathbb{R}_{+}^{n+1}},X\neq Y,$
+with equality if and only if either $X\in\mathbb{R}^{n}$ or $Y\in\mathbb{R}^{n}$.
 Theorem 1.3.
-For $n\neq 3$, the representation formula of Green function $\bar{G}^{(i,j)}$ to (1.8) is
-$\bar{G}^{(0,1)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\frac{1}{2}\Gamma(|\xi|\cdot|\xi^{\ast}-\eta|)\left[(n-3)\frac{|\xi-\eta|^{2}}{|\xi|^{2}\cdot|\xi^{\ast}-\eta|^{2}}-(n-1)\right];$
-$\bar{G}^{(2,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)-\frac{1}{2}\Gamma(|\xi|\cdot|\xi^{\ast}-\eta|)\left[(n-3)\frac{|\xi-\eta|^{2}}{|\xi|^{2}\cdot|\xi^{\ast}-\eta|^{2}}-(n-1)\right];$
-$\bar{G}^{(0,2)}(\xi,\eta)=$ $\Gamma(\xi-\eta)-\Gamma(|\xi|\cdot|\xi^{*}-\eta|);$
-$\bar{G}^{(1,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\Gamma(|\xi|\cdot|\xi^{*}-\eta|);$
-where $\xi^{\ast}=|\xi|^{-2}\xi$ for $0\neq\xi\in\overline{\mathbb{B}^{n+1}}$. Moreover, there holds
-$0\leq\bar{G}^{(0,1)}(\xi,\eta)\leq\bar{G}^{(0,2)}(\xi,\eta)\leq\bar{G}^{(1,3)}(\xi,\eta)\leq\bar{G}^{(2,3)}(\xi,\eta),\quad\forall~{}\xi\neq\eta\in\overline{\mathbb{B}^{n+1}},$
-with equality if and only if either $\xi\in\mathbb{S}^{n}$ or $\eta\in\mathbb{S}^{n}$.
-In the critical dimension that $n=3$, we have
-Theorem 1.4.
-Suppose $i,j\in\{0,1,2,3\}$ and $i<j$. The following are explicit formulae of biharmonic Poisson kernels on $\mathbb{R}_{+}^{4}$ and $\mathbb{B}^{4}$.
-  • (1)
-    Under the same assumptions as in Theorem 1.1 when $n=3$, then the same formulae of $P_{i}^{3}$ are true except for
-    $P_{3}^{3}(X)=-\frac{1}{4|\mathbb{S}^{3}|}\log(t^{2}+|x|^{2})+C,\qquad C\in\mathbb{R}$
-    for $X\in\mathbb{R}_{+}^{4}$, and the biharmonic Poisson kernel for (1.6) is $P_{i}^{3}\oplus P_{j}^{3}$ when $i+j\neq 3$.
-  • (2)
-    Under the same assumptions as in Theorem 1.2 when $n=3$, then the same formulae of $\bar{P}_{i}^{3}$ are true except for
-    $\bar{P}_{3}^{3}(\xi,\eta)=-\frac{1}{2|\mathbb{S}^{3}|}\log|\xi-\eta|+C,\qquad C\in\mathbb{R}$
-    for $\xi\in\mathbb{B}^{4},\eta\in\mathbb{S}^{3}$. Moreover, we have
-    (i) $i=1,2$ and $j=3$, the solution of (1.7) is
-    $U(\xi)-\frac{1}{|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}U\mathrm{d}V_{{\mathbb{S}^{3}}}$
-    $=$ $\int_{\mathbb{S}^{3}}\bar{P}^{3}_{i}(\xi-\eta)\bar{f}_{i}(\eta)\mathrm{d}V_{{\mathbb{S}^{3}}}(\eta)-\frac{1}{2|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}\log|\xi-\eta|\bar{f}_{3}(\eta)\mathrm{d}V_{\mathbb{S}^{3}}(\eta),\quad\forall~{}\xi\in\overline{\mathbb{B}^{4}}.$
-    (ii) $(i,j)=(0,1)\mathrm{~{}or~{}}(0,2)$, the unique solution of (1.7) is
-    $U(\xi)=\int_{\mathbb{S}^{3}}\bar{P}^{3}_{i}(\xi-\eta)\bar{f}_{i}(\eta)+\bar{P}^{3}_{j}(\xi-\eta)\bar{f}_{j}(\eta)\mathrm{d}V_{{\mathbb{S}^{3}}}(\eta),\quad\forall~{}\xi\in\overline{\mathbb{B}^{4}}.$
-Theorem 1.5.
-The representation formula of Green function $\bar{G}^{(i,j)}$ to (1.8) is
-$\bar{G}^{(0,1)}(\xi,\eta)=$ $\Gamma(\xi-\eta)-\Gamma(|\xi|\cdot|\xi^{*}-\eta|)+\frac{1}{8|\mathbb{S}^{3}|}\left(\frac{|\xi-\eta|^{2}}{|\xi|^{2}|\xi^{*}-\eta|^{2}}-1\right);$
-$\bar{G}^{(2,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\Gamma(|\xi|\cdot|\xi^{*}-\eta|)-\frac{1}{8|\mathbb{S}^{3}|}\left(\frac{|\xi-\eta|^{2}}{|\xi|^{2}|\xi^{*}-\eta|^{2}}-1\right)+C,\quad C\in\mathbb{R};$
-$\bar{G}^{(0,2)}(\xi,\eta)=$ $\Gamma(\xi-\eta)-\Gamma(|\xi|\cdot|\xi^{*}-\eta|);$
-$\bar{G}^{(1,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\Gamma(|\xi|\cdot|\xi^{*}-\eta|)+C,\quad C\in\mathbb{R};$
-for $\xi,\eta\in\overline{\mathbb{B}^{4}}$, where $\xi^{\ast}=|\xi|^{-2}\xi$ for $0\neq\xi\in\overline{\mathbb{B}^{4}}$.
-We shall assemble the biharmonic Poisson kernel part in Section 3 and the biharmonic Green function part in Section 4.
-We next turn to prove classification theorems of nonnegative solutions to (1.5) in $\mathbb{R}_{+}^{n+1}$.
-Before presenting the proof of our results, in Section 2 we first need to study the nontrivial solutions of an associated homogeneous problem to (1.6) under suitable assumptions on the decay rate at infinity of the initial data, and also describe asymptotic behavior of two typical singular integrals originating from the biharmonic Poisson integrals for $\mathbb{R}_{+}^{n+1}$.
-For simplicity, we introduce a ‘geometric bubble function’ associated to above biharmonic boundary value problem by
-$U_{\varepsilon,x_{0}}(x,t)=\left(\frac{2\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}},\qquad(x,t)\in\mathbb{R}_{+}^{n+1},$
-and its boundary function by $u_{\varepsilon,x_{0}}(x)=U_{\varepsilon,x_{0}}(x,0)$, where $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$.
-For the pair $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{3})$, under appropriate decay rate at infinity of initial data, the intrinsic nature of $\mathscr{B}^{3}_{3}$ boundary condition enables us to determine the explicit representation of $u$ on $\partial\mathbb{R}_{+}^{n+1}$, see Lemma 5.1.
-Using basic properties of $U_{\varepsilon,x_{0}}$ and biharmonic Poisson kernel, we establish that every nonnegative smooth solution is a sum of a bubble function and a polynomial.
-When $p_{3}=p_{3}^{\ast}$, a striking feature is that the bubble function (in general, without geometric meaning) does not vanish even if $p_{1}\neq p_{1}^{\ast}$, and is exact $U_{\varepsilon,x_{0}}$ when $p_{1}=p_{1}^{\ast}$; whereas for $p_{1}<p_{1}^{\ast}$, the bubble function vanishes.
+The representation formulae of Green functions for $\mathbb{R}_{+}^{4}$ are
+$G^{(0,1)}(X,Y)=$ $\Gamma(X-Y)-\Gamma(\bar{X}-Y)+\frac{1}{8|\mathbb{S}^{3}|}\left(\frac{|X-Y|^{2}}{|\bar{X}-Y|^{2}}-1\right);$
+$G^{(2,3)}(X,Y)=$ $\Gamma(X-Y)+\Gamma(\bar{X}-Y)-\frac{1}{8|\mathbb{S}^{3}|}\left(\frac{|X-Y|^{2}}{|\bar{X}-Y|^{2}}-1\right)+C;$
+$G^{(0,2)}(X,Y)=$ $\Gamma(X-Y)-\Gamma(\bar{X}-Y);$
+$G^{(1,3)}(X,Y)=$ $\Gamma(X-Y)+\Gamma(\bar{X}-Y)+C;$
+for $X,Y\in\overline{\mathbb{R}_{+}^{4}},X\neq Y$ and $C\in\mathbb{R}$.
+Since $\mathbb{R}_{+}^{n+1}$ and $\mathbb{B}^{n+1}$ are conformally equivalent, we extend to introduce the notions of the biharmonic Poisson kernel $\bar{P}_{k}^{3},k\in\{0,1,2,3\}$ and Green function $\bar{G}^{(i,j)},i,j\in\{0,1,2,3\},i<j,i+j\neq 3$ for $\mathbb{B}^{n+1}$, and derive their explicit formulae as well.
+Until now, we have succeeded in obtaining a fourth-order analogue of the classical Poisson kernel and Green function for elliptic equations.
+For $\mathbb{R}_{+}^{n+1}$ and $\mathbb{B}^{n+1}$, we assemble the biharmonic Poisson kernel part in Section 3 and the biharmonic Green function part in Section 4.
+To continue, we introduce a ‘geometric bubble’ by
+$U_{x_{0},\varepsilon}(x,t)=\left(\frac{2\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}},\qquad(x,t)\in\mathbb{R}_{+}^{n+1},$
+and its restriction to the boundary by letting $U_{x_{0},\varepsilon}(x)=U_{x_{0},\varepsilon}(x,0)$, where $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$.
+For $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{3})$, under appropriate decay rate of the initial data, the intrinsic $\mathscr{B}^{3}_{3}$ equation with critical/subcritical growth enables us to determine the explicit formula for $u$ on $\partial\mathbb{R}_{+}^{n+1}$, see Lemma 5.1.
+With the aid of the basics of $U_{x_{0},\varepsilon}$ and biharmonic Poisson integral, we establish that every nonnegative smooth solution is a sum of a bubble and a polynomial.
+In sharp contrast to second-order elliptic equations, fix $p_{3}=p_{3}^{\ast}$, a striking feature is that a nontrivial bubble (in general, without geometric meaning) exists even for $p_{1}\neq p_{1}^{\ast}$, and coincides with $U_{x_{0},\varepsilon}$ for $p_{1}=p_{1}^{\ast}$; whereas for $p_{1}<p_{1}^{\ast}$, the solution is trivial.
 Theorem 1.4.
 For $n\geq 4$ and $p_{1},p_{3}>0$, let $u$ be a nonnegative smooth solution of
-$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}_{1}^{3}(u)=\mathbb{T}^{3}_{1}u^{p_{1}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}(u)=\mathbb{T}^{3}_{3}u^{p_{3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
-Assume $u(x,0)=O(|x|^{-c})$ as $|x|\to\infty$ for some constant $c>\max\{3/p_{3},1/p_{1}\}$. Then we have
-  • (1)
-    If $p_{3}=p_{3}^{\ast}=\frac{n+3}{n-3}$ and $p_{1}>\frac{1}{n-3}$, then there exist some $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$ and nonnegative constant $c_{2}$ such that
+$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\qquad~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}_{1}^{3}(u)=\mathbb{T}^{3}_{1}u^{p_{1}}&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}(u)=\mathbb{T}^{3}_{3}u^{p_{3}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
+Assume that $u(x,0)=O(|x|^{-c})$ as $|x|\to\infty$ for some $c>\max\{3/p_{3},1/p_{1}\}$. Then we have
+  1. (1)
+    If $p_{3}=p_{3}^{\ast}=\frac{n+3}{n-3}$ and $p_{1}>\frac{1}{n-3}$, then either $u=0$ or there exist some $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$ and nonnegative constant $c_{2}$ such that
     $u(x,t)=M_{1}(x,t)+c_{2}t^{2},$
     where
-    $M_{1}(x,t)=\left[\mathbb{T}_{1}^{3}P_{1}^{3}\ast u_{\varepsilon,x_{0}}^{p_{1}}+\mathbb{T}_{3}^{3}P_{3}^{3}\ast u_{\varepsilon,x_{0}}^{p_{3}^{\ast}}\right](x,t).$
-    In particular, if $p_{1}=p_{1}^{\ast}=\frac{n-1}{n-3}$, then
-    $u(x,t)=U_{\varepsilon,x_{0}}(x,t)+c_{2}t^{2}.$
-  • (2)
+    $M_{1}(x,t)=\left[\mathbb{T}_{1}^{3}P_{1}^{3}\ast U_{x_{0},\varepsilon}^{p_{1}}+\mathbb{T}_{3}^{3}P_{3}^{3}\ast U_{x_{0},\varepsilon}^{p_{3}^{\ast}}\right](x,t).$
+    In particular, when $p_{1}=p_{1}^{\ast}=\frac{n-1}{n-3}$, then
+    $u(x,t)=U_{x_{0},\varepsilon}(x,t)+c_{2}t^{2}.$
+  2. (2)
     If $0<p_{3}<p_{3}^{\ast}$, then $u(x,t)=c_{2}t^{2}$ for every nonnegative constant $c_{2}$.
-For a nonlinear boundary conditions involving the exceptional pair $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{2})$, there still exists a surprising classification theorem for problem (1.5) relying not on $p_{1},p_{2}$ but on $p_{2}/p_{1}$.
+Although the BVP $(\Delta^{2};\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{2})$ on $\mathbb{R}_{+}^{n+1}$ is not well-posed, instead we take the derivative $\partial_{t}$ of biharmonic Poisson integral of type $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{2})$ to obtain the integral equation of $u^{p_{1}}(x,0)$, and the classification theorem by Y.
+Y.
+Li [30] and Chen-Li-Ou [12]) gives the explicit formula of $u(x,0)$.
+Our classification result relies not on $p_{1}$ and $p_{2}$ but on the condition whether $p_{2}/p_{1}$ is critical or subcritical Sobolev exponents.
 Theorem 1.5.
 For $n\geq 4$ and $p_{1},p_{2}>0$, let $u$ be a nonnegative smooth solution of
-$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{1}(u)=\mathbb{T}^{3}_{1}u^{p_{1}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}^{3}_{2}(u)=\mathbb{T}^{3}_{2}u^{p_{2}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
-Assume $u(x,0)=O(|x|^{-c})$ as $|x|\to\infty$ for some constant $c>2/p_{2}$. Then we have
-  • (1)
-    If $\frac{p_{2}}{p_{1}}=\frac{n+1}{n-1}$ and $p_{1}>0$, then there exist some $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$ and nonnegative constant $c_{3}$ such that
+$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}u=0,&\mathrm{~{}~{}in~{}~{}}\qquad~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{1}(u)=\mathbb{T}^{3}_{1}u^{p_{1}}&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}^{3}_{2}(u)=\mathbb{T}^{3}_{2}u^{p_{2}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
+Assume that $u(x,0)=O(|x|^{-c})$ as $|x|\to\infty$ for some $c>2/p_{2}$. Then we have
+  1. (1)
+    If $\frac{p_{2}}{p_{1}}=\frac{n+1}{n-1}$ and $p_{1}>0$, then either $u=0$ or there exist some $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$ and nonnegative constant $c_{3}$ such that
     $u(x,t)=M_{2}(x,t)+c_{3}t^{3},$
     where
-    $M_{2}(x,t)=\left[P_{0}^{3}\ast u_{\varepsilon,x_{0}}^{\frac{p_{1}^{\ast}}{p_{1}}}+\mathbb{T}_{2}^{3}P_{2}^{3}\ast u_{\varepsilon,x_{0}}^{p_{2}^{\ast}}\right](x,t).$
-    In particular, $p_{1}=p_{1}^{\ast}=\frac{n-1}{n-3}$, then
-    $u(x,t)=U_{\varepsilon,x_{0}}(x,t)+c_{3}t^{3}.$
-  • (2)
+    $M_{2}(x,t)=\Big{[}P_{0}^{3}\ast U_{x_{0},\varepsilon}^{\frac{p_{1}^{\ast}}{p_{1}}}+\mathbb{T}_{2}^{3}P_{2}^{3}\ast U_{x_{0},\varepsilon}^{p_{2}^{\ast}}\Big{]}(x,t).$
+    In particular, when $p_{1}=p_{1}^{\ast}=\frac{n-1}{n-3}$, then
+    $u(x,t)=U_{x_{0},\varepsilon}(x,t)+c_{3}t^{3}.$
+  2. (2)
     If $\frac{p_{2}}{p_{1}}<\frac{n+1}{n-1}$ , then $u(x,t)=c_{3}t^{3}$ for every nonnegative constant $c_{3}$.
-The pair $(\mathscr{B}^{3}_{2},\mathscr{B}^{3}_{3})$ becomes more challenging and $p_{3}$ is still the dominant exponent.
-Especially, when $p_{i}=p_{i}^{\ast}$ for $i=1,2,3$, the bubble function happens to be the same $U_{\varepsilon,x_{0}}$; meanwhile, the polynomial part varies according to distinct circumstances.
+For $(\mathscr{B}^{3}_{2},\mathscr{B}^{3}_{3})$, the analysis of the polynomial part becomes more challenging, and requires the techniques developed in Theorems 1.4 and 1.5.
 Theorem 1.6.
 For $n\geq 4$ and $p_{2},p_{3}>0$, let $u$ be a nonnegative smooth solution of
-$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{2}(u)=\mathbb{T}^{3}_{2}u^{p_{2}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}(u)=\mathbb{T}^{3}_{3}u^{p_{3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
-Assume $u(x,0)=O(|x|^{-c})$ as $|X|\to\infty$ for some constant $c>\max\{2/p_{2},3/p_{3}\}$. Then we have
-  • (1)
-    If $p_{3}=p_{3}^{\ast}=\frac{n+3}{n-3}$ and $p_{2}>\frac{2}{n-3}$, then there exist some $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$ and nonnegative constant $c_{1}$ such that
+$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\qquad~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{2}(u)=\mathbb{T}^{3}_{2}u^{p_{2}}&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}(u)=\mathbb{T}^{3}_{3}u^{p_{3}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
+Assume that $u(x,0)=O(|x|^{-c})$ as $|X|\to\infty$ for some $c>\max\{2/p_{2},3/p_{3}\}$. Then we have
+  1. (1)
+    If $p_{3}=p_{3}^{\ast}=\frac{n+3}{n-3}$ and $p_{2}>\frac{2}{n-3}$, then either $u=0$ or there exist some $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$ and nonnegative constant $c_{1}$ such that
     $u(x,t)=M_{3}(x,t)+c_{1}t,$
     where
-    $M_{3}(x,t)=\left[P_{0}^{3}\ast u_{\varepsilon,x_{0}}+\mathbb{T}_{2}^{3}P_{2}^{3}\ast u_{\varepsilon,x_{0}}^{p_{2}}\right](x,t).$
-    In particular, $p_{2}=p_{2}^{\ast}=\frac{n+1}{n-3}$, then
-    $u(x,t)=U_{\varepsilon,x_{0}}(x,t)+c_{1}t.$
-  • (2)
+    $M_{3}(x,t)=\left[P_{0}^{3}\ast U_{x_{0},\varepsilon}+\mathbb{T}_{2}^{3}P_{2}^{3}\ast U_{x_{0},\varepsilon}^{p_{2}}\right](x,t).$
+    In particular, when $p_{2}=p_{2}^{\ast}=\frac{n+1}{n-3}$, then
+    $u(x,t)=U_{x_{0},\varepsilon}(x,t)+c_{1}t.$
+  2. (2)
     If $0<p_{3}<\frac{n+3}{n-3}$, then $u(x,t)=c_{1}t$ for every nonnegative constant $c_{1}$.
-These classification results demonstrate that the dominant impact of $\mathscr{B}^{3}_{i}$ is different, accordingly, $\mathscr{B}^{3}_{1}\preceq\mathscr{B}^{3}_{2}\preceq\mathscr{B}^{3}_{3}$.
-In Section 5 we shall present the proof of Theorems 1.6-1.8.
-The above classification results in $\mathbb{R}_{+}^{n+1}$ enable us to classify nonnegative solutions to biharmonic boundary value problems on $\mathbb{B}^{n+1}$ with one boundary point singularity.
+In conclusion, $p_{3}$ is the dominant exponent and the dominant impacts of $\mathscr{B}^{3}_{k}$ are different, precisely, $\mathscr{B}^{3}_{1}\preceq\mathscr{B}^{3}_{2}\preceq\mathscr{B}^{3}_{3}$.
+Especially, for $p_{i}=p_{i}^{\ast},i\in\{1,2,3\}$, the bubble happens to be the same $U_{x_{0},\varepsilon}$; meanwhile, the corresponding polynomial varies according to various circumstances.
+Theorems 1.4-1.6 give a complete picture of Liouville-type theorems to the study of the $Q$-flat and constant $T$-curvature problem on compact manifolds with boundary in the future.
+In Section 2 we study the nontrivial solutions of a homogeneous biharmonic BVP to (1.6) (i.e., $f_{i}=f_{j}=0$) under suitable decay assumption of the initial data, and also describe asymptotic behavior of two typical singular integrals originating from the biharmonic Poisson integrals for $\mathbb{R}_{+}^{n+1}$.
+In Section 5 we present the proof of classification Theorems 1.4-1.6.
+Theorems 1.4-1.6 enable us to classify nonnegative solutions to the $Q$-flat and constant $T$-curvature equations on $\mathbb{B}^{n+1}$ with one boundary point singularity.
 Theorem 1.7.
 Suppose $n\geq 4$ and $U\in C^{4}(\overline{\mathbb{B}^{n+1}}\backslash\{-\mathbf{e}_{n+1}\})$ is a nonnegative solution to
-$\begin{cases}\displaystyle\Delta^{2}U=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(U)=\mathbb{T}^{3}_{i}U^{\frac{n+2i-3}{n-3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\mathbb{S}^{n}\backslash\{-\mathbf{e}_{n+1}\},\\ \displaystyle\mathscr{B}^{3}_{j}(U)=\mathbb{T}^{3}_{j}U^{\frac{n+2j-3}{n-3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\mathbb{S}^{n}\backslash\{-\mathbf{e}_{n+1}\},\end{cases}$ (1.9)
-where $i,j\in\{1,2,3\}$ with $i<j$. Assume that there exist $\alpha\in(0,\frac{n(n-3)}{n+3})$ and $C\in\mathbb{R}_{+}$ such that
+$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}U=0&\mathrm{~{}~{}in~{}~{}}\qquad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(U)=\mathbb{T}^{3}_{i}U^{\frac{n+2i-3}{n-3}}&\mathrm{~{}~{}on~{}~{}}\qquad\mathbb{S}^{n}\backslash\{-\mathbf{e}_{n+1}\},\\ \displaystyle\mathscr{B}^{3}_{j}(U)=\mathbb{T}^{3}_{j}U^{\frac{n+2j-3}{n-3}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\qquad\mathbb{S}^{n}\backslash\{-\mathbf{e}_{n+1}\},\end{cases}$ (1.10)
+where $i,j\in\{1,2,3\},i<j$. Assume that there exist $\alpha\in(0,\frac{n(n-3)}{n+3})$ and $C\in\mathbb{R}_{+}$ such that
 $\sup_{\xi\in\mathbb{S}^{n}\backslash\{-\mathbf{e}_{n+1}\}}U(\xi)|\xi+\mathbf{e}_{n+1}|^{\alpha}\leq C.$ (1.11)
@@
 If $i=1,j=2$, then the assumption of $\alpha$ can be relaxed to $\alpha\in(0,\frac{(n-1)(n-3)}{n+1})$, which ensures that $n-3-\alpha>\frac{2}{p_{2}^{\ast}}$ by virtue of Theorem 1.5.
-A direct consequence of Theorem 1.9 is to prove the uniqueness of positive smooth solutions to a certain type of geometric partial differential equations on $\mathbb{B}^{n+1}$.
-Corollary 1.1.
-Suppose $n\geq 4$ and $U$ is a positive smooth solution in $\overline{\mathbb{B}^{n+1}}$ to
-$\begin{cases}\displaystyle\Delta^{2}U=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(U)=\mathbb{T}^{3}_{i}U^{\frac{n+2i-3}{n-3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\mathbb{S}^{n},\\ \displaystyle\mathscr{B}^{3}_{j}(U)=\mathbb{T}^{3}_{j}U^{\frac{n+2j-3}{n-3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\mathbb{S}^{n},\end{cases}$
-where $i<j\in\{1,2,3\}$. Then
-$U(\xi)=\left(\frac{1-|\xi_{0}|^{2}}{|\xi|^{2}|\xi_{0}|^{2}-2\xi_{0}\cdot\xi+1}\right)^{\frac{n-3}{2}}$
-for some $\xi_{0}\in\mathbb{B}^{n+1}$.
-We now pause to comment on the decay assumptions at infinity of initial data that $u(x,0)=O(|x|^{-c})$ as $|x|\to\infty$ in Theorems 1.6-1.8.
-For critical exponents $p_{i}^{\ast}$, such assumptions are to ensure that involved Poisson integrals are well-defined.
-As a result, this deduces that $c>\frac{3(n-3)}{n+3}$ in Theorems 1.6 and 1.8; $c>\frac{2(n-3)}{n+1}$ in Theorem 1.7.
-As will show in the proof of Theorem 1.9, we know $c=n-3-\alpha$, then the assumption (1.10) leads to $\alpha<\frac{n(n-3)}{n+1}$ and $\alpha<\frac{(n-1)(n-3)}{n+1}$, respectively.
-Another interpretation of assumption of $\alpha$ is as follows: If $\mathscr{B}^{3}_{3}(U)=\frac{n-3}{2}\tilde{T}^{3}_{3}U^{\frac{n+3}{n-3}}\in L^{1}(\mathbb{S}^{n})$ for a smooth conformal metric $\tilde{g}=U^{4/(n-3)}|\mathrm{d}\xi|^{2}$, then near the singular point $-\mathbf{e}_{n+1}$,
+We would digress for the moment to comment on the decay assumption of initial data that $u(x,0)=O(|x|^{-c})$ as $|x|\to\infty$ in Theorems 1.4-1.6.
+For critical exponents $p_{i}^{\ast}$, such assumptions are to ensure that involved biharmonic Poisson integrals are well-defined.
+Consequently, this leads to $c>\frac{3(n-3)}{n+3}$ in Theorems 1.4 ,1.6 and $c>\frac{2(n-3)}{n+1}$ in Theorem 1.5.
+As will show in the proof of Theorem 1.7, we indeed have $c=n-3-\alpha$, so (1.11) gives $\alpha<\frac{n(n-3)}{n+1}$ and $\alpha<\frac{(n-1)(n-3)}{n+1}$, respectively.
+Besides this, if $\mathscr{B}^{3}_{3}(U)=\frac{n-3}{2}\tilde{T}^{3}_{3}U^{\frac{n+3}{n-3}}\in L^{1}(\mathbb{S}^{n})$ for a conformal metric $\tilde{g}=U^{4/(n-3)}|\mathrm{d}\xi|^{2}$, then near $-\mathbf{e}_{n+1}$ we have
 $U^{\frac{n+3}{n-3}}(\xi)\lesssim|\xi+\mathbf{e}_{n+1}|^{-\frac{\alpha(n+3)}{n-3}}\in L^{1}(\mathbb{S}^{n}),$
@@
 Intuitively, $\mathscr{B}^{3}_{3}(U)\in L^{1}(\mathbb{S}^{n})$ means $U\in W^{3,1}(\mathbb{S}^{n})\hookrightarrow W^{2,\frac{n}{n-1}}(\mathbb{S}^{n})$.
-This roughly gives $\mathscr{B}^{3}_{2}(U)\in L^{\frac{n}{n-1}}(\mathbb{S}^{n})$, and also
+This roughly yields $\mathscr{B}^{3}_{2}(U)\in L^{\frac{n}{n-1}}(\mathbb{S}^{n})$, and also
 $\left(U^{\frac{n+1}{n-3}}(\xi)\right)^{\frac{n}{n-1}}\lesssim|\xi+\mathbf{e}_{n+1}|^{\frac{\alpha n(n+1)}{(n-1)(n-3)}}\in L^{1}(\mathbb{S}^{n}),$
 which requires $\alpha<\frac{(n-1)(n-3)}{n+1}$.
-In Section 6, we give the proof of Theorem 1.9, and as applications we also establish two sharp geometric inequalities over some $Q$-curvature flat conformal class of $(\mathbb{B}^{n+1},|\mathrm{d}\xi|^{2})$ for $n\geq 4$, including a sharp inequality of isoperimetric type.
+Corollary 1.1.
+Suppose $n\geq 4$ and $U$ is a positive smooth solution in $\overline{\mathbb{B}^{n+1}}$ to
+$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}U=0&\mathrm{~{}~{}in~{}~{}}\qquad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(U)=\mathbb{T}^{3}_{i}U^{\frac{n+2i-3}{n-3}}&\mathrm{~{}~{}on~{}~{}}\qquad\mathbb{S}^{n},\\ \displaystyle\mathscr{B}^{3}_{j}(U)=\mathbb{T}^{3}_{j}U^{\frac{n+2j-3}{n-3}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\qquad\mathbb{S}^{n},\end{cases}$
+where $i,j\in\{1,2,3\},i<j$. Then
+$U(\xi)=\left(\frac{1-|\xi_{0}|^{2}}{|\xi|^{2}|\xi_{0}|^{2}-2\xi_{0}\cdot\xi+1}\right)^{\frac{n-3}{2}}$
+for some $\xi_{0}\in\mathbb{B}^{n+1}$.
+A geometric interpretation of Corollary 1.1 is that, Let $g$ be a $Q$-flat conformal metric on $(\mathbb{B}^{n+1},|\mathrm{d}\xi|^{2})$ for $n\geq 4$.
+Assume that $\big{(}(T^{3}_{i})_{g},(T^{3}_{j})_{g}\big{)}=\big{(}(T^{3}_{i})_{\mathbb{S}^{n}},(T^{3}_{j})_{\mathbb{S}^{n}}\big{)}$ for $i,j\in\{1,2,3\},i<j$.
+Then there exists a conformal diffeomorphism $\Psi$ on $\overline{\mathbb{B}^{n+1}}$ such that $g=\Psi^{\ast}(|\mathrm{d}\xi|^{2})$.
+In Section 6, we give the proof of Theorem 1.7.
+As applications we establish two sharp geometric inequalities over certain $Q$-flat conformal class of $(\mathbb{B}^{n+1},|\mathrm{d}\xi|^{2})$ for $n\geq 4$, including a sharp fourth-order isoperimetric inequality.
+In Section 7, we raise a rigidity conjecture for a geodesic ball in $\mathbb{S}^{n+1}$, which is a natural extension of Corollary 1.1, and could be reduced to the classification of positive solutions to the constant $Q$-curvature and constant $T$-curvature equation in $\mathbb{B}^{n+1}$.
+We prove a uniqueness result of positive solutions to a fourth-order geometric ODE in Appendix B.
+This allows us to answer this conjecture affirmatively for radial solutions.
+In Appendix A, we give an alternative proof of Proposition 3.1.
 Acknowledgments: The authors thank Professor Qianqiao Guo for bringing [22] into their sight.
 2 Preliminaries
-We give asymptotic behavior of two prototypes of singular integrals, coming from biharmonic Poisson integrals on $\mathbb{R}_{+}^{n+1}$.
+This section is devote to building some technical lemmas and giving asymptotic behavior of two prototypes of singular integrals, which are two variants of biharmonic Poisson integrals for $\mathbb{R}_{+}^{n+1}$.
 Lemma 2.1.
-Suppose $f\in L_{\mathrm{loc}}^{1}(\mathbb{R}^{n})$ with $f(x)=O(|x|^{-a})$ as $|x|\to\infty$ for some $a\in\mathbb{R}_{+}$. Assume $20<\alpha<n/$ 2 and $a+2\alpha>n$. We define
+Suppose $f\in L_{\mathrm{loc}}^{1}(\mathbb{R}^{n})$ with $f(x)=O(|x|^{-a})$ as $|x|\to\infty$ for some $a\in\mathbb{R}_{+}$. Assume that $20<\alpha<n/$ 2 and $a+2\alpha>n$. We define
 $v(x,t)=\int_{\mathbb{R}^{n}}\frac{f(y)}{(t^{2}+|x-y|^{2})^{\alpha}}\mathrm{d}y.$
 Then there holds
-$\left|v(X)\right|\lesssim$$\begin{cases}|X|^{n-2\alpha-\min\{a,n\}}\quad&\mathrm{if~{}~{}}a\neq n\\ |X|^{-2\alpha}\log|X|&\mathrm{if~{}~{}}a=n\end{cases},\qquad\quad\mathrm{~{}~{}as~{}~{}}\quad|X|\to\infty.$
+$\left|v(X)\right|\lesssim$$\begin{cases}|X|^{n-2\alpha-\min\{a,n\}}\quad&\mathrm{if~{}~{}}a\neq n\\ |X|^{-2\alpha}\log|X|&\mathrm{if~{}~{}}a=n\end{cases}\qquad\quad\mathrm{~{}~{}as~{}~{}}\quad|X|\to\infty.$
 Proof..
 We decompose $\mathbb{R}^{n}=A_{1}\cup A_{2}\cup A_{3}$ with
-$A_{1}=$$\left\{y||y|<|x|/2\right\},\qquad\qquad\qquad\qquad A_{2}=\left\{y||x-y|<|x|/2\right\},$
-$A_{3}=$$\left\{y||y|\geq|x|/2,|x-y|\geq|x|/2\right\}.$
+$A_{1}=$$\left\{y;|y|<|x|/2\right\},\qquad\qquad\qquad\qquad A_{2}=\left\{y;|x-y|<|x|/2\right\},$
+$A_{3}=$$\left\{y;|y|\geq|x|/2,|x-y|\geq|x|/2\right\}.$
 Case 1. $|X|\gg 1$ and $t>\frac{|x|}{2}$.
 There hold $t\gg 1$ and $|X|\sim t$ due to
-$1\leq\frac{|X|^{2}}{t^{2}}=\frac{t^{2}+|x|^{2}}{t^{2}}\leq 5$
+$t^{2}\leq|X|^{2}=t^{2}+|x|^{2}\leq 5t^{2}.$
 In $A_{1}$, we have $|x-y|\sim|x|$ and $t^{2}+|x-y|^{2}\sim t^{2}$, hence
-$|v(x,t)|\lesssim t^{-2\alpha}\int_{A_{1}}|f(y)|\mathrm{d}y\lesssim$$\begin{cases}t^{-2\alpha}|x|^{n-\min\{a,n\}}&\mathrm{if~{}~{}}a\neq n\\ t^{-2\alpha}\log(|x|+1)&\mathrm{if~{}~{}}a=n\end{cases}$
-$\lesssim$ $\begin{cases}|X|^{n-2\alpha-\min\{a,n\}}&\mathrm{if~{}~{}}a\neq n\\ |X|^{-2\alpha}\log|X|&\mathrm{if~{}~{}}a=n\end{cases}.$
+$|v(x,t)|\lesssim t^{-2\alpha}\int_{A_{1}}|f(y)|\mathrm{d}y\lesssim\begin{cases}|X|^{n-2\alpha-\min\{a,n\}}&\mathrm{if~{}~{}}a\neq n\\ |X|^{-2\alpha}\log|X|&\mathrm{if~{}~{}}a=n\end{cases}.$
 In $A_{2}$, we have $|x-y|<|x|/2<t$, $|y|\sim|x|$ and $t^{2}+|x-y|^{2}\sim t^{2}$, then
@@
 In $A_{3}$, we have $|x-y|\leq|y|+|x|\leq 3|y|$ and $|y|\leq|x-y|+|x|\leq 3|x-y|$, whence $|x-y|\sim|y|$ We further decompose $A_{3}=A_{31}\cup A_{32}$ with
-$A_{31}=A_{3}\cap\{y||y|>t\},\qquad\qquad A_{32}=A_{3}\cap\{y||y|\leq t\}.$
+$A_{31}=A_{3}\cap\{y;|y|>t\},\qquad\qquad A_{32}=A_{3}\cap\{y;|y|\leq t\}.$
 In $A_{31}$, we have $t^{2}+|x-y|^{2}\sim|y|^{2}$ and
@@
 $|x|^{2}\leq|X|^{2}=t^{2}+|x|^{2}\leq\frac{5}{4}|x|^{2}.$
-In $A_{1}$, we have $\frac{|x|}{2}\leq|x-y|\leq\frac{3|x|}{2}$ and $t^{2}+|x-y|^{2}\sim|x|^{2}$, then,
+In $A_{1}$, we have $\frac{|x|}{2}\leq|x-y|\leq\frac{3|x|}{2}$ and $t^{2}+|x-y|^{2}\sim|x|^{2}$, then
 $|v(x,t)|\lesssim|x|^{-2\alpha}\int_{A_{1}}|f(y)|\mathrm{d}y\lesssim$$\begin{cases}|x|^{-2\alpha}|x|^{n-\min\{a,n\}}&\mathrm{if~{}~{}}a\neq n\\ |x|^{-2\alpha}\log(|x|+1)&\mathrm{if~{}~{}}a=n\end{cases}$
@@
 In $A_{2}$, we have $|x|/2<|y|<3|x|/2$. Further decompose $A_{2}=A_{21}\cup A_{22}$ with
-$A_{21}=A_{2}\cap\{y||x-y|>t\}\qquad\mathrm{and}\qquad A_{22}=A_{2}\cap\{y||x-y|\leq t\}.$
+$A_{21}=A_{2}\cap\{y;|x-y|>t\}\qquad\mathrm{and}\qquad A_{22}=A_{2}\cap\{y;|x-y|\leq t\}.$
 For $A_{21}$, we have $t^{2}+|x-y|^{2}\sim|x-y|^{2}$ and
@@
 Lemma 2.2.
-Suppose $f\in L_{\mathrm{loc}}^{1}(\mathbb{R}^{n})$ with the decay that $f(x)=O(|x|^{-b})$ as $|x|\to\infty$ for some positive constant $b$. We define
-$w(x,t)=\int_{\mathbb{R}^{n}}\frac{t^{\beta}}{(t^{2}+|x-y|^{2})^{\frac{n+\beta}{2}}}f(y)\mathrm{d}y,\quad\beta\in\mathbb{R}_{+}.$
+Suppose $f\in L_{\mathrm{loc}}^{1}(\mathbb{R}^{n})$ satisfies that $f(x)=O(|x|^{-b})$ as $|x|\to\infty$ for some $b\in\mathbb{R}_{+}$. Let
+$w(x,t)=\int_{\mathbb{R}^{n}}\frac{t^{\beta}f(y)}{(t^{2}+|x-y|^{2})^{\frac{n+\beta}{2}}}\mathrm{d}y,\quad\beta\in\mathbb{R}_{+}.$
 Then there holds
-$\left|w(x,t)\right|\lesssim$$\begin{cases}|X|^{-\min\{b,n\}}\qquad&\mathrm{if~{}~{}}b\neq n\\ |X|^{-n}\log|X|&\mathrm{if~{}~{}}b=n\end{cases},\qquad\quad\mathrm{~{}~{}as~{}~{}}\quad|X|\to\infty.$
+$\left|w(x,t)\right|\lesssim$$\begin{cases}|X|^{-\min\{b,n\}}\qquad&\mathrm{if~{}~{}}b\neq n\\ |X|^{-n}\log|X|&\mathrm{if~{}~{}}b=n\end{cases}\qquad\quad\mathrm{~{}~{}as~{}~{}}\quad|X|\to\infty.$
 Since the same trick works for the above second type singular integral, we omit the proof.
 For $r\in\mathbb{R}_{+}$, let $B_{r}^{+}=B_{r}(0)\cap\mathbb{R}^{n+1}_{+}$ and $\partial^{+}B_{r}^{+}=\partial B_{r}^{+}\cap\mathbb{R}^{n+1}_{+}$.
-Let $P_{k}(x)$ denote a polynomial with degree $k$ in the variable $x$, where $k\in\mathbb{N}$.
+Let $P_{k}(x)$ denote a polynomial with degree $\leq k$, where $k\in\mathbb{N}$.
 Given $u\in C^{\infty}(\overline{\mathbb{R}^{n+1}_{+}})$, we introduce a Kelvin transform for $\Delta^{2}$ by
-$u^{\ast}(X)=|X|^{3-n}u\left(\frac{X}{|X|^{2}}\right),\qquad X\in\overline{\mathbb{R}_{+}^{n+1}}\backslash\{0\}.$
-For our purpose, we need to refine Sun-Xiong’s results [30, Proposition 3.1 and Theorem 3.2] on homogeneous biharmonic boundary value problems.
+$u^{\ast}(X)=|X|^{3-n}u\left(X|X|^{-2}\right),\qquad X\in\overline{\mathbb{R}_{+}^{n+1}}\backslash\{0\}.$
+For our purpose, we need to a refinement of Sun-Xiong’s results [36, Proposition 3.1 and Theorem 3.2] for homogeneous biharmonic BVPs on $\mathbb{R}_{+}^{n+1}$.
 Lemma 2.3.
 Let $u\in C^{\infty}(\overline{\mathbb{R}^{n+1}_{+}})$ be a solution of
-$\begin{cases}\displaystyle\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{R}_{+}^{n+1},\\ \displaystyle\partial_{t}u(x,0)=0&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta u(x,0)=0\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (2.1)
+$\begin{cases}\displaystyle~{}\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\quad~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle~{}~{}~{}\partial_{t}u=0&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta u=0\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (2.1)
 Suppose (i) $\lim_{|x|\to\infty}u(x,0)=0$; (ii) there exist some $\delta,C\in\mathbb{R}_{+}$ such that
@@
 Then, $u(x,t)=c_{2}t^{2}$ for some nonnegative constant $c_{2}$.
+For our purpose, the nonnegativity assumption of solutions in Sun-Xiong [36] is redundant.
 Proof..
-It is notable that the nonnegativity of solutions is removed compared with that of Sun-Xiong [30].
 The proof consists of three steps.
 Step 1. $u^{\ast}\in L^{1}(B_{1}^{+})$.
-First, rewrite the boundary conditions as $\mathscr{B}_{1}^{3}u=0,\mathscr{B}_{3}^{3}u=0$, then it follows from conformal covariance that $u^{\ast}$ satisfies the biharmonic boundary value problem (2.1) in $\overline{\mathbb{R}_{+}^{n+1}}\backslash\{0\}$.
+First, rewrite the boundary conditions as $\mathscr{B}_{1}^{3}u=0,\mathscr{B}_{3}^{3}u=0$.
+Then the conformal covariance of $\mathscr{B}_{1}^{3}$ and $\mathscr{B}_{3}^{3}$ yields that $u^{\ast}$ satisfies problem (2.1) in $\overline{\mathbb{R}_{+}^{n+1}}\backslash\{0\}$.
 Define
 $\eta_{\varepsilon}(t)=\begin{cases}\displaystyle\frac{1}{24}(t-\varepsilon)^{4}\qquad&\mathrm{for}\qquad t\geq\varepsilon,\\ \displaystyle 0\qquad&\mathrm{for}\qquad 0\leq t<\varepsilon.\end{cases}$
-Notice that $\eta_{\varepsilon}\in C^{3,1}(\mathbb{R}^{n+1}_{+})$ and $u^{*}$ is biharmonic on +$\overline{\mathbb{R}^{n+1}_{+}}\backslash\{0\}$, then integrating by parts gives
+Notice that $\eta_{\varepsilon}\in C^{3,1}(\mathbb{R}^{n+1}_{+})$ and $u^{*}$ is biharmonic on +$\overline{\mathbb{R}^{n+1}_{+}}\backslash\{0\}$. Then integrating by parts gives
 $\int_{B_{1}^{+}}u^{*}(X)\Delta^{2}\eta_{\varepsilon}\mathrm{d}X$
@@
 Step 2. $u(X)=o(|X|^{4})$ as $|X|\to\infty$.
-For each $r\in\mathbb{R}_{+}$, we set $v_{r}(X)=u^{*}(rX)$, which is biharmonic in +$\overline{\mathbb{R}^{n+1}_{+}}\backslash\{0\}$.
-Now we extend $v_{r}$ by an even reflection to $\mathbb{R}^{n+1}\backslash\{0\}$ and apply the local estimates for biharmonic functions in [27, Proposition 4] to obtain
+For $r\in\mathbb{R}_{+}$, we set $v_{r}(X)=u^{*}(rX)$, which is biharmonic in +$\overline{\mathbb{R}^{n+1}_{+}}\backslash\{0\}$.
+Now we extend $v_{r}$ by an even reflection to $\mathbb{R}^{n+1}\backslash\{0\}$ and apply the local estimates for biharmonic functions (see e.g., [33, Proposition 4]) to obtain
 $\|v_{r}\|_{L^{\infty}(B^{+}_{3/4}\backslash B^{+}_{1/2})}\lesssim\|v_{r}\|_{L^{1}(B^{+}_{5/4}\backslash B^{+}_{1/4})}\lesssim r^{-(n+1)}\|u^{*}\|_{L^{1}(B^{+}_{5r/4}\backslash B^{+}_{r/4})}.$
@@
 $\partial_{t}^{3}u(x,0)=\partial_{t}\Delta u(x,0)-\partial_{t}\overline{\Delta}u(x,0)=0,$
-so $\tilde{u}\in C^{4}(\mathbb{R}^{n+1})$ and is biharmonic in $\mathbb{R}^{n+1}$. By [27, Proposition 4] we obtain
+so $\tilde{u}\in C^{4}(\mathbb{R}^{n+1})$ and is biharmonic in $\mathbb{R}^{n+1}$. By [33, Proposition 4] we obtain
 $\|\nabla^{4}\tilde{u}\|_{L^{\infty}(B_{R})}\leq\frac{C}{R^{4}}\frac{1}{|B_{2R}|}\int_{B_{2R}}|\tilde{u}|\mathrm{d}X\leq\frac{o(R^{4})}{R^{4}}\to 0\qquad\mathrm{as}\qquad R\to+\infty.$
-Thus, we have
+Thus we have
 $u(x,t)=\sum_{i=0}^{3}t^{i}P_{3-i}(x).$
@@
 Remark 2.1.
-If we relax the above assumption (2.2) to $u(X)\geq-C|X|^{\delta}$ for some constant $40<\delta<$ 4, which ensures that $u^{\ast}_{-}\in L^{1}(B_{1}^{+})$, then we apply the argument above to show that
+If we relax the above (2.2) to the assumption that $u(X)\geq-C|X|^{\delta}$ for some constant $40<\delta<$ 4, which ensures that $u^{\ast}_{-}\in L^{1}(B_{1}^{+})$, then the above argument yields
 $u(x,t)=t^{2}P_{1}(x).$
@@
 Let $u\in C^{\infty}(\overline{\mathbb{R}^{n+1}_{+}})$ be a solution of
-$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle u(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\Delta u(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (2.4)
+$\begin{cases}\displaystyle\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\qquad~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\quad~{}u=0&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle~{}~{}\Delta u=0\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (2.4)
 Suppose there exist some $\delta,C\in\mathbb{R}_{+}$ such that
-$u(X)\geq-C|X|^{-\delta}\qquad\mathrm{as}\quad|X|\to\infty.$ (2.5)
+$u(X)\geq-C|X|^{-\delta}\quad\mathrm{as}\quad|X|\to\infty.$ (2.5)
 Then,
@@
 Proof..
-Similarly, the boundary conditions become $\mathscr{B}_{0}^{3}u=0,\mathscr{B}_{3}^{3}u=0$ on $\partial\mathbb{R}_{+}^{n+1}$.
-By conformal covariance, $u^{\ast}$ satisfies the biharmonic boundary value problem (2.4) on +$\overline{\mathbb{R}^{n+1}_{+}}\backslash\{0\}$.
+Notice that these boundary conditions are equivalent to $\mathscr{B}_{0}^{3}u=0,\mathscr{B}_{3}^{3}u=0$ on $\partial\mathbb{R}_{+}^{n+1}$.
+By conformal covariance, $u^{\ast}$ satisfies problem (2.4) on +$\overline{\mathbb{R}^{n+1}_{+}}\backslash\{0\}$.
 Following the same lines as before, we have $u^{\ast}\in L^{1}(B_{1}^{+})$ and $u(x,t)=o(|X|^{4})$ as $|X|\to\infty$.
-Finally, we extend $u$ by an odd reflection
-$\tilde{u}(x,t)=\begin{cases}\displaystyle u(x,t)\qquad&\mathrm{for}\qquad t\geq 0,\\ \displaystyle-u(x,-t)\qquad&\mathrm{for}\qquad t<0.\end{cases}$
-A similar argument as in Lemma 2.3 combined with boundary conditions shows
+Finally, we extend $u$ to $\mathbb{R}^{n+1}$ by an odd reflection
+$\tilde{u}(x,t)=\begin{cases}\displaystyle u(x,t)\qquad&\mathrm{~{}~{}for~{}~{}}\quad t\geq 0,\\ \displaystyle-u(x,-t)\qquad&\mathrm{~{}~{}for~{}~{}}\quad t<0.\end{cases}$
+A similar argument as in Lemma 2.3 combined with the boundary conditions shows
 $u(x,t)=tP_{2}(x)+c_{3}t^{3}.$
@@
 Remark 2.2.
-If we relax the above assumption (2.5) to $u(X)\geq-C|X|^{\delta}$ as $|X|\to\infty$ for some $30<\delta<$ 3, then the same result in Lemma 2.4 still holds.
+If we relax the above assumption (2.5) to $u(X)\geq-C|X|^{\delta}$ as $|X|\to\infty$ for some $30<\delta<$ 3, then the same result in Lemma 2.4 is true.
 3 Biharmonic Poisson kernel
-Our purpose is to derive explicit formulae of biharmonic Poisson kernels for $\mathbb{R}_{+}^{n+1}$ and $\mathbb{B}^{n+1}$.
-3.1 Noncritical dimension: The upper half-space
-We start with the standard Poisson kernel for $-\Delta$ in $\mathbb{R}_{+}^{n+1}$:
+For a smooth compact Riemannian manifold $(X^{n+1},g)$ with boundary $M^{n}=\partial X$ and $n\geq 4$, under the assumption that a certain conformal invariant $\lambda_{1}(P_{4}^{g})>0$ (see [9]), the existence and uniqueness of smooth solutions to the following BVP was studied in [9, Proposition 1.3]: For $\bar{f}_{0},\bar{f}_{1}\in C^{\infty}(M)$,
+$\begin{cases}\displaystyle~{}P_{4}^{g}U=0&\mathrm{~{}~{}in~{}~{}}\qquad~{}X,\\ \displaystyle\mathscr{B}^{3}_{0}U=\bar{f}_{0}&\mathrm{~{}~{}on~{}~{}}\qquad M,\\ \displaystyle\mathscr{B}^{3}_{1}U=\bar{f}_{1}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\qquad M.\end{cases}$
+Our purpose is to derive explicit representation formulae of biharmonic Poisson kernels for $\mathbb{R}_{+}^{n+1}$ and $\mathbb{B}^{n+1}$.
+On $(\mathbb{B}^{n+1},\mathbb{S}^{n},|\mathrm{d}\xi|^{2})$, for each proper conformal boundary operator pair, such a biharmonic BVP is expected to be better understood, serving as a first step to the prescribing boundary curvature problem on manifolds with boundary.
+3.1 The half-space
+We start with the classical Poisson kernel for $-\Delta$ in $\mathbb{R}_{+}^{n+1}$:
 $P(x,t)=\frac{2}{|\mathbb{S}^{n}|}\frac{t}{(t^{2}+|x|^{2})^{\frac{n+1}{2}}}.$
@@
 Lemma 3.1.
-Suppose $f\in C^{2}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$ and $\alpha>1$, then for every fixed $x\in\mathbb{R}^{n}$,
+Suppose $f\in C^{2}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$ and $\alpha>1$. Then for every fixed $x\in\mathbb{R}^{n}$, we have
 $\lim_{t\searrow 0}\partial_{t}\int_{\mathbb{R}^{n}}\frac{t^{\alpha}}{\left(t^{2}+|x-y|^{2}\right)^{\frac{n+\alpha}{2}}}f(y)\mathrm{d}y=0.$
-It is not hard to see that the above result is true provided both $f$ and $|\nabla f|$ are bounded.
+It is not hard to see that the above result is still true provided both $f$ and $|\nabla f|$ are bounded. This elementary lemma will be used later.
 Proof..
@@
 $\int_{\mathbb{R}^{n}}\frac{f(x-ty)}{(1+|y|^{2})^{\frac{n+\alpha+2}{2}}}\mathrm{d}y=\frac{|\mathbb{S}^{n-1}|}{2}B(\frac{n}{2},\frac{\alpha+2}{2})f(x)+O(t^{\sigma(\alpha+2)})+O(t^{2}).$
-Therefore, with the above choice of $\sigma$ we combine these estimates to conclude that for $t\ll 1$,
+Therefore, with the choice of $\sigma$ we combine these estimates to conclude that for $t\ll 1$,
 $\partial_{t}\int_{\mathbb{R}^{n}}\frac{t^{\alpha}}{\left(t^{2}+|x-y|^{2}\right)^{\frac{n+\alpha}{2}}}f(y)\mathrm{d}y$
@@
 $u(x,t)=P\ast f(x,t)$
-solves $\Delta u=0$ in $\mathbb{R}_{+}^{n+1}$ and $u=f(x)$ on $\partial\mathbb{R}_{+}^{n+1}$. However, for each $x\notin\mathrm{supp}(f)$, the Hopf boundary point lemma yields $\partial_{t}u(x,0)>0$.
-Proof of Theorem 1.1. It suffices to show that fix $i,k\in\{0,1,2,3\}$, we define
+solves $\Delta u=0$ in $\mathbb{R}_{+}^{n+1}$ and $u=f(x)$ on $\partial\mathbb{R}_{+}^{n+1}$.
+However, for each $x\notin\mathrm{supp}(f)$, the Hopf boundary point lemma yields $\partial_{t}u>0$ on $\partial\mathbb{R}_{+}^{n+1}$.
+Proof of Theorem 1.1. It suffices to show that fix $i,k\in\{0,1,2,3\}$, let
 $[(\mathscr{B}_{i}^{3}P_{k}^{3})\ast f_{k}](x,t)=\int_{\mathbb{R}^{n}}(\mathscr{B}_{i}^{3}P_{k}^{3})(x-y,t)f_{k}(y)\mathrm{d}y,$
-then we have
-  • (i)
+then we need to show
+  1. (i)
     $\Delta^{2}P_{i}^{3}=0$ in $\mathbb{R}_{+}^{n+1}$.
-  • (ii)
+  2. (ii)
     For each pair $(i,k)$ other than $(0,2),(1,3)$, and fixed $x\in\mathbb{R}^{n}$ there hold
@@
 In the following, we prove assertion (ii) case by case.
-Case 1. Type $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{1})$ boundary condition.
+Case 1. $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{1})$-boundary condition.
 The kernel $P_{0}^{3}$ shares the common property as $P$ that
@@
 $\lim_{t\to 0}[(\mathscr{B}_{1}^{3}P_{1}^{3})\ast f_{1}](x,t)=-\lim_{t\to 0}[(\partial_{t}P_{1}^{3})\ast f_{1}](x,t)=\lim_{t\to 0}[(2P-P_{0}^{3})\ast f_{1}](x,t)=f_{1}(x).$
-Case 2. Type $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{2})$ boundary condition.
+Case 2. $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{2})$-boundary condition.
 Rewrite
@@
 $P_{0}^{3}(X)=\frac{2(n+1)}{|\mathbb{S}^{n}|}t^{3}|X|^{-3-n},\quad P_{2}^{3}(X)=-\frac{1}{(n-1)|\mathbb{S}^{n}|}t|X|^{1-n}.$
-It has been known in (3.2) that
+Also (3.2) yields
 $\lim_{t\to 0}(P_{0}^{3}\ast f_{0})(x,t)=f_{0}(x).$
@@
 A direct computation yields
-$\partial_{t}\Delta|X|^{3-n}=$ $(3-n)\Delta(t|X|^{1-n})=2(n-1)(n-3)t|X|^{-1-n};$
+$\partial_{t}\Delta|X|^{3-n}=$ $(3-n)\Delta(t|X|^{1-n})=2(n-1)(n-3)t|X|^{-1-n},$
 $2\overline{\Delta}\partial_{t}|X|^{3-n}=$ $2(3-n)\overline{\Delta}(t|X|^{1-n}).$
@@
 $\lim_{t\to 0}[(\mathscr{B}^{3}_{3}P_{3}^{3})\ast f_{3}](x,t)=$ $f_{3}(x)-\overline{\Delta}\lim_{t\to 0}[(P-P_{0}^{3})\ast f_{3}](x,t)=f_{3}(x).$ (3.6)
-Case 4. Type $(\mathscr{B}^{3}_{2},\mathscr{B}^{3}_{3})$ boundary condition.
+Case 4. $(\mathscr{B}^{3}_{2},\mathscr{B}^{3}_{3})$-boundary condition.
 Since we already have (3.5) and (3.6), it remains to show
@@
 $\lim_{t\to 0}[(\mathscr{B}^{3}_{3}P_{2}^{3})\ast f_{2}](x,t)=\overline{\Delta}\lim_{t\to 0}[tP\ast f_{2}(x,t)]=0.$
+For $n=3$, we only need to point out several significant differences. Observe that
+$\mathscr{B}^{3}_{3}P^{3}_{3}(X)=(\partial_{t}\Delta+2\overline{\Delta}\partial_{t})P^{3}_{3}(X)=\frac{8}{|\mathbb{S}^{3}|}\frac{t^{3}}{(t^{2}+|x|^{2})^{3}}=P^{3}_{0}(X),$
+so, $\lim_{t\to 0}[(\mathscr{B}^{3}_{3}P^{3}_{3})\ast f_{3}](x,t)=f_{3}(x)$. Also, we can apply
+$\mathscr{B}^{3}_{1}P^{3}_{3}(X)=$ $\frac{t}{2|\mathbb{S}^{3}|(t^{2}+|x|^{2})}=-P_{2}^{3}(X),$
+$\mathscr{B}^{3}_{2}P^{3}_{3}(X)=$ $\frac{2}{|\mathbb{S}^{3}|}\frac{t^{2}}{(t^{2}+|x|^{2})^{2}}=-P_{1}^{3}(X)$
+to show
+$\lim_{t\to 0}[(\mathscr{B}^{3}_{1}P^{3}_{3})\ast f_{3}](x,t)=0\qquad\mathrm{and}\qquad\lim_{t\to 0}[(\mathscr{B}^{3}_{2}P^{3}_{3})\ast f_{3}](x,t)=0.$
 $\Box$
-3.2 Noncritical dimension: The unit ball
-Our purpose is to give explicit representation formulae of biharmonic Poisson kernels for $\mathbb{B}^{n+1}$ with the help of conformal equivalence between $\mathbb{B}^{n+1}$ and $\mathbb{R}_{+}^{n+1}$,
+3.2 The unit ball
+Our purpose is to give explicit representation formulae of biharmonic Poisson kernels for $\mathbb{B}^{n+1}$ with the help of conformal equivalence between $\mathbb{B}^{n+1}$ and $\mathbb{R}_{+}^{n+1}$.
 Let $F:(\mathbb{R}^{n+1}_{+},|\mathrm{d}X|^{2})\to(\mathbb{B}^{n+1},|\mathrm{d}\xi|^{2})$
@@
 $F^{\ast}(|\mathrm{d}\xi|^{2})=\left(\frac{2}{|X+\mathbf{e}_{n+1}|^{2}}\right)^{2}|\mathrm{d}X|^{2}:=U_{0}(X)^{\frac{4}{n-3}}|\mathrm{d}X|^{2}.$
-Clearly, $F$ is exact an inversion of the sphere centered at $-\mathbf{e}_{n+1}$ of radius $\sqrt{2}$ and thus $F=F^{-1}$.
-The following important but elementary identity will be frequently used:
+Clearly, $F$ is an inversion of the sphere centered at $-\mathbf{e}_{n+1}$ of radius $\sqrt{2}$ and its inverse $F^{-1}=F$.
+The following elementary identity will be frequently used:
 $2|\xi+\mathbf{e}_{n+1}|^{-1}|\eta+\mathbf{e}_{n+1}|^{-1}|\xi-\eta|=|F^{-1}(\xi)-F^{-1}(\eta)|,\quad\forall~{}\xi,\eta\in\overline{\mathbb{B}^{n+1}}\setminus\{-\mathbf{e}_{n+1}\}.$ (3.8)
-This can be verified by a straightforward calculation.
-Given $U\in C^{\infty}(\overline{\mathbb{B}^{n+1}})$, we set
-$u=U\circ FU_{0}.$ (3.9)
-Clearly, $u(X)=O^{(k)}(|X|^{3-n})$ as $|X|\to\infty$, where $O^{(k)}(|X|^{3-n})$ means the $k$-th derivatives of $v$ belong to $O(|X|^{3-n-k})$ for every $k\in\mathbb{N}$.
+This follows by a straightforward calculation.
+For $U\in C^{\infty}(\overline{\mathbb{B}^{n+1}})$ we let
+$u=U_{0}(U\circ F).$ (3.9)
+Clearly, $u(X)=O^{(k)}(|X|^{3-n})$ as $|X|\to\infty$, where $O^{(k)}(|X|^{3-n})$ means the $k$-th derivatives of $u$ belong to $O(|X|^{3-n-k})$ for $k\in\mathbb{N}$.
 Moreover, by conformal covariance of the Paneitz operator, we have
-$\Delta^{2}u=P_{4}^{|\mathrm{d}X|^{2}}(u)=P_{4}^{|\mathrm{d}X|^{2}}(U\circ FU_{0})=$ $U_{0}^{\frac{n+5}{n-3}}P_{4}^{U_{0}^{\frac{4}{n-3}}|\mathrm{d}X|^{2}}(U\circ F)$
+$\Delta^{2}u=P_{4}^{|\mathrm{d}X|^{2}}(u)=P_{4}^{|\mathrm{d}X|^{2}}(U_{0}U\circ F)=$ $U_{0}^{\frac{n+5}{n-3}}P_{4}^{U_{0}^{\frac{4}{n-3}}|\mathrm{d}X|^{2}}(U\circ F)$
 $=$ $U_{0}^{\frac{n+5}{n-3}}(P_{4}^{|\mathrm{d}\xi|^{2}}U)\circ F=U_{0}^{\frac{n+5}{n-3}}(\Delta^{2}U)\circ F.$
 Similarly, by (1.4) we have
-$\left(\mathscr{B}_{k}^{3}\right)_{|\mathrm{d}X|^{2}}(u)=U_{0}^{\frac{n+2k-3}{n-3}}\left[\left(\mathscr{B}_{k}^{3}\right)_{|\mathrm{d}\xi|^{2}}U\right]\circ F\qquad\mathrm{~{}~{}for~{}~{}}\quad 0\leq k\leq 3.$ (3.10)
-Proof of Theorem 1.2. For assertion $(i)$, we first handle $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{2})$ type boundary condition.
+$\left(\mathscr{B}_{k}^{3}\right)_{|\mathrm{d}X|^{2}}(u)=U_{0}^{\frac{n+2k-3}{n-3}}\left[\left(\mathscr{B}_{k}^{3}\right)_{|\mathrm{d}\xi|^{2}}U\right]\circ F\qquad\mathrm{~{}~{}for~{}~{}}\quad k\in\{0,1,2,3\}.$ (3.10)
+Theorem 3.1.
+Suppose $n\geq 2$ and $n\neq 3$, and $\bar{f}_{i}\in C^{\infty}(\mathbb{S}^{n})$ for $i\in\{0,1,2,3\}$. For $\xi\in\mathbb{B}^{n+1}$ and $\eta\in\mathbb{S}^{n}$, we let
+$\bar{P}_{0}^{3}(\xi,\eta)=$ $\frac{n+1}{4|\mathbb{S}^{n}|}\frac{\left(1-|\xi|^{2}\right)^{3}}{|\xi-\eta|^{n+3}};$
+$\bar{P}_{1}^{3}(\xi,\eta)=$ $-\frac{1}{2|\mathbb{S}^{n}|}\frac{(1-|\xi|^{2})^{2}}{|\xi-\eta|^{n+1}};$
+$\bar{P}_{2}^{3}(\xi,\eta)=$ $-\frac{1}{2(n-1)|\mathbb{S}^{n}|}\frac{1-|\xi|^{2}}{|\xi-\eta|^{n-1}};$
+$\bar{P}_{3}^{3}(\xi,\eta)=$ $\frac{1}{(n-1)(n-3)|\mathbb{S}^{n}|}\frac{1}{|\xi-\eta|^{n-3}}.$
+Then for $i,j\in\{0,1,2,3\},i<j,i+j\neq 3$, the solution of the following BVP
+$\begin{cases}\displaystyle\Delta^{2}U=0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}U=\bar{f}_{i}&\mathrm{~{}~{}on~{}~{}}\quad\mathbb{S}^{n},\\ \displaystyle\mathscr{B}^{3}_{j}U=\bar{f}_{j}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\mathbb{S}^{n},\end{cases}$ (3.11)
+is unique and has the form of
+$U(\xi)=\int_{\mathbb{S}^{n}}\bar{P}_{i}^{3}(\xi,\eta)\bar{f}_{i}(\eta)\mathrm{d}V_{\mathbb{S}^{n}}(\eta)+\int_{\mathbb{S}^{n}}\bar{P}_{j}^{3}(\xi,\eta)\bar{f}_{j}(\eta)\mathrm{d}V_{\mathbb{S}^{n}}(\eta).$
+Proof..
+We first handle $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{2})$ type boundary condition.
 Let $u$ define as in (3.9), equivalently,
@@
 By virtue of (3.10) we set
-$f_{0}(x)=\left(\frac{2}{1+|x|^{2}}\right)^{\frac{n-3}{2}}\bar{f}_{0}\circ F\left((x,0)\right)\quad\mathrm{and}\quad f_{2}(x)=\left(\frac{2}{1+|x|^{2}}\right)^{\frac{n+1}{2}}\bar{f}_{2}\circ F((x,0)).$
+$f_{0}(x)=\left(\frac{2}{1+|x|^{2}}\right)^{\frac{n-3}{2}}\bar{f}_{0}\circ F(x,0)\quad\mathrm{and}\quad f_{2}(x)=\left(\frac{2}{1+|x|^{2}}\right)^{\frac{n+1}{2}}\bar{f}_{2}\circ F(x,0).$
 Then for $|X|,|x|\gg 1$, we know
@@
 and $u$ solves
-$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle u(x,0)=f_{0}\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial^{2}_{t}u(x,0)-\overline{\Delta}u(x,0)=f_{2}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (3.11)
-As a direct consequence of Remark 2.2, the solution to the above problem (3.11) is unique. This together with Theorem 1.1 gives
+$\begin{cases}\displaystyle\qquad~{}\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\qquad~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\qquad\quad~{}~{}u=f_{0}&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial^{2}_{t}u-\overline{\Delta}u=f_{2}\qquad\quad&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (3.12)
+As a direct consequence of Remark 2.2, the above problem (3.12) admits a unique solution. This together with Theorem 1.1 gives
 $u(X)=\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}f_{0}(y)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y-\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{tf_{2}(y)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y.$
@@
 $\xi=$ $F(X)=\left(\frac{2x}{(1+t)^{2}+|x|^{2}},\frac{1-t^{2}-|x|^{2}}{(1+t)^{2}+|x|^{2}}\right),$
-$\eta=$ $F((y,0))=\left(\frac{2y}{1+|y|^{2}},\frac{1-|y|^{2}}{1+|y|^{2}}\right),$
+$\eta=$ $F(y,0)=\left(\frac{2y}{1+|y|^{2}},\frac{1-|y|^{2}}{1+|y|^{2}}\right),$
 then
 $t\circ F^{-1}=\frac{1-|\xi|^{2}}{|\xi+\mathbf{e}_{n+1}|^{2}}$
-and from (3.8) that
+and
+$|\xi+\mathbf{e}_{n+1}|^{2}=\frac{4}{(1+t)^{2}+|x|^{2}},\quad|\eta+\mathbf{e}_{n+1}|^{2}=\frac{4}{1+|y|^{2}}.$
+Moreover, it follows from (3.8) that
 $2|\xi+\mathbf{e}_{n+1}|^{-1}|\eta+\mathbf{e}_{n+1}|^{-1}|\xi-\eta|=|(x,t)-(y,0)|.$
@@
 $=$ $\frac{(n+1)\left(1-|\xi|^{2}\right)^{3}}{4|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\bar{f}_{0}(\eta)}{|\xi-\eta|^{n+3}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta)-\frac{(1-|\xi|^{2})}{2(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\bar{f}_{2}(\eta)}{|\xi-\eta|^{n-1}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta).$
-As indicated in the preceding discussion, for other pairs of conformal boundary operators it suffices to show the uniqueness.
-The uniqueness for $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{1})$ is trivial.
+As indicated in the preceding discussion, for other proper conformal boundary operator pairs it remains to show the uniqueness.
+The uniqueness for $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{1})$ directly follows by an integration by parts.
 For $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{3})$, the uniqueness follows from a similar argument above combined with Lemma 2.3 or Remark 2.1.
@@
 and let $u$ define as in (3.9), then $u$ solves
-$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\partial_{t}^{2}u(x,0)-\overline{\Delta}u(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta u(x,0)+2\overline{\Delta}\partial_{t}u(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\end{cases}$
+$\begin{cases}\displaystyle\qquad\qquad~{}\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\qquad~{}~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\qquad\partial_{t}^{2}u-\overline{\Delta}u=0&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta u+2\overline{\Delta}\partial_{t}u=0\qquad\quad&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+},\end{cases}$
 with $u(X)=O(|X|^{3-n})$ as $|X|\to\infty$.
@@
 $\lim_{t\to 0}\frac{2t}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t\partial_{t}u(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\mathrm{d}y=0.$
-So, by the uniqueness for $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{2})$ due to Lemma 2.4 we obtain $u=0$ and $U=0$ as well.
-For assertion $(ii)$, the multiplicity of smooth solutions for the remaining two boundary conditions is a direct consequence of the nex Lemma 3.2. $\Box$
+So, by the uniqueness for $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{2})$ due to Lemma 2.4 we obtain $u=0$ and $U=0$ as well. $\Box$
+Remark 3.2.
+It is notable that restricting $\xi$ to $\mathbb{S}^{n}$, $2\bar{P}_{3}^{3}(\xi,\eta)$ coincides with the Green function for the fractional GJMS operator $P_{3}$ of order three on $\mathbb{S}^{n}$ (cf.
+the first author and Shi [13, Theorem 1(3)]).
 Lemma 3.2.
-The following homogeneous biharmonic boundary value problems:
-$\begin{cases}\displaystyle\Delta^{2}U=0\qquad&\mathrm{in}\qquad\mathbb{B}^{n+1}\\ \displaystyle\mathscr{B}^{3}_{0}(U)=0\qquad&\mathrm{on}\qquad\mathbb{S}^{n}\\ \displaystyle\mathscr{B}^{3}_{3}(U)=0\qquad&\mathrm{on}\qquad\mathbb{S}^{n}\end{cases}$ (3.12)
+The following homogeneous biharmonic BVPs:
+$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}U=0&\mathrm{in}\qquad\mathbb{B}^{n+1}\\ \displaystyle\mathscr{B}^{3}_{0}(U)=0&\mathrm{on}\qquad\mathbb{S}^{n}\\ \displaystyle\mathscr{B}^{3}_{3}(U)=0\qquad\quad&\mathrm{on}\qquad\mathbb{S}^{n}\end{cases}$ (3.13)
 and
-$\begin{cases}\displaystyle\Delta^{2}U=0\qquad&\mathrm{in}\qquad\mathbb{B}^{n+1}\\ \displaystyle\mathscr{B}^{3}_{1}(U)=0\qquad&\mathrm{on}\qquad\mathbb{S}^{n}\\ \displaystyle\mathscr{B}^{3}_{2}(U)=0\qquad&\mathrm{on}\qquad\mathbb{S}^{n}\end{cases}$ (3.13)
+$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}U=0&\mathrm{in}\qquad\mathbb{B}^{n+1}\\ \displaystyle\mathscr{B}^{3}_{1}(U)=0&\mathrm{on}\qquad\mathbb{S}^{n}\\ \displaystyle\mathscr{B}^{3}_{2}(U)=0\qquad\quad&\mathrm{on}\qquad\mathbb{S}^{n}\end{cases}$ (3.14)
 have nontrivial smooth solutions.
@@
 for the boundary condition $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{2})$.
-As in the proof of Theorem 1.1, we know that $u=u_{\varphi}$ solves
-$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1}\\ \displaystyle u(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\\ \displaystyle\partial_{t}\Delta u(x,0)+2\overline{\Delta}\partial_{t}u(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\end{cases}$
+As in the proof of Theorem 1.1, we know that $u=u_{\varphi}$ satisfies both
+$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1}\\ \displaystyle\quad~{}u=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\\ \displaystyle\partial_{t}\Delta u+2\overline{\Delta}\partial_{t}u=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\end{cases}$
 and
-$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1}\\ \displaystyle\partial_{t}u(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\\ \displaystyle\partial^{2}_{t}u(x,0)-\overline{\Delta}u(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\end{cases},$
+$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1}\\ \displaystyle~{}~{}\partial_{t}u=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\\ \displaystyle\partial^{2}_{t}u-\overline{\Delta}u=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\end{cases},$
 respectively. Finally, we may obtain nontrivial smooth solutions of (3.13) and (3.14) via
@@
 $\Box$
-This also reflects the loss of ‘the complementing condition’ for $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{3})$ and $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{2})$, under which boundary conditions problem (3.12) or (3.13) fails to have the uniqueness of the solution as explained in [1].
-3.3 Critical dimension
-Motivated by Chen-Shi [10, Theorem 1(1)] we are able to derive explicit formulae of biharmonic Poisson kernels on $\mathbb{R}_{+}^{4}$ and $\mathbb{B}^{4}$.
-Proof of Theorem 1.4. Based on the proof of Theorems 1.1 and 1.2, it suffices to point out several significant differences.
-(1) Observe that
-$\mathscr{B}^{3}_{3}P^{3}_{3}(X)=(\partial_{t}\Delta+2\overline{\Delta}\partial_{t})P^{3}_{3}(X)=\frac{8}{|\mathbb{S}^{3}|}\frac{t^{3}}{(t^{2}+|x|^{2})^{3}}=P^{3}_{0}(X),$
-so, $\lim_{t\to 0}[(\mathscr{B}^{3}_{3}P^{3}_{3})\ast f_{3}](x,t)=f_{3}(x)$. Also, we can apply
-$\mathscr{B}^{3}_{1}P^{3}_{3}(X)=$ $\frac{t}{2|\mathbb{S}^{3}|(t^{2}+|x|^{2})}=-P_{2}^{3}(X),$
-$\mathscr{B}^{3}_{2}P^{3}_{3}(X)=$ $\frac{2}{|\mathbb{S}^{3}|}\frac{t^{2}}{(t^{2}+|x|^{2})^{2}}=-P_{1}^{3}(X)$
-to show
-$\lim_{t\to 0}[(\mathscr{B}^{3}_{1}P^{3}_{3})\ast f_{3}](x,t)=0\qquad\mathrm{and}\qquad\lim_{t\to 0}[(\mathscr{B}^{3}_{2}P^{3}_{3})\ast f_{3}](x,t)=0.$
-(2) Suppose $U$ is a solution of (1.7). In critical dimension four, we introduce
+This also reflects the loss of ‘the complementing condition’ for $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{3})$ and $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{2})$.
+Due to this reason, problem (3.13) or (3.14) fails to have the uniqueness.
+Motivated by Shi and the first author [13, Theorem 1(1)] we are able to derive explicit formulae of biharmonic Poisson kernels for $\mathbb{R}_{+}^{4}$ and $\mathbb{B}^{4}$.
+Theorem 3.2.
+Suppose $\bar{f}_{i}\in C^{\infty}(\mathbb{S}^{3})$ for $i\in\{0,1,2,3\}$.
+Then the biharmonic Poisson kernels $\bar{P}_{i}^{3}$ for $\mathbb{B}^{4}$ have the same representation formulae as in Theorem 3.1 except for
+$\bar{P}_{3}^{3}(\xi,\eta)=-\frac{1}{2|\mathbb{S}^{3}|}\log|\xi-\eta|+C,\qquad C\in\mathbb{R}$
+where $\xi\in\mathbb{B}^{4},\eta\in\mathbb{S}^{3}$. Moreover, we have
+  1. (i)
+    $i=1,2$ and $j=3$, the solution of (3.11) is
+    $U(\xi)-\frac{1}{|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}U\mathrm{d}V_{{\mathbb{S}^{3}}}$
+    $=$ $\int_{\mathbb{S}^{3}}\bar{P}^{3}_{i}(\xi-\eta)\bar{f}_{i}(\eta)\mathrm{d}V_{{\mathbb{S}^{3}}}(\eta)-\frac{1}{2|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}\log|\xi-\eta|\bar{f}_{3}(\eta)\mathrm{d}V_{\mathbb{S}^{3}}(\eta),\quad\forall~{}\xi\in\overline{\mathbb{B}^{4}}.$
+  2. (ii)
+    $(i,j)=(0,1)\mathrm{~{}or~{}}(0,2)$, the unique solution of (3.11) is
+    $U(\xi)=\int_{\mathbb{S}^{3}}\bar{P}^{3}_{i}(\xi-\eta)\bar{f}_{i}(\eta)+\bar{P}^{3}_{j}(\xi-\eta)\bar{f}_{j}(\eta)\mathrm{d}V_{{\mathbb{S}^{3}}}(\eta),\quad\forall~{}\xi\in\overline{\mathbb{B}^{4}}.$
+Proof..
+Suppose $U$ is a smooth solution of (3.11). For $n=3$, we introduce
 $u(x,t)=U\circ F(x,t)+\log\frac{2}{(1+t)^{2}+|x|^{2}},\qquad\hat{U}_{0}(x,t):=\log\frac{2}{(1+t)^{2}+|x|^{2}}$
@@
 $f_{0}(x)=$ $\bar{f}_{0}\circ F(x,0)+\hat{U}(x,0),$
-$f_{k}(x)=$ $\left[\bar{f}_{k}\circ F(x,0)+(T_{k}^{3})_{g_{\mathbb{S}^{3}}}\right]e^{k\hat{U}_{0}(x,0)},k\in\{1,2,3\}.$
-This essentially follows from conformal covariance of $P_{4}^{g}$ and $\mathscr{B}_{k}^{3}$ that
+$f_{k}(x)=$ $\left[\bar{f}_{k}\circ F(x,0)+(T_{k}^{3})_{\mathbb{S}^{3}}\right]e^{k\hat{U}_{0}(x,0)},k\in\{1,2,3\}.$
+By the conformal covariance of $P_{4}^{g}$ and $\mathscr{B}_{k}^{3}$ we have
 $e^{2u}|\mathrm{d}X|^{2}=F^{\ast}(e^{2U}|\mathrm{d}\xi|^{2}),\quad e^{2\hat{U}_{0}}|\mathrm{d}X|^{2}=F^{\ast}(|\mathrm{d}\xi|^{2})$
@@
 […]
-$=$ $e^{k\hat{U}_{0}}(\mathscr{B}_{k}^{3})_{F^{\ast}(|\mathrm{d}\xi|^{2})}(U\circ F)+(T_{k}^{3})_{g_{\mathbb{S}^{3}}}e^{k\hat{U}_{0}}$
-$=$ $e^{k\hat{U}_{0}}[(\mathscr{B}_{3}^{3})_{|\mathrm{d}\xi|^{2}}(U)]\circ F+(T_{k}^{3})_{g_{\mathbb{S}^{3}}}e^{k\hat{U}_{0}}$
-$=$ $e^{k\hat{U}_{0}}(\bar{f}_{k}\circ F+(T_{k}^{3})_{g_{\mathbb{S}^{3}}}).$
-In particular, $(T_{1}^{3})_{g_{\mathbb{S}^{3}}}=1,(T_{2}^{3})_{g_{\mathbb{S}^{3}}}=2,(T_{3}^{3})_{g_{\mathbb{S}^{3}}}=4$.
-Suppose $u$ is a solution of (1.6). By the biharmonic Poisson kernel on $\mathbb{R}_{+}^{4}$ we have
+$=$ $e^{k\hat{U}_{0}}(\mathscr{B}_{k}^{3})_{F^{\ast}(|\mathrm{d}\xi|^{2})}(U\circ F)+(T_{k}^{3})_{\mathbb{S}^{3}}e^{k\hat{U}_{0}}$
+$=$ $e^{k\hat{U}_{0}}[(\mathscr{B}_{3}^{3})_{|\mathrm{d}\xi|^{2}}(U)]\circ F+(T_{k}^{3})_{\mathbb{S}^{3}}e^{k\hat{U}_{0}}$
+$=$ $e^{k\hat{U}_{0}}(\bar{f}_{k}\circ F+(T_{k}^{3})_{\mathbb{S}^{3}}).$
+In particular, $(T_{1}^{3})_{\mathbb{S}^{3}}=1,(T_{2}^{3})_{\mathbb{S}^{3}}=2,(T_{3}^{3})_{\mathbb{S}^{3}}=4$.
+Suppose $u$ is a solution of (1.6). Using the biharmonic Poisson kernel for $\mathbb{R}_{+}^{4}$ we have
 $u(x,t)=$ $\int_{\partial\mathbb{R}_{+}^{4}}[P_{i}^{3}(x-y,t)f_{i}(y)+P_{j}^{3}(x-y,t)f_{j}(y)]\mathrm{d}y$
@@
 where
-$I(x,t)=\begin{cases}\int_{\partial\mathbb{R}_{+}^{4}}[P_{0}^{3}(x-y,t)\hat{U}_{0}(y,0)+P_{j}^{3}(x-y,t)e^{j\hat{U}_{0}(y,0)}(T_{j}^{3})_{g_{\mathbb{S}^{3}}}]\mathrm{d}y,\\ \int_{\partial\mathbb{R}_{+}^{4}}[P_{i}^{3}(x-y,t)e^{i\hat{U}_{0}(y,0)}f_{i}\circ F(y)+P_{j}^{3}(x-y,t)e^{j\hat{U}_{0}(y,0)}f_{j}\circ F(y)]\mathrm{d}y\quad\mathrm{~{}~{}if~{}~{}}i\neq 0.\end{cases}$
+$I(x,t)=\begin{cases}\displaystyle\int_{\partial\mathbb{R}_{+}^{4}}[P_{0}^{3}(x-y,t)\hat{U}_{0}(y,0)+P_{j}^{3}(x-y,t)e^{j\hat{U}_{0}(y,0)}(T_{j}^{3})_{\mathbb{S}^{3}}]\mathrm{d}y,\\ \displaystyle\int_{\partial\mathbb{R}_{+}^{4}}[P_{i}^{3}(x-y,t)e^{i\hat{U}_{0}(y,0)}f_{i}\circ F(y)+P_{j}^{3}(x-y,t)e^{j\hat{U}_{0}(y,0)}f_{j}\circ F(y)]\mathrm{d}y~{}~{}\mathrm{~{}~{}if~{}~{}}i\neq 0.\end{cases}$
 Then it is not hard to see that $I=\hat{U}_{0}$ when $(i,j)=(0,1),(0,2)$; $I=\hat{U}_{0}+C,C\in\mathbb{R}$ when $(i,j)=(1,3),(2,3)$.
-Hence, using variable changes as in the proof of Theorem 1.2 we obtain
+Hence, using the variable changes as in the proof of Theorem 3.1 we obtain
 $U(\xi)=\int_{\mathbb{S}^{3}}\bar{P}^{3}_{i}(\xi-\eta)\bar{f}_{i}(\eta)+\bar{P}^{3}_{j}(\xi-\eta)\bar{f}_{j}(\eta)\mathrm{d}V_{{\mathbb{S}^{3}}}(\eta)$
@@
 when $(i,j)=(1,3),(2,3)$.
-It remains to determine the constant $C$ in the following way: On $\mathbb{S}^{3}$, a compatible condition of $\bar{f}_{3}$ arises from the fact that $\mathrm{ker}P_{3}=\mathbb{R}$ by virtue of [10, Theorem 1(1)], more explicitly,
+It remains to determine the constant $C$.
+On $\mathbb{S}^{3}$, a compatible condition of $\bar{f}_{3}$ arises from the fact that $\mathrm{ker}P_{3}=\mathbb{R}$ by virtue of [13, Theorem 1(1)], more explicitly,
 $\int_{\mathbb{S}^{3}}\bar{f}_{3}\mathrm{d}V_{\mathbb{S}^{3}}=\int_{\mathbb{S}^{3}}(-\frac{\partial\Delta U}{\partial r}-2\Delta_{\mathbb{S}^{3}}\frac{\partial U}{\partial r}-2\Delta_{\mathbb{S}^{3}}U)\mathrm{d}V_{\mathbb{S}^{3}}=0.$ (3.16)
@@
 $U(\xi)=-\frac{1}{2|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}\log|\xi-\eta|\bar{f}_{3}(\eta)\mathrm{d}V_{\mathbb{S}^{3}}+C.$
-Integrating the above equation over $\mathbb{S}^{3}$ to give
+Integrating the above equation over $\mathbb{S}^{3}$ gives
 $\int_{\mathbb{S}^{3}}U\mathrm{d}V_{\mathbb{S}^{3}}=-\frac{1}{2|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}\bar{f}_{3}(\eta)\mathrm{d}V_{\mathbb{S}^{3}}(\eta)\int_{\mathbb{S}^{3}}\log|\xi-\eta|\mathrm{d}V_{\mathbb{S}^{3}}(\xi)+C|\mathbb{S}^{3}|=C|\mathbb{S}^{3}|,$
-where we have used (3.15) and the fact that $\int_{\mathbb{S}^{3}}\log|\xi-\eta|\mathrm{d}V_{\mathbb{S}^{3}}(\xi),\forall~{}\eta\in\mathbb{S}^{3}$ is constant. This follows that
+where we have used (3.16) and the fact that $\forall~{}\eta\in\mathbb{S}^{3}$, $\int_{\mathbb{S}^{3}}\log|\xi-\eta|\mathrm{d}V_{\mathbb{S}^{3}}(\xi)$ is constant. This follows that
 $C=\frac{1}{|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}U\mathrm{d}V_{{\mathbb{S}^{3}}}.$
-Then the desired assertion follows by inserting the above constant into (3.14). $\Box$
-It is not hard to see that Theorem 1.4 could give a direct proof of [6, Proposition 2.2, Lemma 2.3]).
-We point out that the proof of Theorems 1.1 and 1.2 is still valid for the case of $n=2$. However, we shall not go deep into this, since it has been well studied by authors [11].
-Thanks to biharmonic Poisson kernels on $\mathbb{R}^{n+1}$, we can recover a special case in [7, Theorem 1.4], which clarifies the relationship between extrinsic and intrinsic GJMS operators, and leave its proof to Appendix A for clarity.
-Something about $P_{1}$ and $P_{3}$.
-Proposition 3.1 (cf. [7, Theorem 1.4] and [11, Porposition 2.2]).
+Then the desired assertion follows by substituting the above constant into (3.15). $\Box$
+Finally, we point out that the proof of Theorems 1.1 and 3.1 is still valid for $n=2$. However, we shall not go deep into this, since it has been well studied by the authors [15].
+For a Poincaré-Einstein manfold $(X^{n+1},g_{+})$ with conformal infinity $(\partial X=M,[h])$, and assume $(n^{2}-1)/4,(n^{2}-9)/4\notin\sigma_{\mathrm{pp}}(-\Delta_{g_{+}})$, with the factorization of $P_{4}^{g_{+}}$ and explicit formulae of extrinsic GJMS boundary operators $\mathscr{B}_{k}^{3}$, Case [9] used the scattering operator to find fractional GJMS operators $P_{1}^{g},P_{3}^{g}$ from directly computing $(\mathscr{B}_{k}^{3})_{g}u$, where $P_{4}^{g}u=0$ in $X$ for $g=r^{2}g_{+}$ with $r$ being a geodesic defining function.
+As for the Poincaré ball $(\mathbb{B}^{n+1},g_{+}=(\frac{2}{1-|x|^{2}})^{2}|\mathrm{d}x|^{2})$, the spectrum $\sigma(-\Delta_{g_{+}})=[n^{2}/4,+\infty)$ and [9, Theorem 1.4] is applicable.
+Proposition 3.1.
 Suppose $U$ is a biharmonic function in $\mathbb{B}^{n+1}$ for $n\geq 2$, then there hold
 $\mathscr{B}^{3}_{3}U=2P_{3}U\qquad\mathrm{and}\qquad\mathscr{B}^{3}_{2}\left(U\right)=2P_{1}(\mathscr{B}^{3}_{1}U)\qquad\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n},$
-where $P_{1},P_{3}$ are the fractional GJMS operators with respect to the round metric, see [7, 10].
-Proposition 3.1 indicates that a Dirichlet datum is equivalent to a $\mathscr{B}_{3}^{3}$ datum since $P_{3}$ is invertible; a $\mathscr{B}_{1}^{3}$ datum is equivalent to a $\mathscr{B}_{2}^{3}$ datum since $P_{1}$ is invertible.
-That is exact the reason why we can construct multiple smooth solutions for a biharmonic equation on $\mathbb{B}^{n+1}$ with these two types of boundary conditions.
+where $P_{1},P_{3}$ are the first- and third-order GJMS operators with respect to the round metric, see [9, 13].
+An alternative proof of Proposition 3.1 using the biharmonic Poisson kernel is available in Appendix A.
 4 Biharmonic Green function
-For $n\geq 2$, we let $X=(x,t)\in\mathbb{R}_{+}^{n+1},Y=(y,s)\in\mathbb{R}_{+}^{n+1}$. We define $\bar{X}=(x,-t)$, and $\xi=F(X),\eta=F(Y)$ with $F$ being the conformal map as in (3.7).
-4.1 Noncritical dimension
-Recall that the fundamental solution of $\Delta^{2}$ in $\mathbb{R}^{n+1}$ is given by
+For $n\geq 2$, let $X=(x,t)\in\mathbb{R}_{+}^{n+1},Y=(y,s)\in\mathbb{R}_{+}^{n+1}$. We define $\bar{X}=(x,-t)$, and $\xi=F(X),\eta=F(Y)$ with $F$ being the conformal map as in (3.7).
+We first introduce a biharmonic Green function on $\mathbb{B}^{n+1}$ with each of four proper conformal boundary operator pairs.
+Consider an nonhomogeneous biharmonic BVP with vanishing boundary conditions
+$\begin{cases}\Delta^{2}U=\bar{f}\qquad\qquad&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{B}^{n+1},\\ \mathscr{B}_{i}^{3}U=0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{S}^{n},\\ \mathscr{B}_{j}^{3}U=0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{S}^{n}.\end{cases}$ (4.1)
+Definition 4.1 (Biharmonic Green function).
+Let $\bar{f}\in C^{\infty}(\overline{\mathbb{B}^{n+1}})$ and $i,j\in\{0,1,2,3\},i+j\neq 3$.
+A function $\bar{G}^{(i,j)}:(\xi,\eta)\in\overline{\mathbb{B}^{n+1}}\times\overline{\mathbb{B}^{n+1}}\mapsto\mathbb{R}\cup\{\infty\}$ for (4.1) is called a Green function, if $\bar{G}^{(i,j)}$ satisfies
+  • $\xi\mapsto\bar{G}^{(i,j)}(\xi,\eta)-\Gamma(\xi-\eta)\in C^{\infty}(\overline{\mathbb{B}^{n+1}})$ for all $\eta\in\mathbb{B}^{n+1}$;
+  • $\Delta_{\xi}^{2}\left(\bar{G}^{(i,j)}(\xi,\eta)-\Gamma(\xi-\eta)\right)=0$ for all $(\xi,\eta)\in\mathbb{B}^{n+1}\times\mathbb{B}^{n+1}$;
+  • $\mathscr{B}_{i}^{3}\bar{G}^{(i,j)}=0,\mathscr{B}_{j}^{3}\bar{G}^{(i,j)}=0$ on $\mathbb{S}^{n}$.
+Formally, the above Green function enables us to express a unique solution to (4.1) as
+$U(\xi)=\bar{G}^{(i,j)}\ast\bar{f}(\xi)=\int_{\mathbb{B}^{n+1}}\bar{G}^{(i,j)}(\xi,\eta)\bar{f}(\eta)\mathrm{d}\eta.$
+We notice that $\bar{G}^{(0,1)}$ coincides with the Boggio formula ( e.g., Boggio [4] or Gazzola-Grunau-Sweers [20, Lemma 2.27]).
+We next combine this and the biharmonic Poisson kernel to derive explicit formulae for other Green functions.
+4.1 Noncritical dimensions
+Recall the fundamental solution of $\Delta^{2}$ in $\mathbb{R}^{n+1}$:
 $\Gamma(X-Y)=\Gamma(|X-Y|)=\frac{1}{2(n-1)(n-3)|\mathbb{S}^{n}|}|X-Y|^{3-n}.$
-Proof of Theorem 1.3. We denote by $G^{(i,j)}$ and $\bar{G}^{(i,j)}$ the Green functions in +$\mathbb{R}^{n+1}_{+}$ and $\mathbb{B}^{n+1}$, respectively.
-Fix $X\in\mathbb{R}_{+}^{n+1}$, then for all $Y\in\partial\mathbb{R}_{+}^{n+1}$, direct computations (cf. calculations in the proof of Theorem 1.1) show
+For brevity, we let $\xi^{\ast}=|\xi|^{-2}\xi$ for $0\neq\xi\in\overline{\mathbb{B}^{n+1}}$. Notice that
+$|\xi|^{2}|\xi^{\ast}-\eta|^{2}=|\xi-\eta|^{2}+(1-|\xi|^{2})(1-|\eta|^{2})=|\eta|^{2}|\eta^{\ast}-\xi|^{2}$
+can be extended to define for all $\xi,\eta\in\mathbb{B}^{n+1}$.
+Theorem 4.1.
+For $n\neq 3$, the representation formulae of Green functiona $\bar{G}^{(i,j)}$ to (4.1) are
+$\bar{G}^{(0,1)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\frac{1}{2}\Gamma(|\xi|\cdot|\xi^{\ast}-\eta|)\left[(n-3)\frac{|\xi-\eta|^{2}}{|\xi|^{2}\cdot|\xi^{\ast}-\eta|^{2}}-(n-1)\right];$
+$\bar{G}^{(2,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)-\frac{1}{2}\Gamma(|\xi|\cdot|\xi^{\ast}-\eta|)\left[(n-3)\frac{|\xi-\eta|^{2}}{|\xi|^{2}\cdot|\xi^{\ast}-\eta|^{2}}-(n-1)\right];$
+$\bar{G}^{(0,2)}(\xi,\eta)=$ $\Gamma(\xi-\eta)-\Gamma(|\xi|\cdot|\xi^{*}-\eta|);$
+$\bar{G}^{(1,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\Gamma(|\xi|\cdot|\xi^{*}-\eta|);$
+for all $\xi,\eta\in\mathbb{B}^{n+1},\xi\neq\eta$, where $\xi^{\ast}=|\xi|^{-2}\xi$ for $0\neq\xi\in\overline{\mathbb{B}^{4}}$. Moreover, there holds
+$0\leq\bar{G}^{(0,1)}(\xi,\eta)\leq\bar{G}^{(0,2)}(\xi,\eta)\leq\bar{G}^{(1,3)}(\xi,\eta)\leq\bar{G}^{(2,3)}(\xi,\eta),\quad\forall~{}\xi,\eta\in\overline{\mathbb{B}^{n+1}},\xi\neq\eta,$
+with equality if and only if either $\xi\in\mathbb{S}^{n}$ or $\eta\in\mathbb{S}^{n}$.
+Proof of Theorems 1.2 and 4.1. Fix $X\in\mathbb{R}_{+}^{n+1}$. Then for all $Y\in\partial\mathbb{R}_{+}^{n+1}$, direct computations (cf. calculations in the proof of Theorem 1.1) show
 $\begin{split}\mathscr{B}^{3}_{0}(\Gamma(X-Y))=&\frac{1}{2}P^{3}_{3}(x-y,t);\\ \mathscr{B}^{3}_{1}(\Gamma(X-Y))=&-\frac{t}{2(n-1)|\mathbb{S}^{n}|(|x-y|^{2}+t^{2})^{\frac{n-1}{2}}}=\frac{1}{2}P^{3}_{2}(x-y,t);\\ \mathscr{B}^{3}_{2}(\Gamma(X-Y))=&\frac{1}{|\mathbb{S}^{n}|}\frac{t^{2}}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}=-\frac{1}{2}P^{3}_{1}(x-y,t);\\ \mathscr{B}^{3}_{3}(\Gamma(X-Y))=&-\frac{(n+1)}{|\mathbb{S}^{n}|}\frac{t^{3}}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}=-\frac{1}{2}P^{3}_{0}(x-y,t).\end{split}$ (4.2)
-The above formulae demonstrate an intimate connection between the Poisson kernel and fundamental function.
-By definition of Green functions on both +$\mathbb{R}^{n+1}_{+}$ and $\mathbb{B}^{n+1}$, we invoke the conformal invariance of $P_{4}^{g}$ to show
-$\bar{G}^{(i,j)}(\xi,\eta)=G^{(i,j)}(X,Y)\left(\frac{2}{|\eta+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}$ (4.2)
-by setting $u=U\circ FU_{0}$ and $f=\bar{f}\circ FU_{0}^{(n+5)/(n-3)}$. For brevity, we let
+Here each $\mathscr{B}^{3}_{k},k\in\{1,2,3\}$ acts on $\Gamma$ with respect to the variable $Y$.
+By definition of Green functions on +$\mathbb{R}^{n+1}_{+}$ and $\mathbb{B}^{n+1}$, we invoke the conformal invariance of $P_{4}^{g}$ to show
+$\bar{G}^{(i,j)}(\xi,\eta)=G^{(i,j)}(X,Y)\left(\frac{2}{|\eta+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}.$ (4.3)
+For brevity, we let
 $H^{(i,j)}(X,Y)=G^{(i,j)}(X,Y)-\Gamma(X-Y)$
-and
-$\bar{H}^{(i,j)}(\xi,\eta)=\bar{G}^{(i,j)}(\xi,\eta)-\Gamma(\xi-\eta).$
 Then $H^{(i,j)}$ solves
-$\begin{cases}\displaystyle\Delta^{2}H^{(i,j)}=0\qquad&\mathrm{in}\qquad\mathbb{R}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(H^{(i,j)})=-\mathscr{B}^{3}_{i}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{j}(H^{(i,j)})=-\mathscr{B}^{3}_{j}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}^{n+1}.\end{cases}$
+$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}H^{(i,j)}=0\qquad&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(H^{(i,j)})=-\mathscr{B}^{3}_{i}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{j}(H^{(i,j)})=-\mathscr{B}^{3}_{j}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}_{+}^{n+1}.\end{cases}$
 For $(i,j)=(0,2)$, we find
@@
 $G^{(0,2)}(X,Y)=\Gamma(X-Y)-\Gamma(\bar{X}-Y).$
-Let $\xi^{\ast}=|\xi|^{-2}\xi$ whenever $\xi\neq 0$, then it is not hard to see that $F(\bar{X})=\xi^{\ast}$. Using
+For $(i,j)=(1,3)$, similarly we have
+$G^{(1,3)}(X,Y)=\Gamma(X-Y)+\Gamma(\bar{X}-Y).$
+For $(i,j)=(0,1),(2,3)$, the Boggio formula (see, Boggio [4] or Gazzola-Grunau-Sweers [20, Lemma 2.27]) gives
+$G^{(0,1)}(X,Y)=$ $\frac{1}{4|\mathbb{S}^{n}|}|X-Y|^{3-n}\int_{1}^{\frac{|\bar{X}-Y|}{|X-Y|}}(s^{2}-1)s^{-n}\mathrm{d}s$
+$=$ $\Gamma(X-Y)+\frac{1}{2}\Gamma(|\bar{X}-Y|)\left[(n-3)\frac{|X-Y|^{2}}{|\bar{X}-Y|^{2}}-(n-1)\right].$
+Set
+$f_{2}(y)=-\mathscr{B}^{3}_{2}(\Gamma(X-Y))=\frac{1}{2}P^{3}_{1}(x-y,t)$
+and
+$f_{3}(y)=-\mathscr{B}^{3}_{3}(\Gamma(X-Y))=\frac{1}{2}P^{3}_{0}(x-y,t).$
+Thanks to Theorem 1.1 and (4.2), a clever observation is
+$H^{(2,3)}(X,Y)=$ $\int_{\mathbb{R}^{n}}P^{3}_{2}(y-z,s)f_{2}(z)\mathrm{d}z+\int_{\mathbb{R}^{n}}P^{3}_{3}(y-z,s)f_{3}(z)\mathrm{d}z$
+$=$ $\frac{1}{2}\int_{\mathbb{R}^{n}}P^{3}_{2}(y-z,s)P^{3}_{1}(x-z,t)\mathrm{d}z+\frac{1}{2}\int_{\mathbb{R}^{n}}P^{3}_{3}(y-z,s)P^{3}_{0}(x-z,t)\mathrm{d}z$
+$=$ $\int_{\mathbb{R}^{n}}P^{3}_{0}(x-z,t)\mathscr{B}^{3}_{0}\big{(}\Gamma(y-z,s)\big{)}\mathrm{d}z+\int_{\mathbb{R}^{n}}P^{3}_{1}(x-z,t)\mathscr{B}^{3}_{1}\big{(}\Gamma(y-z,s)\big{)}\mathrm{d}z$
+$=$ $-H^{(0,1)}(X,Y).$
+Then it follows
+$G^{(2,3)}(X,Y)=$ $\Gamma(X-Y)-\frac{1}{2}\Gamma(|\bar{X}-Y|)\left[(n-3)\frac{|X-Y|^{2}}{|\bar{X}-Y|^{2}}-(n-1)\right].$
+Observe that
+$|\bar{X}-Y|^{2}=|X-Y|^{2}+4st,\qquad\forall~{}X,Y\in\overline{\mathbb{R}_{+}^{n+1}}.$
+Then we have
+$G^{(2,3)}(X,Y)\geq\Gamma(X-Y)+\Gamma(|\bar{X}-Y|)=$ $G^{(1,3)}(X,Y)\geq G^{(0,2)}(X,Y)\geq 0,$
+$[G^{(0,2)}-G^{(0,1)}](X,Y)=$ $\frac{n-3}{2}\Gamma(|\bar{X}-Y|)\left(1-\frac{|X-Y|^{2}}{|\bar{X}-Y|^{2}}\right)\geq 0.$
+This finish the proof of Theorem 1.2.
+Let $\xi^{\ast}=|\xi|^{-2}\xi$ whenever $\xi\neq 0$. Noitice that $F(\bar{X})=\xi^{\ast}$. Using
 $|X+\mathbf{e}_{n+1}||\xi+\mathbf{e}_{n+1}|=2,\quad|\bar{X}+\mathbf{e}_{n+1}||\xi^{\ast}+\mathbf{e}_{n+1}|=2,$
@@
 $2|\xi||\xi^{*}-\eta|=|\xi+\mathbf{e}_{n+1}||\eta+\mathbf{e}_{n+1}||\bar{X}-Y|.$ (4.4)
-Consequently, by (4.2), (3.8) and (4.3) we obtain
-$\bar{G}^{(0,2)}(\xi,\eta)=$ $\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\left(\frac{2}{|\eta+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}G^{(0,2)}(X,Y)$
-$=$ $\Gamma(\xi-\eta)-\Gamma(|\xi|\cdot|\xi^{*}-\eta|).$
-For $(i,j)=(1,3)$, similarly we have
-$G^{(1,3)}(X,Y)=\Gamma(X-Y)+\Gamma(\bar{X}-Y)$
-and
-$\bar{G}^{(1,3)}(\xi,\eta)=\Gamma(\xi-\eta)+\Gamma(|\xi|\cdot|\xi^{*}-\eta|).$
-For $(i,j)=(0,1),(2,3)$, the Boggio formula (see, Boggio [2] or Gazzola-Grunau-Sweers [15, Lemma 2.27]) gives
-$\bar{G}^{(0,1)}(\xi,\eta)=$ $\frac{1}{4|\mathbb{S}^{n}|}|\xi-\eta|^{3-n}\int_{1}^{\frac{|\xi||\xi^{\ast}-\eta|}{|\xi-\eta|}}(s^{2}-1)s^{-n}\mathrm{d}s$
-$=$ $\Gamma(\xi-\eta)+\frac{1}{2}\Gamma(|\xi|\cdot|\xi^{\ast}-\eta|)\left[(n-3)\frac{|\xi-\eta|^{2}}{|\xi|^{2}\cdot|\xi^{\ast}-\eta|^{2}}-(n-1)\right].$
-Set
-$f_{2}(y)=-\mathscr{B}^{3}_{2}(\Gamma(X-Y))=\frac{1}{2}P^{3}_{1}(x-y,t)$
-and
-$f_{3}(y)=-\mathscr{B}^{3}_{3}(\Gamma(X-Y))=\frac{1}{2}P^{3}_{0}(x-y,t).$
-We apply Theorem 1.1 together with (4.1) to solve
-$H^{(2,3)}(X,Y)=$ $\int_{\mathbb{R}^{n}}P^{3}_{2}(y-z,s)f_{2}(z)\mathrm{d}z+\int_{\mathbb{R}^{n}}P^{3}_{3}(y-z,s)f_{3}(z)\mathrm{d}z$
-$=$ $\frac{1}{2}\int_{\mathbb{R}^{n}}P^{3}_{2}(y-z,s)P^{3}_{1}(x-z,t)\mathrm{d}z+\frac{1}{2}\int_{\mathbb{R}^{n}}P^{3}_{3}(y-z,s)P^{3}_{0}(x-z,t)\mathrm{d}z.$
-A similar trick gives
-$H^{(0,1)}(X,Y)=-H^{(2,3)}(X,Y).$
-Hence, by (3.8) and (4.2) we have
-$\bar{H}^{(2,3)}(\xi,\eta)=$ $\left(\frac{2}{|\eta+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}H^{(2,3)}(X,Y)$
-$=$ $-\left(\frac{2}{|\eta+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}H^{(0,1)}(X,Y)=-\bar{H}^{(0,1)}(\xi,\eta),$
-from which we conclude that
-$\bar{G}^{(2,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)-\frac{1}{2}\Gamma(|\xi|\cdot|\xi^{\ast}-\eta|)\left[(n-3)\frac{|\xi-\eta|^{2}}{|\xi|^{2}\cdot|\xi^{\ast}-\eta|^{2}}-(n-1)\right].$
-Until now, we have already obtained biharmonic Green functions on $\mathbb{R}_{+}^{n+1}$ at the same time.
-Observe that
-$|\xi|^{2}|\xi^{\ast}-\eta|^{2}=|\xi-\eta|^{2}+(1-|\xi|^{2})(1-|\eta|^{2}),\qquad\forall~{}\xi,\eta\in\overline{\mathbb{B}^{n+1}},$
-then it follows that
+Therefore, using (4.3), (3.8) and (4.4) the explicit representation formulae $\bar{G}^{(i,j)}$ for $\mathbb{B}^{n+1}$ follow from those of $G^{(i,j)}$ for $\mathbb{R}_{+}^{n+1}$.
+Notice that
+$|\xi|^{2}|\xi^{\ast}-\eta|^{2}=|\xi-\eta|^{2}+(1-|\xi|^{2})(1-|\eta|^{2}),\qquad\forall~{}\xi,\eta\in\overline{\mathbb{B}^{n+1}}.$
+Then
 $\bar{G}^{(2,3)}(\xi,\eta)\geq$ $\Gamma(\xi-\eta)+\Gamma(|\xi|\cdot|\xi^{*}-\eta|)$
@@
 $[\bar{G}^{(0,2)}-\bar{G}^{(0,1)}](\xi,\eta)=\frac{n-3}{2}\Gamma(|\xi|\cdot|\xi^{\ast}-\eta|)\left(1-\frac{|\xi-\eta|^{2}}{|\xi|^{2}\cdot|\xi^{\ast}-\eta|^{2}}\right)\geq 0.$
-Consequently, we combine these facts together to conclude the second assertion. $\Box$
+Hence the second assertion follows. This completes the proof of Theorem 4.1. $\Box$
 4.2 Critical dimension
-The fundamental solution for $\Delta^{2}$ on $\mathbb{R}_{+}^{4}$ is given by
+The purpose is to derive explicit formulae of biharmonic Poisson kernels for $\mathbb{R}_{+}^{4}$ and $\mathbb{B}^{4}$.
+Recall that the fundamental solution for $\Delta^{2}$ on $\mathbb{R}_{+}^{4}$ is
 $\Gamma(X-Y)=-\frac{1}{4|\mathbb{S}^{3}|}\log|X-Y|.$
-Proof of Theorem 1.5. Fix $X\in\mathbb{R}^{4}_{+}$, then for all $Y\in\partial\mathbb{R}^{4}_{+}$, it is straightforward to show
-$\mathscr{B}^{3}_{0}(\Gamma(X-Y))=$ $\frac{1}{2}P^{3}_{3}(x-y,t);$
-$\mathscr{B}^{3}_{1}(\Gamma(X-Y))=$ $\frac{1}{2}P^{3}_{2}(x-y,t);$
-$\mathscr{B}^{3}_{2}(\Gamma(X-Y))=$ $-\frac{1}{2}P^{3}_{1}(x-y,t);$
-$\mathscr{B}^{3}_{3}(\Gamma(X-Y))=$ $-\frac{1}{2}P^{3}_{0}(x-y,t).$
-We next derive the Green function on $\mathbb{B}^{4}$ associated with the unique smooth solution of
+Theorem 4.2.
+The representation formula of Green function $\bar{G}^{(i,j)}$ to (4.1) is
+$\bar{G}^{(0,1)}(\xi,\eta)=$ $\Gamma(\xi-\eta)-\Gamma(|\xi|\cdot|\xi^{*}-\eta|)+\frac{1}{8|\mathbb{S}^{3}|}\left(\frac{|\xi-\eta|^{2}}{|\xi|^{2}|\xi^{*}-\eta|^{2}}-1\right);$
+$\bar{G}^{(2,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\Gamma(|\xi|\cdot|\xi^{*}-\eta|)-\frac{1}{8|\mathbb{S}^{3}|}\left(\frac{|\xi-\eta|^{2}}{|\xi|^{2}|\xi^{*}-\eta|^{2}}-1\right)+C;$
+$\bar{G}^{(0,2)}(\xi,\eta)=$ $\Gamma(\xi-\eta)-\Gamma(|\xi|\cdot|\xi^{*}-\eta|);$
+$\bar{G}^{(1,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\Gamma(|\xi|\cdot|\xi^{*}-\eta|)+C;$
+for $\xi,\eta\in\overline{\mathbb{B}^{4}},\xi\neq\eta$ and $C\in\mathbb{R}$.
+Proof of Theorems 1.3 and 4.2. By definition of biharmonic Green function, we have
+$H^{(i,j)}(X,Y)=G^{(i,j)}(X,Y)-\Gamma(X-Y)$
+satisfy
+$\begin{cases}\displaystyle\Delta^{2}H^{(i,j)}=0\qquad&\mathrm{~{}~{}in~{}~{}}\qquad~{}~{}\mathbb{R}^{4}_{+},\\ \displaystyle\mathscr{B}^{3}_{i}(H^{(i,j)})=-\mathscr{B}^{3}_{i}(\Gamma)\qquad&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{4}_{+},\\ \displaystyle\mathscr{B}^{3}_{j}(H^{(i,j)})=-\mathscr{B}^{3}_{j}(\Gamma)\qquad&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{4}_{+}.\end{cases}$
+For $(i,j)=(0,2)$, we find
+$H^{(0,2)}(X,Y)=-\Gamma(\bar{X}-Y)$
+whence,
+$G^{(0,2)}(X,Y)=\Gamma(X-Y)-\Gamma(\bar{X}-Y).$
+By (1.7) and symmetry we find
+$G^{(1,3)}(X,Y)=\Gamma(X-Y)+\Gamma(|\bar{X}-Y|)+C,\quad C\in\mathbb{R}.$
+For $(i,j)=(0,1),(2,3)$, again the Boggio formula gives
+$G^{(0,1)}(X,Y)=$ $\frac{1}{4|\mathbb{S}^{3}|}\int_{1}^{\frac{|\bar{X}-Y|}{|X-Y|}}(s^{2}-1)s^{-3}\mathrm{d}s$
+$=$ $\Gamma(X-Y)-\Gamma(|\bar{X}-Y|)+\frac{1}{8|\mathbb{S}^{3}|}\left(\frac{|X-Y|^{2}}{|\bar{X}-Y|^{2}}-1\right).$
+Notice that
+$H^{(2,3)}(X,Y)=$ $\frac{1}{2}\int_{\mathbb{R}^{3}}P^{3}_{2}(y-z,s)P^{3}_{1}(x-z,t)\mathrm{d}z+\frac{1}{2}\int_{\mathbb{R}^{3}}(P^{3}_{3}(y-z,s)+c)P^{3}_{0}(x-z,t)\mathrm{d}z$
+$=$ $\frac{1}{2}\int_{\mathbb{R}^{3}}P^{3}_{2}(y-z,s)P^{3}_{1}(x-z,t)\mathrm{d}z+\frac{1}{2}\int_{\mathbb{R}^{3}}P^{3}_{3}(y-z,s)P^{3}_{0}(x-z,t)\mathrm{d}z+\frac{C}{2}.$
+Then
+$H^{(2,3)}(X,Y)=-H^{(0,1)}(X,Y)+\frac{C}{2}.$
+Consequently, we have
+$G^{(2,3)}(X,Y)=$ $\Gamma(X-Y)+\Gamma(|\bar{X}-Y|)-\frac{1}{8|\mathbb{S}^{3}|}\left[\frac{|X-Y|^{2}}{|\bar{X}-Y|^{2}}-1\right]+C.$
+This finishes the proof of Theorem 1.3.
+Next, we derive the Green function on $\mathbb{B}^{4}$ associated with the unique smooth solution of
 $\begin{cases}\displaystyle\Delta^{2}U=f\qquad\quad&\mathrm{in}\qquad\mathbb{B}^{4},\\ \displaystyle\mathscr{B}^{3}_{i}U=0\qquad&\mathrm{on}\qquad\mathbb{S}^{3},\\ \displaystyle\mathscr{B}^{3}_{j}U=0\qquad&\mathrm{on}\qquad\mathbb{S}^{3},\end{cases}$
@@
 $\bar{G}^{(i,j)}(\xi,\eta)=G^{(i,j)}(F^{-1}(\xi),F^{-1}(\eta)).$
-As before, we let
-$H^{(i,j)}(X,Y)=G^{(i,j)}(X,Y)-\Gamma(X-Y)$
-satisfy
-$\begin{cases}\displaystyle\Delta^{2}H^{(i,j)}=0\qquad&\mathrm{in}\qquad\mathbb{R}^{4}_{+},\\ \displaystyle\mathscr{B}^{3}_{i}(H^{(i,j)})=-\mathscr{B}^{3}_{i}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}^{4}_{+},\\ \displaystyle\mathscr{B}^{3}_{j}(H^{(i,j)})=-\mathscr{B}^{3}_{j}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}^{4}_{+}.\end{cases}$
-For $(i,j)=(0,2)$, we find
-$H^{(0,2)}(X,Y)=-\Gamma(\bar{X}-Y)$
-whence,
-$G^{(0,2)}(X,Y)=\Gamma(X-Y)-\Gamma(\bar{X}-Y).$
-This together with (3.8), (4.3) gives
-$\bar{G}^{(0,2)}(\xi,\eta)=\Gamma(\xi-\eta)-\Gamma(|\xi||\xi^{*}-\eta|).$
-We introduce
-$\bar{G}^{(1,3)}(\xi,\eta)=\Gamma(\xi-\eta)+\Gamma(|\xi||\xi^{*}-\eta|)+C,\quad C\in\mathbb{R}.$
-For $(i,j)=(0,1),(2,3)$, again the Boggio formula gives
-$\bar{G}^{(0,1)}(\xi,\eta)=$ $\frac{1}{4|\mathbb{S}^{3}|}\int_{1}^{\frac{|\xi||\xi^{*}-\eta|}{|\xi-\eta|}}(s^{2}-1)s^{-3}\mathrm{d}s$
-$=$ $\Gamma(\xi-\eta)-\Gamma(|\xi||\xi^{*}-\eta|)+\frac{1}{8|\mathbb{S}^{3}|}\left(\frac{|\xi-\eta|^{2}}{|\xi|^{2}|\xi^{*}-\eta|^{2}}-1\right).$
-Notice that
-$H^{(2,3)}(X,Y)=$ $\frac{1}{2}\int_{\mathbb{R}^{3}}P^{3}_{2}(y-z,s)P^{3}_{1}(x-z,t)\mathrm{d}z+\frac{1}{2}\int_{\mathbb{R}^{3}}(P^{3}_{3}(y-z,s)+c)P^{3}_{0}(x-z,t)\mathrm{d}z$
-$=$ $\frac{1}{2}\int_{\mathbb{R}^{3}}P^{3}_{2}(y-z,s)P^{3}_{1}(x-z,t)\mathrm{d}z+\frac{1}{2}\int_{\mathbb{R}^{3}}P^{3}_{3}(y-z,s)P^{3}_{0}(x-z,t)\mathrm{d}z+\frac{C}{2},$
-then
-$H^{(2,3)}(X,Y)=-H^{(0,1)}(X,Y)+\frac{C}{2}.$
-Consequently, we have
-$\bar{G}^{(2,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\Gamma(|\xi||\xi^{*}-\eta|)-\frac{1}{8|\mathbb{S}^{3}|}\left[\frac{|\xi-\eta|^{2}}{|\xi|^{2}|\xi^{*}-\eta|^{2}}-1\right]+C.$
-$\Box$
-Eventually, combining the biharmonic Poisson kernel and Green function, we are in a position to present a Green formula for an nonhomogeneous biharmonic boundary value problem on $\mathbb{B}^{n+1}$.
-Corollary 4.1 (Green formula for a biharmonic boundary value problem).
+This completes the proof of Theorem 4.2. $\Box$
+Combining the biharmonic Poisson kernel and Green function, we are in a position to present a Green formula for an nonhomogeneous biharmonic BVP on $\mathbb{B}^{n+1}$.
+Corollary 4.1 (Biharmonic Poisson and Green integral).
 For $i,j\in\{0,1,2,3\}$ and $i+j\neq 3$, suppose $\bar{f}\in C^{\infty}(\overline{\mathbb{B}^{n+1}}),\bar{f}_{i},\bar{f}_{j}\in C^{\infty}(\mathbb{S}^{n})$. Then the unique solution of
@@
 $U(\xi)=[\bar{G}^{(i,j)}\ast\bar{f}+\bar{P}_{i}^{3}\ast\bar{f}_{i}+\bar{P}_{j}^{3}\ast\bar{f}_{j}](\xi).$
-As an application of the Green formula, we are able to obtain the following comparison principles for biharmonic boundary value problems on $\mathbb{B}^{n+1}$.
+As an application of the Green formula, we are able to obtain the following comparison principles for biharmonic BVPs on $\mathbb{B}^{n+1}$.
 Proposition 4.1.
@@
 Proof..
-This is a direct consequent of Corollary 4.1, Theorem 1.2 and Theorem 1.3. $\Box$
+This is a direct consequence of Corollary 4.1, Theorems 3.1 and 4.1. $\Box$
 5 Classification theorems
-As above, let $P_{k}(X)$ and $P_{k}(x)$ denote a polynomial with degree $k$ in the variables $X$ and $x$, where $k\in\mathbb{N}$.
+Our approach builds on the intrinsic third-order GJMS operator $P_{3}=\frac{1}{2}\mathscr{B}_{3}^{3}$ in our setting, and the basics of the geometric bubble and biharmonic Poisson integral.
+An important property of the geometric bubble $U_{x_{0},\varepsilon}$ is that for $n\geq 4$ and $i,j\in\{1,2,3\},i<j$, $U_{x_{0},\varepsilon}$ satisfies the BVP (1.5) with critical growth that $p_{i}=p_{i}^{\ast},p_{j}=p_{j}^{\ast}$.
+There are two ways to see this.
+One is through a straightforward computation.
+The other is in a geometric way of exhausting all possible geometric bubbles: A conformal transformation on $\overline{\mathbb{B}^{n+1}}$ can be expressed as
+$\zeta:=\psi_{a}(\xi)=\frac{\xi-a-|\xi|^{2}a+2(a\cdot\xi)a-|a|^{2}\xi}{1-2a\cdot\xi+|a|^{2}|\xi|^{2}},\quad\forall~{}\xi\in\mathbb{B}^{n+1}$
+for some $a\in\mathbb{B}^{n+1}$ (see e.g., [28, Chapter 1]), and there holds
+$\psi_{a}^{\ast}(|\mathrm{d}\zeta|^{2})=\left(\frac{1-|a|^{2}}{|\xi|^{2}|a|^{2}-2a\cdot\xi+1}\right)^{2}|\mathrm{d}\xi|^{2}.$
+Indeed, one can verify that
+$F^{\ast}(\psi_{a}^{\ast}(|\mathrm{d}\zeta|^{2}))=U_{x_{0},\varepsilon}^{\frac{4}{n-3}}(X)|\mathrm{d}X|^{2}$
+through the relation
+$a=F(x_{0},\varepsilon)\qquad\mathrm{by~{}~{}regarding~{}~{}}(x_{0},\varepsilon)\in\mathbb{R}_{+}^{n+1}.$
+Equivalently, the BVP (1.5) with critical growth can interpreted as the pullback of the metric $(\mathbb{B}^{n+1},\psi_{a}^{\ast}(|\mathrm{d}\zeta|^{2}))$ via the map $F$ has constant curvatures $(T_{i}^{3})_{\mathbb{S}^{n}},i\in\{1,2,3\}$.
+As before, let $P_{k}(x)$ denote a polynomial with degree $\leq k$ for $k\in\mathbb{N}$.
 5.1 Proof of Theorem 1.4
@@
 $\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}_{3}^{3}(u)=\mathbb{T}^{3}_{3}u^{p_{3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\end{cases}$
-for $0<p_{3}\leq p_{3}^{\ast}=\frac{n+3}{n-3}$. Assume
+for $0<p_{3}\leq p_{3}^{\ast}=\frac{n+3}{n-3}$. Assume that
 $u(x,0)=O(|x|^{-c}),\qquad\partial_{t}u(x,0)=O(|x|^{-d})\qquad\mathrm{~{}~{}as~{}~{}}|x|\to\infty$
 for some constants $c>3/p_{3}$ and $d>1$. Then we have
-  • (i)
+  1. (i)
     If $p_{3}=p_{3}^{\ast}$, then there exist $\varepsilon>0$ and $x_{0}\in\mathbb{R}^{n}$ such that
     $u(x,0)=\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}}.$
-  • (ii)
+  2. (ii)
     If $0<p_{3}<p_{3}^{\ast}$, then $u(x,0)=0$.
@@
 By Theorem 1.1, $w:=u-v$ solves
-$\begin{cases}\displaystyle\Delta^{2}w=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\partial_{t}w(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta w(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
+$\begin{cases}\displaystyle~{}\Delta^{2}w=0&\mathrm{~{}~{}in~{}~{}}\qquad~{}~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle~{}~{}~{}\partial_{t}w=0&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta w=0\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
 Moreover, since $u$ is nonnegative, by (5.1) we have
@@
 $u(x,0)=\frac{n+1}{4|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u^{p_{3}}(y,0)}{|x-y|^{n-3}}\mathrm{d}y.$
-The classification theorem (see Chen-Li-Ou [9] and Y. Y. Li [24]) for the integral equation above gives
+The classification theorem (e.g., Y. Y. Li [30] and Chen-Li-Ou [12]) for the integral equation above gives
 $u(x,0)=\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}},\quad\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}\qquad\qquad\mathrm{~{}~{}for~{}~{}}p_{3}=p_{3}^{\ast}$
@@
 Now $u$ solves
-$\begin{cases}\displaystyle\Delta^{2}u=0&\mathrm{in}\quad\mathbb{R}_{+}^{n+1},\\ \displaystyle\partial_{t}u(x,0)=-\frac{n-3}{2}\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{(n-3)p_{1}}{2}}&\mathrm{on}\quad\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}u=\frac{(n^{2}-1)(n-3)}{4}\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{n+3}{2}}\quad&\mathrm{on}\quad\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (5.3)
+$\begin{cases}\displaystyle\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\quad~{}~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle~{}~{}\partial_{t}u=-\frac{n-3}{2}\Big{(}\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\Big{)}^{\frac{(n-3)p_{1}}{2}}&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}u=\frac{(n^{2}-1)(n-3)}{4}\Big{(}\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\Big{)}^{\frac{n+3}{2}}\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (5.3)
 By (5.2) we know
-$u(x,t)=M_{1}(x,t)+c_{2}^{\prime}t^{2}.$ (5.4)
+$u(x,t)=M_{1}(x,t)+c_{2}^{\prime}t^{2},$ (5.4)
 where $c_{2}^{\prime}$ is a nonnegative constant and
-$M_{1}(x,t)=$ $\frac{n+1}{4|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{1}{(t^{2}+|x-y|^{2})^{\frac{n-3}{2}}}\left(\frac{2\varepsilon}{\varepsilon^{2}+|y-x_{0}|^{2}}\right)^{\frac{n+3}{2}}\mathrm{d}y$
-$-\frac{n-3}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\left(\frac{2\varepsilon}{\varepsilon^{2}+|y-x_{0}|^{2}}\right)^{\frac{(n-3)p_{1}}{2}}\mathrm{d}y.$
+$M_{1}(x,t)=$ $\frac{n+1}{4|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{1}{(t^{2}+|x-y|^{2})^{\frac{n-3}{2}}}\Big{(}\frac{2\varepsilon}{\varepsilon^{2}+|y-x_{0}|^{2}}\Big{)}^{\frac{n+3}{2}}\mathrm{d}y$
+$-\frac{n-3}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\Big{(}\frac{2\varepsilon}{\varepsilon^{2}+|y-x_{0}|^{2}}\Big{)}^{\frac{(n-3)p_{1}}{2}}\mathrm{d}y.$
 Especially, assume $p_{1}=p_{1}^{\ast}=\frac{n-1}{n-3}$. On one hand, by Lemmas 2.1 and 2.2 we have
@@
 On the other hand, we consider
-$w(x,t)=u(x,t)-\left(\frac{2\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}},$
+$w(x,t)=u(x,t)-U_{x_{0},\varepsilon}(x,t),$
 then $w$ satisfies
-$\begin{cases}\displaystyle\Delta^{2}w=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\partial_{t}w(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta w(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\end{cases}$ (5.6)
+$\begin{cases}\displaystyle~{}\Delta^{2}w=0\qquad\qquad\qquad&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle~{}~{}~{}\partial_{t}w=0&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta w=0&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\end{cases}$ (5.6)
 and for $|X|,|x|\gg 1$,
@@
 Hence, it follows from Lemma 2.3 that the solution of (5.3) has the form of
-$u(x,t)=\left(\frac{2\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}}+c_{2}t^{2}.$ (5.7)
-Eventually, we combine (5.4), (5.5) and (5.7) to conclude that $c_{2}^{\prime}=c_{2}$.
+$u(x,t)=U_{x_{0},\varepsilon}(x,t)+c_{2}t^{2}.$ (5.7)
+Therefore, we combine (5.4), (5.5) and (5.7) to conclude that $M_{1}=U_{x_{0},\varepsilon}$ and $c_{2}^{\prime}=c_{2}$.
 (2) If $p_{3}<p_{3}^{\ast}$, then again by Lemma 5.1 we obtain $u(x,0)=0$.
-In this case, $u$ also solves the same homogeneous biharmonic boundary value problem (5.6), thanks to Lemma 2.3, we conclude that $u(x,t)=c_{2}t^{2}$ for every nonnegative constant $c_{2}$.
+In this case, $u$ also solves the same homogeneous biharmonic BVP (5.6), thanks to Lemma 2.3, we conclude that $u(x,t)=c_{2}t^{2}$ for every nonnegative constant $c_{2}$.
 $\Box$
 5.2 Proof of Theorem 1.5
+From the viewpoint of the Lopatinski-Shapiro or complement condition, the conformal boundary operator pair $(\mathscr{B}_{1}^{3},\mathscr{B}_{2}^{3})$ is exceptional.
+Nevertheless, some different technique needs to be developed for treating this type of $T$-curvature equations.
 Proof of Theorem 1.5. For $cp_{2}>2$, $f_{0}(x)=u(x,0)=O(|x|^{-c})$ and $f_{2}(x)=\mathbb{T}^{3}_{2}u^{p_{2}}=O(|x|^{-p_{2}c})$ as $|x|\to\infty$. Let
-$v(x,t)=\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}f_{0}(y)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y-\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{tf_{2}(y)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y.$
+$v(x,t)=\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}f_{0}(y)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y-\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{tf_{2}(y)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y$
 with
-$|v(x,t)|\lesssim\begin{cases}|X|^{2-\min\{p_{2}c,n\}}&\mathrm{if~{}~{}}p_{2}c\neq n\\ |X|^{2-n}\log|X|&\mathrm{if~{}~{}}p_{2}c=n\end{cases}+\begin{cases}|X|^{-\min\{c,n\}}&\mathrm{if~{}~{}}p_{2}c\neq n\\ |X|^{-n}\log|X|&\mathrm{if~{}~{}}p_{2}c=n\end{cases}\to 0$ (5.8)
+$|v(x,t)|\lesssim\begin{cases}|X|^{-\min\{c,n\}}&\mathrm{if~{}~{}}c\neq n\\ |X|^{-n}\log|X|&\mathrm{if~{}~{}}c=n\end{cases}+\begin{cases}|X|^{2-\min\{p_{2}c,n\}}&\mathrm{if~{}~{}}p_{2}c\neq n\\ |X|^{2-n}\log|X|&\mathrm{if~{}~{}}p_{2}c=n\end{cases}\to 0$ (5.8)
 as $|X|\to\infty$ by virtue of Lemmas 2.1 and 2.2.
-By Theorem 1.1, $w:=u-v$ satisfies
-$\begin{cases}\displaystyle\Delta^{2}w=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle w(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\Delta w(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
-Moreover, by (5.8) we have
+By the assumption and Theorem 1.1, $w:=u-v$ satisfies
+$\begin{cases}\displaystyle\Delta^{2}w=0\qquad\qquad\qquad&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\quad~{}w=0&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle~{}~{}\Delta w=0&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
+Moreover, using the nonnegativity of $u$ and (5.8) we have
 $w(x,t)\geq-C|X|^{-\delta}\qquad\mathrm{as}\qquad|X|\to\infty,$
-where $\delta=\min\left\{\min\{p_{2}c,n\}-2,\min\{c,n\}\right\}/2>0$. By Lemma 2.4 we have
+where $\delta=\min\left\{\min\{p_{2}c,n\}-2,\min\{c,n\}\right\}/2>0$. Then we apply Lemma 2.4 to obtain
 $u(x,t)=$ $\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}u(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y-\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{tu^{p_{2}}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y$
 $+tP_{2}(x)+c_{3}t^{3},$ (5.9)
-where $c_{3}$ is a nonnegative constant.
+where $c_{3}$ is a nonnegative constant and $\liminf_{|x|\to\infty}P_{2}(x)\geq 0.$
 From Lemma 3.1 that
-$\lim_{t\to 0}\partial_{t}\int_{\mathbb{R}^{n}}\frac{t^{3}}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}u(y,0)\mathrm{d}y=0,$
-and the first boundary condition we obtain
+$\lim_{t\to 0}\partial_{t}\int_{\mathbb{R}^{n}}\frac{t^{3}u(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y=0,$
+together with the first boundary condition we obtain
 $\frac{n-3}{2}u^{p_{1}}(x,0)=-\partial_{t}u(x,0)=\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u^{p_{2}}(y,0)}{|x-y|^{n-1}}\mathrm{d}y-P_{2}(x).$ (5.10)
-Since $u(x,0)=o(1)$ as $|x|\to\infty$, by Lemma 2.1 we have
+Using the assumption that $u(x,0)=o(1)$ as $|x|\to\infty$ and Lemma 2.1 we have
 $P_{2}(x)=o(1)\qquad\mathrm{as}\qquad|x|\to+\infty,$
@@
 $k(x)=\frac{1}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{k^{\frac{p_{2}}{p_{1}}}(y)}{|x-y|^{n-1}}\mathrm{d}y.$
-(1) If $\frac{p_{2}}{p_{1}}=\frac{n+1}{n-1}$, then we apply the classification theorem (see [9] and [24]) to know
+(1) If $\frac{p_{2}}{p_{1}}=\frac{n+1}{n-1}$, then the classification theorem (e.g., [12] and [30]) yields
 $k(x)=\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-1}{2}},\qquad\varepsilon\in\mathbb{R}_{+},~{}x_{0}\in\mathbb{R}^{n}.$
@@
 $\mathscr{B}^{3}_{2}u=\frac{(n-1)(n-3)}{2}\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{n+1}{2}}.$
-By (5.2) we have
-$u(x,t)=M_{2}(x,t)+c_{3}t^{3}$
-and
+Going back to (5.2) we conclude that
+$u(x,t)=M_{2}(x,t)+c_{3}t^{3},$
+where
 $M_{2}(x,t)=$ $\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\left(\frac{2\varepsilon}{\varepsilon^{2}+|y-x_{0}|^{2}}\right)^{\frac{n-1}{2p_{1}}}\mathrm{d}y$
@@
 Especially, when $p_{1}=p_{1}^{\ast}=\frac{n-1}{n-3}$, we consider
-$w(x,t)=u(x,t)-\left(\frac{2\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}},$
-which satisfies
-$\begin{cases}\displaystyle\Delta^{2}w=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle w(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\Delta w(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\end{cases}$ (5.11)
-with
+$w(x,t)=u(x,t)-U_{x_{0},\varepsilon}(x,t).$
+Then $w$ solves
+$\begin{cases}\displaystyle\Delta^{2}w=0\qquad\qquad\qquad&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\quad~{}w=0&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle~{}~{}\Delta w=0&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\end{cases}$ (5.11)
+with some $C\in\mathbb{R}_{+}$ such that
 $w(X)\geq-C|X|^{3-n}\qquad\mathrm{as~{}~{}}|X|\to\infty.$
@@
 Hence, by Lemma 2.4 we obtain
-$u(x,t)=\left(\frac{2\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}}+c_{3}t^{3}.$
+$u(x,t)=U_{x_{0},\varepsilon}(x,t)+c_{3}t^{3}.$
 (2) If $\frac{p_{2}}{p_{1}}<\frac{n-1}{n-3}$, then the uniqueness theorem in Y.
 Y.
-Li [24] yields $u(x,0)=0$ in $\mathbb{R}^{n}$, thus $u(x,t)$ satisfies the homogeneous boundary value problem (5.11).
+Li [30] yields $u(x,0)=0$ in $\mathbb{R}^{n}$, so $u$ satisfies problem (5.11).
 This together with Lemma 2.4 yields $u(x,t)=c_{3}t^{3}$ in $\mathbb{R}_{+}^{n+1}$ for every nonnegative constant $c_{3}$.
@@
 5.3 Proof of Theorem 1.6
-Proof of Theorem 1.8. We now consider
-$\begin{cases}\displaystyle\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}_{2}^{3}(u)=\frac{(n-1)(n-3)}{2}u^{p_{2}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}(u)=\frac{(n^{2}-1)(n-3)}{4}u^{p_{3}}&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (5.12)
+Proof of Theorem 1.6. Recall that $u$ solves
+$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\quad~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}_{2}^{3}(u)=\mathbb{T}^{3}_{2}u^{p_{2}}&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}(u)=\mathbb{T}^{3}_{3}u^{p_{3}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (5.12)
 As $cp_{2}>2$, a similar argument as in Theorem 1.5 yields
@@
 $\liminf_{|x|\to\infty}P_{2}(x)\geq 0.$ (5.14)
-Moreover,
+Moreover, by (5.3) and Lemma 3.1 we have
 $-\partial_{t}u(x,0)=\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u^{p_{2}}(y,0)}{|x-y|^{n-1}}\mathrm{d}y-P_{2}(x).$
-We first claim that
-$P_{2}(x)\equiv c\geq 0\qquad\mathrm{and}\qquad c_{3}=0.$
-It is known from (5.14) that $\overline{\Delta}P_{2}(x)$ is a nonnegative constant. With $\tilde{c}_{3}:=c_{3}+\frac{1}{2}\overline{\Delta}P_{2}(x)$ we define
+We claim that there is some nonnegative constant $c$ such that
+$P_{2}(x)\equiv c\qquad\mathrm{~{}~{}and~{}~{}}\qquad c_{3}=0.$
+To this end, (5.14) yields that $\overline{\Delta}P_{2}(x)$ must be a nonnegative constant. With $\tilde{c}_{3}:=c_{3}+\frac{1}{2}\overline{\Delta}P_{2}(x)$ we introduce
 $\tilde{u}(x,t)=u(x,t)-P_{2}(x)t-c_{3}t^{3}+\tilde{c}_{3}t^{3}$
@@
 Clearly, $\tilde{u}$ is biharmonic in $\mathbb{R}_{+}^{n+1}$ and $\tilde{u}(x,0)=u(x,0)$.
-Recall that $\mathscr{B}^{3}_{2}=\partial_{t}^{2}-\overline{\Delta}$ and $\mathscr{B}^{3}_{3}=\partial_{t}\Delta+2\overline{\Delta}\partial_{t}$, then
+Recall that $\mathscr{B}^{3}_{2}=\partial_{t}^{2}-\overline{\Delta}$ and $\mathscr{B}^{3}_{3}=\partial_{t}\Delta+2\overline{\Delta}\partial_{t}$, then $\tilde{u}$ and $u$ satisfy the same boundary conditions, more explicitly,
 $\mathscr{B}^{3}_{3}(\tilde{u})=\mathscr{B}^{3}_{3}(u)+6(\tilde{c}_{3}-c_{3})-3\overline{\Delta}P_{2}(x)=\mathscr{B}^{3}_{3}(u)$
 and $\mathscr{B}^{3}_{2}(\tilde{u})=\mathscr{B}^{3}_{2}(u)$ on $\partial\mathbb{R}_{+}^{n+1}$.
-This means that $\tilde{u}$ also satisfies the biharmonic boundary value problem (5.12), though it might change sign.
-On the other hand, notice that $\tilde{u}(x,t)\geq-C|X|^{3}$ as $|X|\to\infty$. By Lemma 2.1 as $|x|\to\infty$ we have
-$-\partial_{t}\tilde{u}(x,0)=\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{\tilde{u}^{p_{2}}(y,0)}{|x-y|^{n-1}}\mathrm{d}y=$$\begin{cases}O(|x|^{1-\min\{p_{2}c,n\}})&\qquad\mathrm{if~{}~{}}p_{2}c\neq n\\ O(|x|^{1-n}\log|x|)&\qquad\mathrm{if~{}~{}}p_{2}c=n\end{cases}$
-$=$ $O(|x|^{-d}),$
-where $d=\min\{cp_{2}-1,(n-1)/2\}>1$. For $c>\max\{2/p_{2},3/p_{3}\}$, we apply Remark 2.1 to obtain
-$\tilde{u}(x,t)=\frac{2}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}\partial_{t}\tilde{u}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\mathrm{d}y+\frac{n+1}{4|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u^{p_{3}}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n-3}{2}}}\mathrm{d}y+t^{2}P_{1}(x).$ (5.16)
-By (5.3) and (5.16), we can apply Lemmas 2.1 and 2.2 to conclude that
+This means that $\tilde{u}$ also satisfies problem (5.12), though it might change sign in $\mathbb{R}_{+}^{n+1}$.
+On the other hand, notice that there is some $C\in\mathbb{R}_{+}$ such that $\tilde{u}(x,t)\geq-C|X|^{3}$ as $|X|\to\infty$. By Lemma 2.1 we have
+$-\partial_{t}\tilde{u}(x,0)=$ $\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{\tilde{u}^{p_{2}}(y,0)}{|x-y|^{n-1}}\mathrm{d}y$
+$=$ $\begin{cases}O(|x|^{1-\min\{p_{2}c,n\}})&\qquad\mathrm{if~{}~{}}p_{2}c\neq n\\ O(|x|^{1-n}\log|x|)&\qquad\mathrm{if~{}~{}}p_{2}c=n\end{cases}$
+$=$ $O(|x|^{-d})\qquad\qquad\mathrm{~{}~{}as~{}~{}}|x|\to\infty,$
+where $d=\min\{cp_{2}-1,(n-1)/2\}>1$. As $c>\max\{2/p_{2},3/p_{3}\}$ by assumption, we apply Remark 2.1 to obtain
+$\tilde{u}(x,t)$
+$=$ $\frac{2}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}\partial_{t}\tilde{u}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\mathrm{d}y+\frac{n+1}{4|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u^{p_{3}}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n-3}{2}}}\mathrm{d}y+t^{2}P_{1}(x).$ (5.16)
+Combining (5.3) and (5.3), we can apply Lemmas 2.1 and 2.2 to conclude that
 $\tilde{c}_{3}t^{3}-t^{2}P_{1}(x)=o(1)\qquad\mathrm{~{}~{}as~{}~{}}|X|\to\infty.$
-This implies $\tilde{c}_{3}=0$ and $P_{1}(x)\equiv 0$. Then, (5.16) becomes
+This implies $\tilde{c}_{3}=0$ and $P_{1}(x)\equiv 0$. Thus, (5.3) becomes
 $\tilde{u}(x,t)=\frac{2}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}\partial_{t}\tilde{u}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\mathrm{d}y+\frac{n+1}{4|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u^{p_{3}}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n-3}{2}}}\mathrm{d}y.$ (5.17)
 Keep in mind that
-$0=\tilde{c}_{3}=c_{3}+\frac{\overline{\Delta}P_{2}(x)}{2},$
-so, $c_{3}=0$ and $\overline{\Delta}P_{2}(x)=0$. This together with (5.14) implies $P_{2}(x)\equiv c_{1}\geq 0$. Thus, we prove the above claim.
-Going back to (5.3) and (5.3) we now have
+$0=\tilde{c}_{3}=c_{3}+\frac{1}{2}\overline{\Delta}P_{2}(x),$
+so, $c_{3}=0$ and $\overline{\Delta}P_{2}(x)=0$. This together with (5.14) implies $P_{2}(x)\equiv c_{1}\geq 0$. Hence, we prove the above claim.
+Going back to (5.3) and (5.3) we have
 $u(x,t)=$ $\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}u(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y-\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{tu^{p_{2}}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y$
 $+c_{1}t$ (5.18)
-and
+and there is some $C\in\mathbb{R}_{+}$ such that
 $\tilde{u}(x,t)=u(x,t)-c_{1}t\geq-C|X|\qquad\mathrm{as~{}~{}}|X|\to\infty.$ (5.19)
-Furthermore, letting $t\to 0$ in (5.17) we arrive at
+Furthermore, letting $t\to 0$ in (5.17), by Lemma 2.1 we arrive at
 $u(x,0)=\tilde{u}(x,0)=\frac{n+1}{4|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u^{p_{3}}(y,0)}{|x-y|^{n-3}}\mathrm{d}y.$
-(1) If $p_{3}=\frac{n+3}{n-3}$ and $p_{2}>\frac{2}{n-3}$, then the classification theorem in [9] and [24] gives
+(1) If $p_{3}=\frac{n+3}{n-3}$ and $p_{2}>\frac{2}{n-3}$, then the classification theorem (e.g., [30] and [12]) gives
 $u(x,0)=\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}},\quad x_{0}\in\mathbb{R}^{n},\varepsilon\in\mathbb{R}_{+}.$
@@
 $-\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\left(\frac{2\varepsilon}{\varepsilon^{2}+|y-x_{0}|^{2}}\right)^{\frac{(n-3)p_{2}}{2}}\mathrm{d}y.$
-In particular, when $p_{2}=p_{2}^{\ast}$, as before, we notice that $(\tilde{u}-U_{\varepsilon,x_{0}})(X)\geq-C|X|$ as $|X|\to\infty$, and apply Remark 2.2 to conclude that $M_{3}=U_{\varepsilon,x_{0}}$ and then
-$u(x,t)=U_{\varepsilon,x_{0}}(x,t)+c_{1}t.$
-(2) If $p_{3}<\frac{n+3}{n-3}$, then the uniqueness theorem in [24] implies $u(x,0)=0$.
-This together with (5.3) implies $u(x,t)=c_{1}t$ in $\mathbb{R}_{+}^{n+1}$ for every nonnegative constant $c_{1}$.
+In particular, when $p_{2}=p_{2}^{\ast}$, notice that there is some $C\in\mathbb{R}_{+}$ such that $(\tilde{u}-U_{x_{0},\varepsilon})(X)\geq-C|X|$ as $|X|\to\infty$.
+Then, we apply Remark 2.2 to conclude that $M_{3}=U_{x_{0},\varepsilon}$ and then
+$u(x,t)=U_{x_{0},\varepsilon}(x,t)+c_{1}t.$
+(2) If $p_{3}<\frac{n+3}{n-3}$, then the uniqueness theorem (e.g., [30]) yields $u(x,0)=0$.
+This together with (5.3) follows that $u(x,t)=c_{1}t$ in $\mathbb{R}_{+}^{n+1}$ for every nonnegative constant $c_{1}$.
 $\Box$
 6 Applications to biharmonic boundary value problems on balls
-Although we focus on the conformal geometry nature for biharmonic boundary value problems, we would like to mention that from a physical point of view, the boundary conditions $(\mathscr{B}_{0}^{3},\mathscr{B}_{1}^{3})$ and $(\mathscr{B}_{0}^{3},\mathscr{B}_{2}^{3})$ have a close connection with the clamped plate and hinged plate model (cf.
+Although we focus on the conformal geometry nature for biharmonic BVPs, we would like to mention that from a physical viewpoint, the boundary conditions $(\mathscr{B}_{0}^{3},\mathscr{B}_{1}^{3})$ and $(\mathscr{B}_{0}^{3},\mathscr{B}_{2}^{3})$ have a close connection with the clamped plate and hinged plate model (cf.
 [20]), respectively.
-However, the study of other boundary conditions is far less developed.
-6.1 One boundary point singularity
-Proof of Theorem 1.9. It suffices to give a proof for $i=1$ and $j=3$, since the others are the same in the spirit due to Theorems 1.7 and 1.8.
+However, the study of biharmonic equation coupled with other boundary conditions is far less developed.
+6.1 Solutions with single boundary point singularity
+Proof of Theorem 1.7. It suffices to give the proof for $i=1$ and $j=3$, since the others are the same in the spirit thanks to Theorems 1.5 and 1.6.
 Let $U$ be a nonnegative solution of (1.10). As before, we introduce
-$u(X)=\left(\frac{2}{|X+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}U\circ F.$
+$u(X)=\left(\frac{2}{|X+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}(U\circ F)(X).$
 Then $u$ solves
-$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(u)=\mathbb{T}^{3}_{1}u^{\frac{n-1}{n-3}}&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{j}(u)=\mathbb{T}^{3}_{3}u^{\frac{n+3}{n-3}}&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}_{+}^{n+1}.\end{cases}$
-By definition of the spherical inversion $F$, we know that $|X+\mathbf{e}_{n+1}||\xi+\mathbf{e}_{n+1}|=2$ and $F(\partial\mathbb{R}_{+}^{n+1})=\partial\mathbb{B}^{n+1}$.
+$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{1}(u)=\mathbb{T}^{3}_{1}u^{\frac{n-1}{n-3}}&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{3}(u)=\mathbb{T}^{3}_{3}u^{\frac{n+3}{n-3}}&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}_{+}^{n+1}.\end{cases}$
+By definition of the conformal map $F$, we know that $|X+\mathbf{e}_{n+1}||\xi+\mathbf{e}_{n+1}|=2$ and $F(\partial\mathbb{R}_{+}^{n+1})=\partial\mathbb{B}^{n+1}$.
 By the assumption (1.11) for $U$, we obtain
@@
 6.2 Sharp geometric inequalities
-On a compact manifold with boundary, the maximizing problem of the isometric ratio over a scalar-flat conformal class was studied in [22, 21].
-With comparison principles in Proposition 4.1, we are motivated to study of the maximizers of the isometric ratio and another sharp geometric inequality over some $Q$-curvature flat conformal class on $(\overline{\mathbb{B}^{n+1}},|\mathrm{d}\xi|^{2})$.
+On a smooth compact manifold with boundary, the maximizing problem of the isometric ratio over a scalar-flat conformal class was studied in [27, 26].
+With comparison principles in Proposition 4.1, we are motivated to study of the maximizers of the isometric ratio and another sharp geometric inequality over some $Q$-curvature flat conformal class of the flat metric $[|\mathrm{d}\xi|^{2}]$ on $\overline{\mathbb{B}^{n+1}}$.
 More general, for $n\geq 4$ we define
-$\mathcal{M}_{1}=$$\{g\in[|\mathrm{d}\xi|^{2}]~{}|~{}Q_{g}\leq 0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}(T_{2}^{3})_{g}\geq 0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\};$
-$\mathcal{M}_{2}=$$\{g\in[|\mathrm{d}\xi|^{2}]~{}|~{}Q_{g}\leq 0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}h_{g}\geq 0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\}$
+$\mathcal{M}_{1}=$$\{g\in[|\mathrm{d}\xi|^{2}];Q_{g}\leq 0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}(T_{2}^{3})_{g}\geq 0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\};$
+$\mathcal{M}_{2}=$$\{g\in[|\mathrm{d}\xi|^{2}];Q_{g}\leq 0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}h_{g}\geq 0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\}$
 and
-$\mathring{\mathcal{M}}:=$$\{g\in[|\mathrm{d}\xi|^{2}]~{}|~{}Q_{g}=0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}(T_{2}^{3})_{g}=0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\}$
-$=$ $\{g\in[|\mathrm{d}\xi|^{2}]~{}|~{}Q_{g}=0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}h_{g}=0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\}$
+$\mathring{\mathcal{M}}:=$$\{g\in[|\mathrm{d}\xi|^{2}];Q_{g}=0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}(T_{2}^{3})_{g}=0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\}$
+$=$ $\{g\in[|\mathrm{d}\xi|^{2}];Q_{g}=0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}h_{g}=0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\}$
 by virtue of Proposition 3.1.
@@
 Let $g=U^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2}$ be a smooth metric in $\overline{\mathbb{B}^{n+1}}$. Suppose $g\in\mathcal{M}_{1}\cup\mathcal{M}_{2}$. Then we have
-  • (1)
+  1. (1)
     there exists a conformal metric $\tilde{g}=\tilde{U}^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2}\in\mathring{\mathcal{M}}$ such that $\tilde{U}\geq U$ in $\mathbb{B}^{n+1}$ with $\tilde{U}=U$ on $\mathbb{S}^{n}$;
-  • (2)
+  2. (2)
     there exists a conformal metric $\tilde{g}=\tilde{U}^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2}\in\mathring{\mathcal{M}}$ such that $\tilde{U}\geq U$ in $\mathbb{B}^{n+1}$ with $\mathscr{B}^{3}_{3}(U)=\mathscr{B}^{3}_{3}(\tilde{U})$ on $\mathbb{S}^{n}$.
@@
 Theorem 6.1.
-Let $n\geq 4$. Then for all $g\in\mathcal{M}_{1}\cup\mathcal{M}_{2}$, there hold
+Let $n\geq 4$. Then for all $g\in\mathcal{M}_{1}\cup\mathcal{M}_{2}$,
 $|\mathbb{B}^{n+1}|_{g}\leq d_{n}|\mathbb{S}^{n}|_{g}^{\frac{n+1}{n}}\quad\mathrm{~{}~{}and~{}~{}}\quad|\mathbb{B}^{n+1}|^{\frac{n-3}{2(n+1)}}_{g}\leq e_{n}\|(T_{3}^{3})_{g}\|_{L^{\frac{2n}{n+3}}(\mathbb{S}^{n},g)},$
@@
 $d_{n}:=\sup_{\begin{subarray}{c}0<u\in C^{\infty}(\mathbb{R}^{n})\\ u=O(|x|^{3-n})\mathrm{~{}as~{}}|x|\to\infty\end{subarray}}\frac{\|E_{-2,3}(u)\|_{L^{\frac{2(n+1)}{n-3}}(\mathbb{R}^{n+1}_{+})}}{\|u\|_{L^{\frac{2n}{n-3}}(\mathbb{R}^{n})}}.$ (6.4)
-Gluck [17, Theorem 1.1] states that the optimal constant $d_{n}$ for the extremal problem (6.4) is attained by
+Gluck [22, Theorem 1.1] states that the optimal constant $d_{n}$ for the extremal problem (6.4) is attained by
 $c(|x-x_{0}|^{2}+d^{2})^{\frac{3-n}{2}},\qquad c,d\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n},$
-equivalently, modulo of a positive constant, the extremal metric on $\mathbb{S}^{n}$ is a conformal metric $u_{a}^{4/(n-3)}g_{\mathbb{S}^{n}}$ induced by some conformal transformation on $\mathbb{S}^{n}$, which can be expressed as
+equivalently, modulo of a positive constant, the extremal metric is a conformal metric $u_{a}^{4/(n-3)}g_{\mathbb{S}^{n}}$ induced by some conformal transformation on $\mathbb{S}^{n}$, which can be expressed as
 $u_{a}(\xi)=\Big{(}\frac{1-|a|^{2}}{|a|^{2}|-2a\cdot\xi+1}\Big{)}^{\frac{n-3}{2}}\qquad\mathrm{on~{}~{}}\quad\mathbb{S}^{n}$ (6.5)
@@
 such that $U_{a}^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2}\in\mathring{\mathcal{M}}$ is an extremal metric with $U_{a}=u_{a}$ on $\mathbb{S}^{n}$.
-Indeed, the trick in authors [11, Section 6] can be applied to derive an explicit representation formula of
+Indeed, the trick in authors [15, Section 6] can be applied to derive an explicit representation formula of
 $U_{a}(\xi)=\Big{(}\frac{1-|a|^{2}}{|a|^{2}|\xi|^{2}-2a\cdot\xi+1}\Big{)}^{\frac{n-3}{2}}+\frac{n-3}{4}(1-|\xi|^{2})\Big{(}\frac{1-|a|^{2}}{|a|^{2}|\xi|^{2}-2a\cdot\xi+1}\Big{)}^{\frac{n-1}{2}}.$
@@
 $\sup_{\begin{subarray}{c}0\not\equiv u\in C^{\infty}(\mathbb{R}^{n})\\ u=O(|x|^{3-n})\mathrm{~{}as~{}}|x|\to\infty\end{subarray}}\frac{\|E_{4}(u)\|_{L^{\frac{2(n+1)}{n-3}}(\mathbb{R}^{n+1}_{+})}}{\|u\|_{L^{\frac{2n}{n+3}}(\mathbb{R}^{n})}}.$ (6.7)
-Dou-Zhu [12, Theorem 1.4] states that every maximizer for (6.7) has the form of
+Dou-Zhu [16, Theorem 1.4] states that every maximizer for (6.7) has the form of
 $c(|x-x_{0}|^{2}+d^{2})^{-\frac{n+3}{2}},\qquad c,d\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}.$
-Combining the above with authors [11, Section 6], we conclude from (6.6) that, modulo a positive constant,
+Combining the above discussion and the authors [15, Section 6], we conclude from (6.6) that modulo a positive constant,
 $2P_{3}U=\mathscr{B}^{3}_{3}(U)=\frac{n(n-2)(n-4)}{4}u_{a}^{\frac{n+3}{n-3}}\quad\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n},$
@@
 $\Box$
+7 Perspectives
+Similar to the boundary Yamabe problem (see for example Chen-Sun [14] and Escobar [17, 18]), it is natural to ask whether the round metric on a geodesic ball $\Sigma_{r}^{n+1}$ with radius $r\in(0,\pi)$ in $\mathbb{S}^{n+1}$ is the unique conformal metric (up to a conformal diffeomorphism) of constant $Q$-curvature and constant boundary $T$-curvature under certain geometric constraints on the pair of $T$-curvature constants.
+Consider $\Sigma_{r}^{n+1}$ equipped with the round metric $g_{c}:=g_{\mathbb{S}^{n+1}}$.
+Using the precise formulae of $T$-curvatures in [9, p.296] and $Q$-curvature in [15], we directly compute that $Q_{g_{\mathrm{c}}}=Q_{\mathbb{S}^{n+1}}=\frac{(n+1)((n+1)^{2}-4)}{8}$ and
+$(T^{3}_{1})_{g_{\mathrm{c}}}=$ $h_{g_{\mathrm{c}}}=\cot r,$
+$(T^{3}_{2})_{g_{\mathrm{c}}}=$ $(n-1)\big{(}\frac{1}{2}+h_{g_{\mathrm{c}}}^{2}\big{)},\qquad(T^{3}_{3})_{g_{\mathrm{c}}}=\frac{n^{2}-1}{2}h_{g_{\mathrm{c}}}\big{(}\frac{3}{2}+h_{g_{\mathrm{c}}}^{2}\big{)}.$ (7.1)
+Notice that $\mathbb{B}^{n+1}$ is conformal equivalent to $\Sigma_{r}^{n+1}$, see for example, the first author and Sun [14, pp.8-9].
+Based on (7), we can transfer the above problem into the following conjecture.
+Conjecture..
+For $n\geq 4$ and $i\in\{1,2\}$, let $U\in C^{4}(\overline{\mathbb{B}^{n+1}})$ be a positive smooth solution to
+$\begin{cases}\displaystyle\quad\Delta^{2}U=\frac{n-3}{2}Q_{\mathbb{S}^{n+1}}U^{p^{*}}&\mathrm{~{}~{}in~{}~{}}\quad~{}~{}\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(U)=c_{i}\mathbb{T}^{3}_{i}U^{p^{*}_{i}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad~{}~{}\mathbb{S}^{n},\\ \displaystyle\mathscr{B}_{3}^{3}(U)=c_{3}\mathbb{T}^{3}_{3}U^{p^{*}_{3}}&\mathrm{~{}~{}on~{}~{}}\quad~{}~{}\mathbb{S}^{n},\end{cases}$ (7.2)
+with $c_{k}\in\mathbb{R},k\in\{1,2,3\}$ and
+$\begin{cases}\displaystyle c_{3}=\tfrac{3}{2}c_{1}+c_{1}^{3}\qquad&\mathrm{for}\qquad i=1,\\ \displaystyle c_{2}\geq\tfrac{1}{2},\quad c_{3}^{2}=(c_{2}-\tfrac{1}{2})(1+c_{2})\qquad&\mathrm{for}\qquad i=2.\end{cases}$
+Then $U$ is of the form
+$U(\xi)=\left(\frac{2\varepsilon(1-|\xi_{0}|^{2})}{(\varepsilon^{2}+1)(|\xi_{0}|^{2}|\xi|^{2}-2\xi_{0}\cdot\xi+1)+(1-|\xi_{0}|^{2})(|\xi|^{2}-1)}\right)^{\frac{n-3}{2}}$ (7.3)
+for some $\xi_{0}\in\mathbb{B}^{n+1}$ and
+$c_{1}=\cot r,\quad c_{1}=\frac{\varepsilon^{2}-1}{2\varepsilon}\qquad\mathrm{~{}~{}for~{}~{}}\qquad i=1,$
+$c_{2}=\frac{\varepsilon^{4}+1}{4\varepsilon^{2}}\qquad\mathrm{~{}~{}for~{}~{}}\qquad i=2.$
+This conjecture can be regarded as an extension of Corollary 1.1 to the constant $Q$-curvature and constant $T$-curvature equations on $\mathbb{B}^{n+1}$.
+Moreover, we can prove a uniqueness theorem of positive solutions to a general 4-th order ODE including the BVP (7.2) for radial solutions.
+See Theorem B.1, whose proof is left to Appendix B.
+On the other hand, it is direct to verify that the radial function $\left(\frac{2\varepsilon}{\varepsilon^{2}+|\xi|^{2}}\right)^{\frac{n-3}{2}}$ corresponding to $\xi_{0}=0$ in (7.3) satisfies (7.2).
+Hence we confirm the above conjecture for all radial solutions.
+We hope that the biharmonic Poisson kernel and Green function will come into play in the resolution of the above conjecture.
 Appendix A Appendix: Proof of Proposition 3.1
-The combination of Theorem 1.2 and [10, Theorem 1(3)] demonstrates that the Green functions of fractional GJMS operators $P_{1}$ and $P_{3}$ are
+The combination of Theorem 3.1 and [13, Theorem 1(3)] demonstrates that the Green functions of fractional GJMS operators $P_{1}$ and $P_{3}$ are
 $G_{P_{1}}(\xi,\eta)=\frac{2}{(n-1)|\mathbb{S}^{n}|}|\xi-\eta|^{1-n},\quad G_{P_{3}}(\xi,\eta)=2\bar{P}_{3}^{3}(\xi,\eta),\qquad\xi,\eta\in\mathbb{S}^{n}.$
@@
 $=$ $\int_{\mathbb{S}^{3}}\bar{P}^{3}_{1}(\xi-\eta)\bar{f}_{1}(\eta)\mathrm{d}V_{{\mathbb{S}^{3}}}(\eta)-\frac{1}{2|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}\log|\xi-\eta|\bar{f}_{3}(\eta)\mathrm{d}V_{\mathbb{S}^{3}}(\eta),\quad\forall~{}\xi\in\overline{\mathbb{B}^{4}}.$
-Using the elementary formula (3.8) and conformal changes we obtain
-$\frac{(1-|\xi|^{2})^{2}}{2|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\bar{f}_{1}(\eta)}{|\xi-\eta|^{n+1}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta)=\frac{2}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}f_{1}(y)}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\mathrm{d}y.$
+Using the elementary formula (3.8) we obtain
+$\frac{(1-|\xi|^{2})^{2}}{2|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\bar{f}_{1}(\eta)}{|\xi-\eta|^{n+1}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta)=\frac{2}{|\mathbb{S}^{n}|}\left(\frac{(1+t)^{2}+|x|^{2}}{2}\right)^{\frac{n-3}{2}}\int_{\mathbb{R}^{n}}\frac{t^{2}f_{1}(y)\mathrm{d}y}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}.$
 Since $|\xi|\to 1$ means $t\to 0$, we have
@@
 $u(x,t)=$ $\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}f_{0}(y)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y-\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{tf_{2}(y)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y$
-$=$ $[P^{3}_{0}\ast f_{0}+P^{3}_{2}\ast f_{2}](x).$
+$=$ $[P^{3}_{0}\ast f_{0}+P^{3}_{2}\ast f_{2}](x,t).$
 Then it is not hard to see that
@@
 which leads to
-$2P_{1}\mathscr{B}^{3}_{1}U=\mathscr{B}^{3}_{2}U\qquad\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}.$
+$2P_{1}\mathscr{B}^{3}_{1}U=\mathscr{B}^{3}_{2}U\qquad\mathrm{~{}~{}on~{}~{}}\quad\mathbb{S}^{n}.$
 $\Box$
+Appendix B Appendix: Uniqueness of positive solutions to fourth-order ODE
+The following uniqueness theorem of the BVP for a fourth-order ODE is of independent interest.
+Theorem B.1.
+Let $n\geq 4$ and $i\in\{1,2\}$. Assume that there is a positive radial solution $U\in C^{\infty}(\overline{\mathbb{B}^{n+1}})$ to
+$\begin{cases}\displaystyle\quad\Delta^{2}U=U^{p^{*}}\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{B}^{n+1},\\ \displaystyle~{}\mathscr{B}^{3}_{i}(U)=C_{i}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\mathbb{S}^{n},\\ \displaystyle~{}\mathscr{B}^{3}_{3}(U)=C_{3}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\mathbb{S}^{n}.\end{cases}$ (B.1)
+Here, $C_{k}\in\mathbb{R},k\in\{1,2,3\}$. Then $U$ is unique.
+Proof..
+Denote by $U=C_{0}>0$ on $\mathbb{S}^{n}$. Let $r=|\xi|\in(0,1]$ and $U=U(r)$. Under the change of variable $t=-\log r\in[0,\infty)$, the conformal metric is
+$U(r)^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2}=U(e^{-t})^{\frac{4}{n-3}}e^{-2t}(\mathrm{d}t^{2}+g_{\mathbb{S}^{n}}):=V(t)^{\frac{4}{n-3}}g_{\mathrm{cyl}},$
+where $g_{\mathrm{cyl}}=\mathrm{d}t^{2}+g_{\mathbb{S}^{n}}$ is the product metric on the cylinder $[0,\infty)\times\mathbb{S}^{n}$ and
+$V(t)=e^{\frac{(3-n)t}{2}}U(e^{-t}),~{}~{}t\in[0,\infty)\qquad\Longleftrightarrow\qquad U(r)=r^{\frac{3-n}{2}}V(-\log r),~{}~{}r\in(0,1].$ (B.2)
+Then the equation $\Delta^{2}U=U^{p^{*}}$ is geometrically interpreted as
+$1=P_{4}^{U(r)^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2}}(1)=$ $U(r)^{-\frac{n+5}{n-3}}P_{4}^{|\mathrm{d}\xi|^{2}}(U)=U(e^{-t})^{-\frac{n+5}{n-3}}P_{4}^{e^{-2t}g_{\mathrm{cyl}}}(U(e^{-t}))$
+$=$ $U(e^{-t})^{-\frac{n+5}{n-3}}e^{\frac{n+5}{2}t}P_{4}^{g_{\mathrm{cyl}}}(e^{\frac{3-n}{2}t}U(e^{-t}))$
+$=$ $V(t)^{-\frac{n+5}{n-3}}P_{4}^{g_{\mathrm{cyl}}}(V),$
+equivalently,
+$P_{4}^{g_{\mathrm{cyl}}}(V)=V^{\frac{n+5}{n-3}}.$
+Notice that
+$Q_{g_{\mathrm{cyl}}}=\frac{(n+1)^{2}(n-3)}{8},\quad P_{4}^{g_{\mathrm{cyl}}}(V)=\left(\partial_{t}^{4}-\frac{(n-1)^{2}+4}{2}\partial_{t}^{2}+\frac{n-3}{2}Q_{g_{\mathrm{cyl}}}\right)V.$
+So, $V=V(t)$ satisfies
+$V^{(4)}-\frac{n^{2}-2n+5}{2}V^{\prime\prime}+\left(\frac{(n+1)(n-3)}{4}\right)^{2}V=V^{p^{*}},\quad t\in[0,+\infty).$ (B.3)
+Similarly, by the conformal invariance of $\mathscr{B}^{3}_{k},k\in\{1,2,3\}$ and the fact that $U(1)=V(0)$, the same argument yields
+$\mathscr{B}^{3}_{k}(V)=C_{k}\qquad\mathrm{~{}~{}on~{}~{}}\quad\{0\}\times\mathbb{S}^{n}.$
+Since the boundary $\{0\}\times\mathbb{S}^{n}$ is totally geodesic with respect to $g_{\mathrm{cyl}}$, it is easy to compute
+$\mathscr{B}^{3}_{1}(V)=$ $-\partial_{t}V;$
+$\mathscr{B}^{3}_{2}(V)=$ $\Big{(}\partial_{t}^{2}+\frac{(n+1)(n-3)}{4}\Big{)}V;$
+$\mathscr{B}^{3}_{3}(V)=$ $\Big{(}\partial_{t}^{3}-\frac{3n^{2}-6n+7}{4}\partial_{t}\Big{)}V;$
+using the explicit formulae $\mathscr{B}^{3}_{k},k\in\{1,2,3\}$ on manifolds with boundary; see [9, pp.295-296].
+Hence, collecting these facts together we conclude that
+$\begin{cases}\displaystyle V^{(4)}-\frac{n^{2}-2n+5}{2}V^{\prime\prime}+\left(\frac{(n+1)(n-3)}{4}\right)^{2}V(t)=V^{p^{*}}&\mathrm{~{}~{}in~{}~{}}[0,+\infty),\\ \displaystyle-V^{\prime}(0)=C_{1},\quad V^{\prime\prime\prime}(0)-\frac{3n^{2}-6n+7}{4}V^{\prime}(0)=C_{3}&\mathrm{~{}~{}for~{}~{}}i=1,\\ \displaystyle V^{\prime\prime}(0)+\frac{(n-3)(n+1)}{4}V(0)=C_{2},\quad V^{\prime\prime\prime}(0)-\frac{3n^{2}-6n+7}{4}V^{\prime}(0)=C_{3}&\mathrm{~{}~{}for~{}~{}}i=2,\end{cases}$ (B.4)
+and $V(0)=C_{0}$.
+Alternatively, problem (B.4) of $V$ can be derived by direct calculations. We leave it as an exercise for the readers.
+It reduces to showing the uniqueness of positive solutions to (B.4).
+In the following, we adapt the strategy of Frank-König [19, Proposition 4] or van den Berg [3, Theorem 1] to the uniqueness result of (B.4).
+However, the 4-order ODE studied in [3] or [19] is defined in $\mathbb{R}$, and their proof hinges on the symmetry of $\mathbb{R}$, so it is impossible to apply their results directly to our setting.
+Suppose not, there exists another positive smooth solution $W=W(t)$ to (B.4) with $W(0)=C_{0}$, which arises from another distinct positive radially symmetric solution to (B.1) through the change of function (B.2).
+Rewrite equation (B.3) of $V$ as
+$(\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}-\mu)(\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}-\lambda)V=V^{p^{\ast}}$
+with
+$\lambda=\big{(}\frac{n-3}{2}\big{)}^{2},\quad\mu=\big{(}\frac{n+1}{2}\big{)}^{2}.$
+We introduce an auxiliary function by
+$\varphi(t)=(V-W)^{\prime\prime}(t)-\lambda(V(t)-W(t)).$
+Notice that $V(0)=W(0)$.
+For $i=1$, the boundary conditions give
+$V^{\prime}(0)=W^{\prime}(0),~{}~{}V^{\prime\prime\prime}(0)=W^{\prime\prime\prime}(0).$
+Without loss of generality, we assume $V^{\prime\prime}(0)>W^{\prime\prime}(0)$. This implies
+$\varphi(0)>0,\quad\varphi^{\prime}(0)=0.$ (B.5)
+For $i=2$, the boundary conditions yield that
+$V^{\prime\prime}(0)=$ $W^{\prime\prime}(0),~{}~{}$
+$V^{\prime\prime\prime}(0)-\frac{3n^{2}-6n+7}{4}V^{\prime}(0)=$ $W^{\prime\prime\prime}(0)-\frac{3n^{2}-6n+7}{4}W^{\prime}(0).$
+Without loss of generality, we assume $V^{\prime}(0)>W^{\prime}(0)$. This yields
+$\varphi(0)=0,\quad\varphi^{\prime}(0)=(V-W)^{\prime\prime\prime}(0)-\lambda(V-W)^{\prime}(0)=\frac{n^{2}-1}{2}(V-W)^{{}^{\prime}}(0)>0.$ (B.6)
+In both cases, we deduce that $V(t)>W(t)$ near zero. Then we define
+$\sigma:=\sup\{s;V(t)>W(t),t\in(0,s)\}\quad\mathrm{~{}~{}and~{}~{}}\quad\sigma\in\mathbb{R}_{+}\cup\{+\infty\}.$
+Thus we deduce from the equations of $V$ and $W$ that
+$V^{p^{*}}-W^{p^{*}}=$ $\left(\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}-\mu\right)\left(\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}-\lambda\right)(V-W)=\varphi^{\prime\prime}-\mu\varphi.$
+This together with the definition of $\sigma$ yields
+$\varphi^{\prime\prime}(t)-\mu\varphi(t)>0\qquad\mathrm{~{}~{}in~{}~{}}\quad(0,\sigma).$ (B.7)
+Next we claim that
+$\varphi(t)>0\qquad\mathrm{in}\qquad(0,\sigma).$
+To this end, it follows from either (B.5) or (B.6) that $\varphi(t)>0$ near zero. Then we define
+$t_{0}:=\sup\{T\in(0,\sigma);\varphi(t)>0,t\in(0,T)\}.$
+It suffices to show $t_{0}=\sigma$.
+Otherwise, if $t_{0}<\sigma$, then $\varphi(t_{0})=0$ by definition of $t_{0}$.
+Then we multiply both sides of (B.7) by $\varphi$ and integrate by parts over $(0,t_{0})$ to show
+$-\int_{0}^{t_{0}}(\varphi^{\prime}(s))^{2}\mathrm{d}s=\int_{0}^{t_{0}}\varphi^{\prime\prime}(s)\varphi(s)\mathrm{d}s>\mu\int_{0}^{t}\varphi^{2}(s)\mathrm{d}s,$
+where we also used either (B.5) or (B.6). A contradiction! This finishes the proof of the above claim.
+The above claim asserts that
+$(V-W)^{\prime\prime}-\lambda(V-W)>0\qquad\mathrm{in}\qquad(0,\sigma).$
+Set $\Psi=V-W$ and then $\varphi=\Psi^{\prime\prime}-\lambda\Psi$. Then the above inequality becomes
+$\Psi^{\prime\prime}-\lambda\Psi>0\qquad\mathrm{~{}~{}in~{}~{}}\quad(0,\sigma),$ (B.8)
+equipped with the boundary conditions that
+$\Psi(0)=\Psi^{\prime}(0)=0\qquad\mathrm{for}\qquad i=1,$
+$\Psi(0)=0,\Psi^{\prime}(0)>0\qquad\mathrm{for}\qquad i=2.$
+By definition of $\sigma$, we know that $\Psi(t)>0$ in $(0,\sigma)$. For any fixed $t\in(0,\sigma)$, we multiply both sides of (B.8) by $\Psi$ and integrate by parts to show
+$\Psi^{\prime}(t)\Psi(t)>\int_{0}^{t}(\Psi^{\prime}(s))^{2}\mathrm{d}s+\lambda\int_{0}^{t}(\Psi(s))^{2}\mathrm{d}s.$
+This leads to
+$\Psi(t)>0,\quad\Psi^{\prime}(t)>0,\quad\Psi^{\prime\prime}(t)>0\qquad\mathrm{~{}~{}in~{}~{}}\quad(0,\sigma).$
+Therefore, we conclude that $\sigma=+\infty$. Otherwise, by definition of $\sigma$ we have $\Psi(\sigma)=0$, and arrive at a contradiction.
+For $\sigma=+\infty$, we repeat the same argument to show
+$\Psi(t)>0,\qquad\Psi^{\prime}(t)>0,\qquad\Psi^{\prime\prime}(t)>0\qquad\mathrm{~{}~{}in~{}~{}}\quad(0,+\infty).$
+This implies $\lim_{t\to+\infty}\Psi(t)=+\infty$. On the other hand, we deduce from the change of function in (B.2) that $V,W$ are bounded in $[0,\infty)$. A contradiction! $\Box$